A Complete Bibliography of Publications in Fundamenta Informaticae: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.02

Title word cross-reference

(1, 1) [837]. (1, 2) [188]. (3, g) [486]. (t, n) [830]. $-Calculus [590]. 1 [721, 240, 372]. 2 [1320]. 3 [1320, 1388, 370, 426, 222]. 4 [1036]. ^{∗} [410]. ^{¬} [1202]. _{−G} [897]. C [1144]. C^{1} [440]. χ [15]. δ [311]. F [992]. g [486]. H [137, 509]. hp [1306]. K [1196, 176, 1296, 619, 1277]. L^{3} [852]. λ [854, 530, 375, 1032]. m [1282]. mc [1282]. μ [203]. N [295, 704, 720, 1028]. [1187]. Ω [1282, 325, 125, 1007, 460, 1008, 719, 518, 1016]. P [160, 352, 764, 161, 1280, 1309, 488, 327, 1121, 784, 163, 712, 1122, 980, 824, 1008, 166, 721, 662, 375, 169, 505, 170, 771, 1123, 174, 175, 514, 728, 706, 177, 178, 1232, 537, 1124, 1395, 1125, 1126, 141]. ∂ [525]. π [634]. π_{1}^{1} [84]. q [313]. S [888]. T [391, 1010, 1232]. W [85]. X [160, 502]. Z [1150].

* [868]. *-continuous [868].

-algebras [1202]. -Anonymization [1277]. -augmentation [1010].

-automaton [1196]. -axioms [295]. -Bayesian [188]. -Bit [240]. -Bit-Communication [372]. -Cages [486]. -calculi [854]. -Calculus [375, 634]. -Communication [1232]. -complete [84]. -Definability [530]. -Deterministic [721]. -expression [1196]. -Finite [1306]. -Grams [313]. -Homogeneous [1036]. -Languages [1008, 1016]. -Logic-based [992]. -Machine [502]. -Machines [160]. -Matching [311]. -Means [1144]. -NN [619]. -norm [391]. -Normal [141]. -Operators [85]. -Power [460]. -Process [125]. -Rational [720]. -Regular [719]. -regularities [1032]. -rewriting [518]. -Subset [176]. -Threshold [830]. -Valued [1296, 222, 426]. -Views [704]. -Words [1007].

1 [556, 141, 1394].

2 [605, 141, 1405]. 2004 [746]. 2007 [1078]. 2008 [1326]. 2CNF [90].

3DM [1221].

= [1058].

Abduction [1024]. Abstract [835, 401, 563, 402, 740, 277, 331, 769, 964, 841, 839, 587]. Abstract-state [841]. Abstraction [1167, 1159, 351, 138, 698]. Abstractions [1189, 403]. AC [1302]. Accelerator [1353]. Accepting [426]. Acceptors [894]. Access [807, 714, 1261, 1049, 1330]. Accessing [1024]. Accuracies [1219]. Acquire [798]. Acquisition [467, 468, 390, 1211]. Action [613, 144, 145, 58, 868, 214, 152, 1105]. Action-at-a-distance [613]. Actions [123, 435, 1373, 1345, 1001, 339, 344, 1348, 1045]. Active [352, 820, 327, 154]. Activity [1177]. ACTLS [469]. Ad [1179]. Ad-Hoc [1179]. Adapt [147]. Adaptive [458, 1147, 1201, 844, 737, 1306, 1307, 828, 906]. Adaptively [1052]. Addition [691]. Additive [952, 871, 1128]. Address [320]. Addressing [1137]. Adhesive [747, 748]. Adjective [1252]. Admission [1337]. Advances [1379, 1371, 812, 543]. Affine [528, 1264, 855, 846]. After [726, 1336]. Again [560]. Against [401]. Agency [145, 152]. Agent [474, 1059, 1345, 647, 278, 910, 45, 46, 1348, 473, 648, 298]. Agent-Oriented [46]. Agents [613, 677, 1284, 145, 345, 590, 681, 985, 475]. Agglomeration [1307]. Aggregation [290, 1273, 975, 432]. Agreement [1179, 446, 334]. Aiding [1387]. AIT [596]. ALC [1377, 1297]. Algebra [734, 639, 435, 656, 1318, 274, 713, 1063, 331, 1286, 858, 1341, 250, 428, 968, 1212, 598, 23, 483]. Algebraic [527, 303, 481, 482, 711, 889, 38, 374, 758, 867, 1108, 1311, 641, 1391, 371, 874, 1089, 1394, 1405, 1027, 261, 938]. Algebras [842, 993, 481, 1198, 972, 755, 756, 95, 757, 1401, 88, 1289, 69, 797, 965, 595, 964, 702, 805, 615, 1202, 1187]. Algorithm [806, 1352, 934, 809, 1272, 1064, 718, 455, 544, 1399, 684, 1098, 636, 1253, 316, 1033, 911, 1080, 320, 844, 600, 947, 11, 1269, 47, 1213, 601, 846, 1368, 1129, 270, 1207, 1125, 1308, 390, 1093].

Algorithmic [851, 1002, 1272, 1058, 1142, 1025]. Algorithms [1030, 1029, 1094, 941, 1228, 315, 1286, 181, 126, 1018, 1071, 1398, 462, 1384, 312, 172, 384, 1314, 1185, 300, 416, 8, 1031, 1134, 449, 216, 1403, 240, 775, 283, 105, 195, 1032, 142, 834, 309]. Alias [1149]. Alignment [316]. Almost [488, 935, 1296, 363]. Alpha [272]. Alphabet [522]. Alphabetic [1399]. Alphabets [616]. ALTARICA [459]. Alternating [365]. Alternative [1094, 864]. Alternatives [1046]. Ambiguity [803]. Ambitious [283]. Among [482]. Analysis [1191, 1344, 36, 765, 207, 1063, 1285, 1249, 787, 86, 103, 646, 40, 408, 758, 467, 468, 25, 944, 878, 1216, 670, 1387, 570, 945, 1110, 508, 46, 1307, 515, 300, 573, 1044, 1055, 65, 1170, 919, 762, 1153, 121, 421, 195, 644, 422, 898, 196, 1383]. Analytic [391, 189]. Analyzing [1283, 51]. Andrzej [949]. Animated [850]. ANN [1094]. Annotated [26]. Annotation [1382]. Anonymization [1277]. Answer [1375, 347, 995, 839]. Answers [331]. Ant [588, 844]. Antidictionaries [495]. Antimirov [79]. Antiport [514]. Antoni [923]. any [958]. Aperiodic [1010]. Application [836, 882, 1164, 1127, 1106, 1334, 819, 781, 146, 752, 982, 209, 1320, 547, 624, 187, 844, 368, 975, 49, 1134, 264, 997, 247, 1336, 885, 1242]. Applications [856, 847, 1002, 1379, 778, 1131, 231, 496, 680, 549, 1288, 1338, 524, 661, 438, 3, 102, 50, 1015, 270]. Applied [5]. Applying [483, 1300, 287]. Appointment [46]. Approach [779, 1059, 1060, 1316, 1328, 864, 589, 710, 754, 810, 1260, 465, 1064, 755, 756, 890, 849, 1267, 539, 1067, 115, 1195, 648, 1324, 1355, 1356, 1333, 1141, 1113, 387, 1298, 1142, 297, 263, 1358, 34, 433, 578, 701, 520, 247, 259, 324, 266]. Approaches [1316, 1383, 1216, 1053, 801]. Approximate [338, 786, 705, 1200, 1053, 213, 265, 652, 919, 580, 135]. Approximating [95, 1098, 83]. Approximation [1028, 1318, 407, 568, 683, 904, 1071, 462, 618, 387, 664, 916, 800, 442, 282, 699, 283, 921, 1090, 356]. Approximations [378, 620, 210, 254, 703]. Araucaria} [977, 774]. Arbitrary [1226, 778, 591, 1098, 438]. Arc [1157]. Architectural [647]. Architecture [242, 1048, 649, 743, 215, 1186, 270]. Architectures [739, 454]. Arcs [1160, 194]. Areas [954, 617]. Arithmetic [1401, 1042, 969]. Arithmetical [854]. Arithmetics [959]. Arrangement [601]. Array [498, 870, 321, 1026, 1232]. Arrays [1129]. Arrow [778]. Artificial [1127, 1379]. ASM [838]. Aspects [474, 161, 1086]. Assembly [780, 1050, 724]. Assertions [1055]. Assigned [764]. Assignment [1041, 807, 1049, 975]. Association [381, 1269, 1314]. Associative [47, 1185, 417, 575, 697]. Associativity [1246]. Asymptotic [522]. Asymptotics [439]. Asynchronous [1165, 740, 284, 552, 193, 859, 627, 1169, 33, 367, 431, 861, 427, 1336, 938, 195, 631, 478, 192]. Asynchrony [985]. Attachment [1034]. Attack [446]. Attacks [1291]. Attempt [147, 958]. Attractor [1101]. Attribute [572, 434, 598, 1093]. Attributed [748]. Attributes [140]. Audio [1385]. augmentation [1010]. Australasian [1029]. Authentication [1197, 907, 1402, 692, 1205, 604]. Authorization [258]. Autoepistemic [896]. Automata [1191, 1127, 77, 1236, 342, 123, 205, 1309, 1131, 185, 971, 1258, 1239, 230, 710, 231, 493, 232, 208, 248, 224, 717, 982, 983, 1001, 1065, 1007, 1130, 870, 871, 1128, 931, 424, 1263, 233, 365, 686, 908, 1073, 1075, 503, 1033, 689, 81, 1293, 1294, 913, 1295, 366, 43, 129, 411, 368, 873, 1101, 369, 235, 227, 1339, 1133, 1256, 989, 1084, 371, 874, 246, 1085, 738, 875, 300, 1103, 1134, 238, 239, 363, 630, 450, 240, 876, 241, 364, 301, 897, 422, 583, 922, 1246, 234, 362, 370, 236]. Automata-Theoretic [248]. Automated [30, 414]. Automatic [9, 1159, 559, 1332, 1195, 570, 1382, 1056, 1392, 1144]. Automating [859]. Automaton [926, 1023, 1313, 427, 372, 229, 289, 1196]. Autonomous [219, 42, 1268]. Average [198]. Avoiding [1335]. Aware [880, 1149]. Awareness [500]. Axiom [322, 1202, 1365, 1044]. Axiomatic [566, 1187, 702, 1405]. Axiomatisation [99]. Axiomatization [842, 78, 186, 618]. Axioms [620, 221, 295].

B [836, 625]. Back [560, 603]. Backbones [725]. Backtracking [917]. Bag [1117]. Baire [536]. Balance [837, 1242, 802]. Balancing [581]. Barabasz [945]. Barycentric [964]. Based [1226, 36, 391, 842, 99, 1342, 1329, 780, 453, 941, 124, 354, 436, 538, 1317, 754, 843, 830, 809, 1147, 979, 208, 68, 1097, 786, 393, 680, 890, 1381, 126, 113, 244, 39, 559, 1066, 904, 225, 1201, 905, 744, 115, 1273, 1252, 380, 736, 1099, 1049, 1253, 1264, 1321, 608, 1077, 823, 688, 1216, 627, 1109, 1402, 118, 1295, 1354, 305, 844, 600, 669, 1254, 1110, 1356, 946, 1255, 1391, 1341, 1141, 1133, 1257, 1213, 516, 622, 707, 801, 816, 1056, 1392, 1204, 52, 708, 1370, 1274, 579, 358, 581, 617, 1205, 326, 135, 470, 665, 833, 1144, 922, 1125, 1308, 1208]. Based [390, 834, 478, 951, 259, 1372, 1328, 934, 755, 756, 1385, 748, 16, 1018, 1320, 1068, 1118, 547, 1291, 878, 1177, 1021, 1194, 1085, 100, 700, 359, 1212, 1105, 1153, 121, 1145, 422, 583, 992]. Bases [737, 140, 341]. Basic [999, 713, 902, 1203, 556, 605, 450, 309]. Basins [1101]. Basis [873]. Bayesian [355, 183, 1178, 188, 629]. BCCSP [1146]. BCH [1033]. BDD [191, 688]. be [850]. Beam [803]. Behaved [141]. Behavior [996, 871, 684, 985, 745]. Behavioral [676, 779, 998]. Behaviour [1247, 1180, 534, 1076, 138, 937, 805]. Behavioural [925, 1193]. Behaviours [1358, 939, 1089, 1394, 1405]. Being [798]. Belief [337, 1310, 344, 463, 1312]. Benefit [1323]. Best [412, 1337]. Best-Effort [1337]. Better [313]. Between [1383, 713, 825, 1252, 365, 438, 504, 1358, 222, 364, 786, 316, 798, 27, 883, 992]. Beyond [1346]. BGS [841]. Bhopalator [168]. Bi [1041]. Bi-intuitionistic [1041]. Bilateral [397]. Bilinear [507]. BiLog [928]. Bimachines [730]. Binary [806, 811, 973, 1052, 935, 729, 802, 804, 6, 1205]. Binding [1043]. Biochemical [573]. Biological [1247, 508, 542]. Biorders [532]. Biosyntax [487]. Bipartite [350, 1035]. Birkhoff [1025]. BISECTION [462]. Bisimilarity [1136, 1304]. Bisimulation [243, 1287, 1263, 294]. Bisimulations [1027]. Bit [1137, 240, 372, 876, 538]. Bitrades [1036].

Bitten [1391]. Black [659, 769]. Blind [1007, 305]. Block [736, 640, 866, 1137, 723, 1205]. Blocking [359]. Blocks [1317]. Bodies [826]. Böhm [73]. Boldface [61]. Bond [1135]. Boolean [198, 1063, 814, 1183, 1035, 846]. Boosting [1364]. Border [1252]. Borel [460, 554, 1346]. Bottlenecks [196]. Boundaries [875]. Boundary [864]. Bounded [590, 931, 1340, 476, 366, 532, 212, 298, 469, 897]. Bounds [808, 637]. Box [1139, 817, 284]. Bracketed [504]. Brackets [504]. Brain [1152, 1392]. Branching [712, 1136, 48, 698, 1304]. Bridging [798]. BTC [865]. Büchi [185, 630]. Bucket [1201]. Buffered [631]. Building [1112, 1399, 1332, 18, 1133]. Built [97, 1364, 1220]. Byzantine [1179].

C [175, 270]. Caching [1297]. Cages [486]. CAKE [340]. Calculi [847, 391, 524, 699, 854]. Calculus [162, 563, 675, 527, 528, 284, 39, 59, 438, 375, 1139, 963, 292, 832, 650, 602, 27, 89, 592, 597, 817, 623, 634]. Call [606]. Can [1284, 850]. Cancellation [1146]. Candidate [1356]. Canonical [104, 289]. Capabilities [759]. Capacity [1115]. Cardinality [1028]. Carry [1158]. Carrying [849]. Case [124, 750, 822, 1068, 201]. Case-based [1068]. Cases [1237]. Categorial [181, 307]. Categorical [66, 99, 747, 758]. Categories [37, 748, 531, 703]. Category [1043, 1155]. Catenation [1009]. Caterpillar [246]. Causal [1108, 1055]. Causality [383]. Cause [275]. Cause-Effect [275]. CD [822, 829]. Cell [161, 168, 1368, 240]. Cells [369, 216]. Cellular [1127, 1236, 613, 1131, 231, 750, 1023, 1130, 870, 871, 1128, 424, 233, 365, 872, 362, 366, 367, 368, 873, 1101, 369, 1313, 235, 1133, 371, 874, 1213, 875, 1134, 238, 239, 363, 240, 372, 876, 241, 364, 229, 234, 370, 236]. Center [326]. Center-Based [326]. century [923]. Certificateless [1154]. Chain [1051, 330]. Chains [1094, 1340]. Chaitin [596]. Change [819, 58, 1054, 1145]. Channel [354]. Channels [1280, 931, 744, 1292]. Chaos [1142]. Chaotic [231, 1213]. Characteristics [1015]. Characterization [106, 144, 711, 889, 343, 1101, 256, 702]. Characterized [765]. Characterizing [1135, 511]. ChariTime [46]. Chase [880]. Chaum [688]. Cheating [1259]. Check [402, 1143]. Checker [1078]. Checkers [1332]. Checking [1165, 1167, 277, 30, 455, 1376, 1340, 907, 476, 688, 410, 1295, 1081, 1298, 212, 298, 1085, 629, 698, 457, 839, 301, 469, 897, 583, 922, 906]. Chemistry [486]. Choice [492, 494, 455, 499, 153, 1044]. Chord [439, 440]. Church [962]. CI [1210]. CILC08 [1371]. Circuit [1169, 1035]. Circuits [1165, 627, 367, 771, 555, 451, 427, 938, 192]. Circular [1254]. Class [1112, 1219, 765, 902, 1175, 1151, 187, 1347, 261]. Classes [1037, 751, 350, 1075, 112, 1294, 989, 536]. Classical [1042, 1300, 876, 76, 1301]. Classification [1319, 290, 1067, 225, 1273, 1384, 368, 285, 1113, 694, 816, 1056, 216, 990, 121]. Classifications [1219]. Classifier [1064, 1098, 1048, 1386, 873, 816, 701, 270]. Classifiers [1381, 1364, 255, 845, 117, 1204, 700, 135]. Clausal [1310, 268, 156]. Clause [1199, 612]. Clinical [255, 388]. Clock [1191]. Clockless [1243]. Closeness [1380]. Closure [353, 1183, 171, 1185]. Cloud [1325]. Cluster [260]. Clustering [1147, 465, 600, 947, 1113, 693, 1056, 1392, 1369, 1403]. Clusters [128]. Co [781, 811, 1152]. Co-Entropies [781]. Co-Occurrence [1152]. Co-processor [811]. Coalition [472]. Code [849, 1011, 330]. Codebook [809, 979, 636, 1253, 1143, 1276]. Codebook-linked [1276]. Coded [1056]. Codes [1197, 718, 1033, 772, 1015, 256, 1100]. Codeword [1143]. Coding [658, 1107, 736, 640, 865, 358, 1088]. Codomain [730]. Cognitive [1216, 1210, 1209, 1217, 1214]. Coherence [1302]. Collatz [518]. Collective [145, 274, 220, 328]. Colony [588, 844]. Color [458, 1362, 1148, 640, 617, 444, 833]. Color-Complexity [833]. Colored [1267]. Colorings [347]. Colors [984]. Coloured [1162]. Combination [760]. Combinatorial [1029, 324]. Combinatorics [1397]. Combined [1009, 1393]. Combining [225, 700, 1347]. Commitment [328]. Common [722]. Commonality [131]. Communicating [373, 137, 931, 138, 415, 427]. Communication [740, 859, 88, 1292, 1307, 1111, 1232, 372, 876, 1335, 195, 631]. Communications [552, 240]. Communities [411]. Commutative [912, 249]. Compact [85, 321]. Compactness [463, 976, 961]. Comparability [589]. Comparative [210, 598, 898]. Comparator [29]. Compared [1157]. Comparing [62, 243, 127, 688, 1274]. Comparison [1400, 1316, 1172, 1216, 384, 216, 1301]. Comparisons [1321]. Compensation [708]. Competence [823]. Compilation [461]. Compile [1174]. Complementarity [197]. Complete [352, 78, 1107, 502, 1138, 186, 3, 1343, 703, 28, 785, 84]. Completeness [37, 1272, 848, 687, 203]. Completion [772]. Complex [1040, 1132, 1238, 1401, 1001, 339, 1034, 418, 577, 580, 1308]. Complexity [1002, 302, 71, 191, 1237, 712, 1122, 659, 1006, 68, 789, 114, 1107, 554, 1009, 80, 1203, 1012, 448, 770, 332, 183, 1140, 285, 430, 1015, 1016, 189, 521, 1265, 833, 523, 90, 251, 735]. Complicated [807]. Component [744, 998, 395]. Components [1249, 1195, 745]. Composable [552, 187, 215]. Composed [996]. Composing [925]. Composite [1306]. Composition [734, 1192, 1195, 308, 1339, 395, 893]. Compositional [900, 419]. Compositionality [997]. Compositions [929]. Compound [734, 1074]. Compressed [754, 466, 318]. Compressible [291]. Compression [319, 436, 495, 1261, 1251, 636, 866, 722, 908, 1141, 586, 360]. Compressor [541]. Computability [1013, 141, 157]. Computable [489]. Computably [1017]. Computation [639, 625, 930, 1196, 769, 943, 1243, 22, 142]. Computational [1371, 1259, 1122, 68, 114, 721, 1050, 1244, 773, 1124, 1211]. Computations [750, 304, 1149, 769, 763, 695]. Compute [941]. Computed [717, 1123]. Computer [800, 428, 967]. Computers [940]. Computing [1119, 1120, 317, 930, 822, 73, 242, 811, 167, 647, 1071, 1398, 826, 794, 12, 13, 728, 920, 1209, 1217, 1214, 1032]. Concatenation [67]. Concept [378, 106, 787, 249, 758, 467, 468, 944, 878, 118, 1305, 669, 800, 1143, 1206, 266]. Concepts [864, 127, 683, 46, 1300, 519, 247]. Concerning [715, 430]. Concrete [402, 596]. Concurrency [673, 899, 856, 1164, 1057, 1106, 1334, 1278, 191, 562, 403, 624, 1108, 1109, 279, 1336]. Concurrent [36, 996, 402, 55, 62, 857, 454, 277, 903, 825, 1075, 431, 413, 414, 937, 579, 654, 1089, 1394, 1405]. Conditional [147, 67]. Conditions [752, 554, 437, 823, 1053]. Configurations [980, 363]. Conflict [336, 1194, 695, 801]. Conflict-free [695]. Conflicts [456, 1109, 1335]. Confluence [635]. Confusion [128]. Congestion [1162]. Congruence [1136, 539]. Congruences [1350]. Conjecture [1271]. Conjoinability [425]. Connected [569, 689]. Connectedness [1005]. Connection [713, 825, 1330, 96, 623]. Connections [814, 793, 7, 421]. connectivity [110]. Connector [999]. Conquer [811]. Consecutions [966]. Consecutive [430]. Conserving [233, 229]. Considerations [276]. Consistency [30, 455, 1107, 437]. Consistent [920]. Constitutive [290]. Constrained [1357, 815, 660]. Constraint [1372, 1066, 115, 42, 1339, 1378]. Constraint-Based [115]. Constraints [1372, 146, 667, 1319, 752, 583]. Constructing [549, 11, 21, 268]. Construction [166, 1343, 63, 796, 1269, 1314, 245, 512, 861, 1105, 1208]. Constructions [527, 1230, 592, 597]. Constructive [1373, 1150, 1394]. Constructs [707, 434]. Consumers [437]. Consumption [238]. Contact [864, 755, 756]. Contagious [485]. Content [665]. Content-Based [665]. Context [971, 24, 848, 1006, 1117, 981, 165, 499, 1183, 1160, 726, 63, 82, 1087, 1206, 834]. Context-Free [1006, 499, 971, 63]. Context-Semantics [848]. Context-Sensitive [1183]. Context-Sensitivity [726]. Contextual [66, 62, 492, 494, 498, 499, 503, 504, 169, 505, 170, 506, 510, 512, 513, 1026, 218]. Contingency [1222, 1212]. Continuous [950, 93, 1130, 94, 1340, 140, 451, 450, 868]. Continuous-Time [1340]. Contracts [1192]. Contrary [147]. Contrary-To-Duties [147]. Contribution [884, 967]. Contributions [1188]. Control [940, 1162, 219, 807, 682, 351, 1049, 531, 611, 855, 49, 247, 357, 376, 1337, 266, 286]. Controllable [860]. Controlled [1065]. Controller [17, 987, 578]. Convex [704, 1035, 6]. Convolutions [1030]. Conway [1095]. Cooperating [821, 713, 823, 43, 129]. Cooperation [621]. Cooperative [479, 791]. Coordination [740, 739, 741, 744, 742]. Coping [44]. Copy [1021, 665]. Core [15]. CoreASM [838]. Corecursive [558]. Corner [1356]. Correct [275, 861]. Correct-by-Construction [861]. Correctness [566]. Correlation [1344, 809]. Correspondence [1272, 786]. Corrigendum [1302]. Cosmic [500]. Cost [1323, 1133, 651, 1105]. Cost/Benefit [1323]. Costs [718]. Could [798]. countable [540]. Counter [1239, 1007, 638]. Counter-Free [638]. Counting [72, 314]. Counts [1382]. Coupled [159, 997]. Covariances [898]. Coverings [781, 1071]. Covers [1248]. CP [1178]. CP-Logic [1178]. Create [963]. Creation [1161]. Creativity [485]. Crises [649]. Criteria [1097]. Criterion [599]. Critical [871]. CRL [203].

Cryptanalysis [253]. Cryptographers [688]. Cryptographic [1049]. Cryptography [1259]. Cryptosystem [1145, 672]. CS&P [673, 899, 1057, 1278]. CSC [1335]. CTL [1376, 410, 212]. Cubes [382]. Cubics [440]. Cumulative [439, 440]. Cumulativity [150]. Curve [5, 622, 1145]. Cut [1102, 76, 1220]. Cut-Elimination [76]. Cuts [1367]. Cycle [1128]. Cyclic [1363, 414].

D [1320, 1388, 370]. Dance [1103]. Data [525, 882, 400, 1171, 926, 1361, 754, 1147, 1115, 978, 1148, 1260, 495, 682, 549, 1201, 548, 115, 1151, 943, 408, 661, 880, 686, 1073, 878, 267, 1158, 382, 1322, 1052, 988, 1254, 1022, 1053, 541, 255, 1084, 140, 413, 515, 263, 586, 1170, 700, 918, 654, 1403, 388, 1221, 23, 421, 587, 922]. Data-driven [1221]. Database [258, 667, 1267, 380, 944, 885, 1290]. Database-theoretic [1267]. Databases [119, 584, 1175, 116, 182, 1305, 892, 184, 884]. Dataflow [1238]. Datagram [1162]. Datalog [707]. Datamodels [638]. Dataspace [743]. Davidson [1045]. Davis [888]. Deadlock [154]. Debit [1079]. Deblocking [359]. Debugging [14]. Decent [937]. Decidability [768, 902, 448, 966, 644]. Decidable [273, 1366, 261]. Deciding [288, 888]. Decision [185, 198, 248, 1097, 244, 549, 1332, 1151, 987, 102, 11, 12, 19, 188, 285, 796, 430, 399, 268, 432, 426, 884, 65, 1325, 216, 990, 1104]. Decisions [1323]. Declarative [345]. Decomposed [654]. Decomposing [81]. Decomposition [836, 974, 267, 134, 862]. Deductive [1000, 584, 182, 892, 1218]. Default [90, 251]. Defeasibility [151]. Defeasible [340]. Definability [319, 530]. Defined [1008, 1075, 1294]. Defining [1089, 1394, 1405]. Definite [156]. Definition [1279, 1104, 286]. Definitions [853, 8]. Deforestation [607]. Degree [462, 1222, 1118]. Delay [193, 1287, 860, 626, 367]. Delay-Insensitive [193, 860, 626, 367]. Delays [857]. deliverability [252]. Delivery [648]. Demands [648]. Demarcated [1380]. Denial [446]. Denial-of-Service [446]. Denotational [1209, 1217, 1214, 480]. Dense [1359, 1088, 1389]. Dense-time [1389]. Density [1273, 600]. Density-Based [600]. Deontic [144, 146, 145, 150]. Dependable [22]. Dependence [1222]. Dependencies [620, 394, 1262, 793, 20, 399, 638]. Dependent [592, 597]. Dependently [529]. Depending [285]. Depends [69]. Depth [198, 11]. Derivable [795]. Derivation [322, 1303]. Derivatives [79]. Derived [307]. Deriving [66]. Descent [917]. Described [654]. Describing [400, 199]. Description [342, 819, 1373, 493, 1198, 1377, 25, 995, 1297, 1370, 329, 376, 992]. Descriptional [302, 1140, 286]. Descriptions [444]. Descriptive [189]. Design [856, 1164, 1106, 1334, 17, 624, 1133, 46, 49, 578, 579, 457, 1336, 240]. Designated [410]. Designing [1190, 718, 1381, 775]. Designs [1082]. Detecting [548, 456]. Detection [1011, 312, 1021, 1356, 1194, 1054, 359, 665]. Determinacy [1169]. Determinants [1212]. Determining [678]. Deterministic [185, 304, 721, 365, 503, 724, 1080, 12, 707, 521, 155, 330]. Devaney [1142].

Development [625, 1193, 1286, 1118]. Developments [818]. Devices [872]. devising [370]. DFCA [490]. Di-Algebraic [641]. Diagnosing [287]. Diagnosis [1383, 1168, 1179, 555, 264, 388]. Diagonal [897, 583]. Diagrams [94, 1332, 884]. Dialogue [783]. Dichotomic [544]. Did [963]. Difference [1254, 262]. Differences [883]. Differencing [632, 1022]. Different [191, 1131, 1383, 373, 1172, 107, 290, 1364, 216]. Differential [1352, 5, 1403]. Differentiating [525]. Diffusion [1280, 241]. Digit [1003]. Digital [120, 1248, 1197, 1213, 360, 470, 1276]. Dijkstra [47]. Dimension [200]. Dimensional [1173, 1201, 775, 1236, 231, 365, 241]. Dimensions [510]. Dining [688]. Diophantine [325]. Direct [166, 500, 1080]. Disabling [738]. Disabling/Enabling [738]. Disagreement [788]. Disambiguation [1252, 520]. Disassembly [131]. Discover [963]. Discovering [119, 379, 116, 413]. Discovery [877, 1074, 649, 431, 1331]. Discrepancies [890]. Discrepancy [1357]. Discrete [836, 766, 686, 1073, 1402, 305, 1341, 1084, 965, 817, 922, 1390]. Discrete-time [1390]. Discretization [1316, 1260]. Discs [1248]. Disjunctive [657, 584, 642, 461]. Disorders [1383]. Dispersed [1353]. Disruption [591]. Dissimilar [1147]. Dissimilarity [904]. Dissolution [1124]. Distance [124, 843, 496, 735, 1097, 5, 153, 1090, 613]. distinguishing [1108]. Distortions [1264]. Distributed [400, 423, 821, 941, 56, 713, 137, 42, 823, 45, 211, 411, 509, 1341, 1194, 1257, 214, 281, 282, 444, 1277]. Distribution [1062, 743]. Distributions [1219, 729]. Divergence [599, 1304]. Divide [811]. Divide-and-Conquer [811]. Division [1124]. DNA [735, 1135, 379, 670, 172, 173, 794, 174]. DNA-Worm-Objects [174]. Do [1174]. Doctrine [1210]. Does [500]. Domain [1372, 632, 544, 387, 875, 1221, 1266]. Domain-oriented [1221]. Domain-Specific [544]. Domains [1198]. Dominance [1317, 94, 432]. Dominance-Based [1317]. Dominant [984]. dominated [1018]. Double [1186]. Double-Tree [1186]. Down [81]. Doyle [1365]. Driven [1040, 1339, 1221]. DRM [1190]. DSD [382]. Dual [682, 168, 690, 441, 334, 1205, 149]. Dual-Pair [1205]. Dualities [965]. Duration [27, 443]. during [328]. Duties [147]. DWT [359]. DWT-based [359]. Dynamic [1059, 819, 610, 1168, 101, 538, 1179, 680, 39, 209, 346, 791, 684, 1338, 42, 648, 280, 1324, 211, 411, 385, 845, 117, 1349]. Dynamical [57, 768]. Dynamically [1360]. Dynamics [1242, 371, 801, 542].

Easy [1184]. ECA [993]. Eco [57, 827, 828]. Eco-Grammar [57, 827, 828]. Economics [728, 1223]. ECTL [1091]. Edge [750, 1021, 359]. Edge-based [1021]. Edges [364]. Editing [980]. Effect [275, 359]. Effective [1180, 1230, 1388, 555, 237, 155]. Effects [1149, 1023]. Efficiency [1030, 1050, 1124]. Efficient [1167, 934, 1196, 614, 466, 318, 1332, 323, 661, 1176, 321, 1185, 140, 1297, 846, 881, 920, 775, 876, 672]. Efficiently [382, 545]. Effort [1337]. Eikonal [5]. Eilenberg [160]. Elastic [116]. Election [934, 392, 582]. Electronic [52, 761]. Element [1306, 1307].

Elementary [71, 528, 871, 1160, 412, 194, 329]. Elements [290]. Eliminating [322, 588, 1378]. Elimination [606, 855, 76]. Eliptic [622]. Elliptic [1145]. Embedded [1247, 545]. Embedding [993, 754, 1147, 1148, 393, 769, 751, 736, 367, 1388, 1052, 1255, 1186, 470, 1155]. Emergence [1211]. Emerging [1365]. EMG [1383]. Emphasizing [1381]. Empirical [1216, 651]. Empiricism [100]. Emptiness [887, 630]. Enabling [738]. Encoding [456, 1109, 192]. Encrypting [672]. Encryption [946, 1213]. Energy [764, 168, 238]. Enforce [1284]. Enforcing [807]. Engine [838]. Engineering [1188]. Enhance [403]. Enhanced [740, 446]. Enquiry [837]. Ensemble [816]. Ensembles [1319, 135]. Entish [400]. Entropies [781]. Entropy [136, 1013, 844, 265, 521]. Enumerable [498, 511, 1017]. Environment [1179, 1149, 827, 968]. Environment-aware [1149]. Environmental [1023]. Environments [97, 346, 791, 1386, 626, 74, 1111, 282]. Epistemic [476, 1365, 298]. Equalities [262]. Equality [851]. Equals [720]. Equation [1196, 299, 133]. Equational [595, 868, 268, 592, 597]. Equations [679, 973, 226, 773, 325]. Equilibrium [343]. Equivalence [36, 482, 1287, 155, 1038, 261]. Equivalences [674, 1193, 420]. Equivalent [808, 1302]. Erratum [977]. Error [1011, 1338, 394]. Error-Robust [394]. Essay [800]. Establishment [1330]. Estimating [263]. Estimation [1094, 193, 1320, 1077, 1015, 651]. Ethics [936]. Etna [1134]. Euler [811, 350]. Evaluation [1344, 1267, 1341, 414, 1134]. Event [674, 993, 591, 40, 738, 51, 836]. Event-B [836]. Eventually [334]. Evolution [328, 5, 1403]. Evolutionary [1225, 1228, 1345, 1260, 1348, 1125, 309]. Evolutions [207]. Evolving [1162]. Exact [639, 200]. Example [223]. Examples [849]. Exceptions [1000]. Exchange [1190]. Exclusive [1345]. Execution [890, 838, 346, 1348]. Exercise [625, 937]. Exercises [272]. Exhaustiveness [788]. existence [325]. Existential [1378, 897]. Expansion [1254]. Experiment [275]. Experimental [1172, 431]. Experiments [1248]. Explanations [890]. Explicit [971, 74, 1304]. Exploiting [663, 1082, 398]. Exploring [1128]. Expressible [1076]. Expressing [677]. Expression [882, 164, 917, 1086, 1300, 1196]. Expressions [585, 616, 255, 1256, 314]. Expressiveness [1157, 1292, 855, 1389, 1390]. EXPTIME [785]. EXPTIME-complete [785]. Extended [1316, 331, 646, 1008, 406, 1033, 355, 966, 1350, 364]. Extensible [838]. Extension [459, 1377, 17, 1139, 916, 142, 251]. Extensionality [308]. Extensions [920, 1272]. extensive [1087]. External [510, 13, 1031, 178]. Extracting [1053]. Extraction [565, 845, 388]. Extractive [1207]. Extreme [574].

Face [1320, 441]. facet [1059]. Factorization [1397]. Factors [522]. Failure [243, 1068, 1387, 334]. Fair [1190]. Fairness [44, 412]. Families [1397, 232, 1014, 171]. family [325]. Fast [495, 1253, 316, 312, 1356, 442, 1143]. Faster [718, 1398]. Fault [1168, 1179].

Faults [304, 555]. FC [84]. FC-Normal [84]. Feasibility [238]. Feasible [472]. Feature [1064, 1264, 845, 117, 708]. Features [1364, 833]. Feedback [639, 451]. Feedforward [86, 103]. Few [1159]. Fifth [1106]. Figures [232]. File [442]. Filter [132]. Filtering [313, 442, 700]. Finding [1094, 29, 890, 1308]. Finitary [929, 460]. Finite [37, 1166, 1372, 224, 634, 717, 929, 983, 1065, 1267, 1008, 501, 1075, 1033, 959, 987, 1343, 772, 595, 1306, 1307, 730, 522, 1143, 1090]. Finite-State [929, 987]. Finiteness [330]. Firing [1327, 233, 51, 775, 1233]. First [551, 924, 1019, 32, 692, 1312, 433, 587]. First-Order [924, 1019, 433, 32, 692, 1312]. Five [1233]. Five-State [1233]. Fixed [144, 814, 530, 531, 369]. Fixed-Point [531, 144]. Fixpoint [292, 895, 896]. Flat [1239, 569, 689]. Flavours [1236]. Flexible [590, 572, 1355]. Flow [402, 277, 107, 682, 685, 1069, 515, 1134, 1170]. Flows [1329, 23]. Folding [612, 1378]. forbidden [317]. Forbidding [165]. Force [1383]. Foreword [869, 361, 158, 310]. Fork [209]. Fork-Join [209]. Form [1374, 846, 1265]. Formal [1279, 625, 787, 758, 467, 468, 944, 878, 670, 910, 33, 236, 650, 1270, 1275, 457, 1218, 1223, 473, 286]. Formalisation [15]. Formalising [926]. Formalism [1387, 45, 579]. Formalisms [302, 127]. Formalization [42, 943, 177]. Formalized [1286]. Formalizing [24, 340]. Formation [576, 472]. Forms [191, 423, 507, 159]. Formulas [407, 568, 1067]. Forward [837, 582]. Foundation [55, 867, 844]. Foundational [1192, 849]. Foundations [739, 374, 668, 892, 556, 605, 1209, 1217]. founded [461]. Four [1231, 1038]. Fourier [1402]. Fourth [856]. Fractal [1141, 200, 360]. Fragile [1197]. Fragment [101, 212, 897]. Fragments [924]. Frames [149, 1133]. Framework [851, 882, 1400, 1132, 1318, 529, 566, 338, 786, 680, 747, 1199, 546, 878, 1081, 395, 415, 1393, 449, 1223, 593, 939, 1089, 1394, 1405]. Fredkin [771]. Free [971, 927, 1006, 718, 455, 499, 350, 974, 1203, 608, 638, 98, 78, 322, 1359, 557, 692, 63, 696, 695]. Free-Choice [455]. Frequent [550, 1172, 546, 545]. Friendly [1362]. frontier [288]. Full [808, 321]. Full-Searching-Equivalent [808]. Full-Text [321]. Fully [740]. Function [1344, 1361, 1099, 873, 1313, 846, 357]. Functional [394, 793, 638, 215]. Functions [765, 666, 198, 717, 337, 118, 536, 153, 65, 7]. Fundamenta [2, 1326]. Fundamental [748, 1223]. Fundamentals [1188, 1315]. Further [1250, 505]. Fusion [132]. Future [648]. Fuzzification [831]. Fuzzy [1131, 106, 353, 244, 646, 406, 437, 539, 337, 954, 1118, 80, 956, 1151, 31, 995, 947, 873, 1101, 1152, 1113, 269, 1056, 1144, 898, 1406, 249].

Galois [814, 793, 421]. GALS [1337]. Game [77, 950, 842, 263]. Games [180, 554, 850, 201]. Gap [1013, 798, 27]. Gapped [313]. GAPS [1369]. GASP [1375]. GDPLL [646]. Gene [882, 164, 1050, 724, 255]. General [1400, 1199, 546, 1242, 161, 990]. Generalised [837]. Generalization [1249, 113, 1312, 975, 612]. Generalizations [530, 1022]. Generalized [551, 210, 118, 1311, 893, 630, 419]. Generated [784, 1350]. Generating [804]. Generation [327, 1010, 1253, 611, 82, 1143, 240, 372, 885, 1290].

Generic [1251, 1330]. Genetic [1132, 1094, 493, 1018, 1098, 1384, 600, 601, 1134, 1056]. Genome [722]. Geodesy [110]. Geometric [1264, 1153]. Geometrical [769, 235]. Geometry [71, 99, 507]. Girard [1266]. Girth [486]. Global [681, 1313, 1297, 214]. Gödel [958]. GOLOG [339]. GP [165]. Graded [1376, 407]. Grammar [818, 302, 821, 57, 713, 822, 824, 825, 823, 827, 1306, 917, 1086, 1300, 828, 829, 1218]. Grammars [492, 494, 1303, 1117, 981, 181, 498, 646, 942, 503, 307, 504, 506, 794, 63, 1140, 139, 82, 510, 511, 245, 512, 513, 1026]. Grams [313]. Granularity [1222]. Granule [134]. Granules [1317, 1074]. Granulometries [4]. Graph [1112, 486, 1359, 614, 747, 748, 942, 549, 349, 749, 351, 547, 1035, 1306, 1220]. Graph-based [547]. Graph-Structured [549]. Graph-Theoretical [486]. Graphs [780, 750, 1285, 751, 748, 752, 455, 1107, 350, 546, 59, 462, 347, 235, 795, 915, 515, 64, 26, 919, 762, 862, 543]. Gray [1356]. Gray-Level [1356]. Grayscale [1115, 636, 866, 865, 1099]. Greedy [1099, 1269, 1314, 1093]. Grey [6]. Grid [810, 1186]. Grids [1362]. Ground [77, 1019, 1178, 1350]. Grounding [1375]. Group [1241]. Groups [576]. Growing [1183]. Growth [1034]. Grzegorczyk [949, 960, 967, 969]. Guarantees [810]. Guided [629]. Gumbel [760].

Haines [1230]. Hairpin [670]. half [923]. Halftoning [594]. Halting [1226, 889, 1114]. Hamming [843]. Handling [1181, 1338, 661, 1257, 140, 155, 700]. Hard [1243]. Hardness [22]. Harmonic [886]. Harmony [1352]. Hashing [633]. having [261]. Hazard [1165, 1134]. Help [646]. Heterogeneous [119]. heuristic [629]. Heuristics [311]. Heyting [481, 965]. Hidden [1094, 1241, 879]. Hiding [354, 538, 1361, 632, 843, 830, 1115, 978, 640, 1099, 988, 1254, 1022, 1255, 360]. Hierachy [1049]. Hierarchical [296, 879, 1332, 25, 280, 1208]. Hierarchies [10, 290]. Hierarchy [971, 1037, 807, 1346, 235, 412, 1017]. High [1115, 552, 534, 747, 752, 682, 554, 1151, 187, 988, 1058]. High-Level [552, 534, 747, 752, 682, 187]. High-Performance [988]. Higher [223, 565, 1060, 276, 607, 638]. Higher-Order [276, 638]. Highly [1240]. Higman [766, 1230]. Hillclimbing [601]. Hirschberg [315]. Histogram [1253, 803, 833]. History [1244, 1072, 1163]. HLR [748]. HMM [1094]. HMM/ANN [1094]. HMMs [1094]. Hoc [1179]. HOL. [1044]. Hole [659, 769]. Homage [800]. Homogeneous [827, 1036, 139, 1395]. Horizontal [364]. Horn [1199]. hp [1307]. hp-Finite [1307]. Huffman [466]. Hulls [6]. Human [1383, 798, 358]. Hunt [315]. Hwang [253]. Hwang-Shi [253]. Hybrid [1094, 1064, 947, 873, 939]. Hyper [642, 1213]. Hyper-chaotic [1213]. Hyper-Resolution [642]. Hyperbolic [424, 324, 1313, 237, 370]. Hypergraph [943, 1367]. Hypergraphs [1251]. Hyperplane [432]. Hypertext [1393]. Hypothesis [1174].

ICGT [746]. ID [1049, 1208]. ID-Based [1049, 1208]. Idea [990]. Ideal [113]. Ideals [663]. Identical [77]. Identification [676, 779]. Identifying [1080]. Identity [946, 150]. Identity-Based [946]. II [168, 1217, 482, 756, 468, 1307, 1275, 597, 1222, 896]. III [1272]. Illustrated [643]. ILP [333]. Image [354, 436, 538, 632, 843, 830, 1197, 242, 811, 984, 1251, 1320, 720, 636, 640, 866, 865, 1099, 3, 1402, 669, 1052, 1021, 1141, 1382, 1213, 1367, 1392, 358, 617, 444, 833, 1207, 445]. Image-based [1320]. Imagery [1113, 1056]. Images [458, 754, 1115, 1116, 1148, 736, 25, 1022, 1356, 1152, 1054, 1205, 470, 1276]. Imbalanced [1219, 700]. Immediate [1139]. Immunological [1345, 1348]. Impact [985, 608]. Imperative [930]. Imperatives [148]. Imperfect [383]. Implementation [1282, 626, 47, 1185, 861, 1368, 289]. Implementations [240]. Implementing [1121, 609, 198]. Implication [783, 1055, 266, 113]. Improve [548, 943]. Improved [1352, 1250, 636, 1382, 1035, 862, 1154]. Improvement [358]. Improvements [422]. Improving [48, 1091]. Incidence [602]. Including [226]. Inclusion [1172, 1118, 31]. Inclusion-degree-based [1118]. Incomplete [781, 467, 468, 267, 1322, 202]. Inconsistency [958, 1257]. Inconsistent [1053, 65]. Incremental [627, 1053, 1324, 390]. Indeed [1136]. Indefinite [182]. Independence [620, 1212, 286]. Independency [729]. independent [306]. Index [658, 978, 1201, 541, 321, 442]. Indexed [1317, 1014]. Indexing [120, 1201, 643, 444, 326]. Indicators [1038]. Indices [486]. Indiscernibility [197, 572]. Induced [353, 729, 1090]. Inducing [1324]. Induction [953, 1302, 225, 264, 1325, 651]. Inductive [527, 1228, 384]. Inequality [808]. Inf [85]. Inf-Compact [85]. Infants [1068]. Inference [942]. Inferences [1262]. Infinigons [324]. Infinitary [868]. Infinite [971, 1342, 460, 1340, 616, 561, 464, 285, 291, 64, 200, 644, 1246]. Infinite-State [1340, 644]. Infinity [962]. Influence [635]. Infomorphisms [282]. Informaticae [1326]. Informatics [1210, 1209, 1217, 1214]. Information [66, 197, 778, 402, 781, 599, 146, 1353, 277, 28, 248, 785, 336, 685, 1069, 793, 168, 844, 1324, 1255, 20, 1024, 285, 291, 935, 1296, 371, 50, 200, 155, 761, 134, 386, 577, 920, 990, 1104, 598, 356]. Information-Hiding [1255]. Information/Energy [168]. Ingredients [683]. Inhabitants [72]. Inheritance [66]. Inhibitor [194]. Inhibitory [1296]. Initialized [1357]. Injectivity [179, 1313, 517, 731]. Inner [266, 1104]. Input [136]. Inputs [22]. Insensitive [193, 860, 626, 367]. Insertion [1303, 684, 1109, 511]. inspirational [923]. Inspired [1197, 1125]. Instance [66, 225]. Instance-Based [225]. instances [257]. Instantaneous [123]. Instantiated [1360]. Instantiation [836, 534, 612]. Institution [306]. Institution-independent [306]. Institutionalized [145]. Instruction [1360]. Instructions [1360]. Instrument [886]. Integer [1331]. Integral [1353]. Integrated [628]. Integrating [423, 1372, 1163]. Integration [845, 117, 1325]. Integrity [667, 1058]. Intelligence [1127, 1379, 1351]. Intelligent [120, 346, 515]. Intent [119]. Intentions [220]. Inter [240]. Inter-Cell [240]. Interacting [230, 1268, 411]. Interaction [71, 423, 1243].

Interactions [1023]. Interactive [54, 590, 732, 592, 597]. Intercode [813]. Interconnect [1337]. Interconnection [107]. Interconnections [222]. Interesting [954]. Interface [30, 1001]. Interfaces [925, 373, 998]. Interference [205, 273]. Interleaving [1001]. Intermediate [1073, 239]. Intermolecular [1050]. Internal [927, 499, 593, 1335]. Internet [1162, 1274]. Interplay [1034]. Interpolating [354]. Interpolation [551, 440, 381]. Interpretation [401, 563, 277, 1042, 997, 969, 1212]. interpretations [526]. Interpreted [901, 716]. Intersection [78, 322, 853]. Intersections [778, 887]. Interval [1406]. Interval-Valued [1406]. Intrinsic [332]. Intrinsically [1236]. Introduction [228, 542, 895, 896]. Introns [588]. Introspection [565]. Intrusion [312]. Intuitionistic [24, 101, 1406, 1041]. Invariance [986]. Invariant [1264, 1153]. Invariants [871, 1082]. Inverse [594]. Inversion [559]. Investigating [1219]. Invisible [470]. Invitation [954]. Invited [1058]. Involution [1100]. IP [320]. Irreducible [796, 1335]. Irrelevance [26]. Irrelevant [548]. Isolated [576]. Isolation [886]. Issue [856, 877, 1164, 1106, 1119, 1120, 1334, 1396, 746]. Issueon [624]. Issues [789, 1104]. Italiano [1185]. Iteration [78, 322, 1095, 38, 497]. Iteration-free [78, 322].

Jacobi [806]. Jacopini [75]. Java [1000, 565, 1060, 1282, 1080, 1299, 1092]. Job [1355]. Job-shop [1355]. Join [209, 1100].

KDS [586]. KDS-transformation [586]. Kernels [1046]. Key [807, 1049, 446]. Keys [638]. KGL [18]. Khepera [578]. Kleene [595]. Knapsack [1125, 257]. knapsack-type [257]. Knapsacks [54]. Know [475]. Knowledge [877, 1315, 467, 468, 1252, 1078, 648, 692, 737, 1324, 18, 387, 214, 698, 341, 390, 478]. Knowledge-Based [1252, 478]. Known [323, 622]. Kolmogorov [1015]. Krivine [1042, 1044].

Labeled [546, 184]. Labels [392, 398]. Lambda [847, 528, 524, 59, 89]. Lambek [307]. Language [852, 400, 709, 1279, 1372, 926, 625, 15, 928, 460, 705, 1073, 1075, 670, 171, 1294, 19, 108, 798, 773, 184, 480, 1211, 236]. Languages [991, 950, 971, 1166, 1342, 1237, 739, 231, 784, 38, 125, 1006, 1363, 1096, 496, 497, 981, 498, 1346, 1008, 499, 719, 813, 974, 1203, 1183, 1216, 616, 307, 504, 872, 569, 689, 82, 511, 696, 802, 1016, 330, 1218, 671, 317]. Large [1148, 140]. Larger [1247]. Largest [1359]. Latency [193]. Latin [1036]. Lattice [663]. Lattices [106, 783, 758, 814, 1305, 868, 1206, 703, 249, 110]. Lava [1134]. Law [151]. Laws [1245]. Layers [1082]. Lazy [1375, 202]. LCS [315]. LDA [441]. LDYIS [1081]. Leader [392]. Learnability [976]. Learning [9, 782, 1319, 181, 86, 103, 621, 544, 879, 225, 1175, 548, 1098, 1048, 547, 332, 1014, 355, 183, 1322, 333, 1178, 384, 255, 694, 664, 761, 433, 520]. Least [21]. Leibniz [963]. LEM2 [1071]. LEM2-like [1071]. Lemmata [38, 497]. Length [439]. Letter [718]. Level [656, 1180, 552, 534, 747, 752, 405, 682, 1356, 187, 130, 457]. Levels [1281].

Leveraging [1178]. Libraries [120]. Library [175]. Life [500, 777]. Light [500]. Like [814, 327, 1071]. Likely [1094]. Limited [230, 946, 1087]. Limiting [1001]. Limits [753]. Linda [480]. Lindenmayer [87]. Line [1186, 705]. Linear [852, 1131, 1167, 38, 39, 1079, 760, 201, 601, 966, 968, 363, 1212, 1331, 1011]. Linearly [815]. Linguistic [867]. linked [1276]. Links [582]. Literals [832]. Living [168]. Load [1116]. Loading [884]. Local [750, 681, 932, 499, 1071, 915, 117, 1348, 356, 855, 582]. Localities [1160]. Locally [795]. Located [435]. Locations [852]. Löf [851]. Logarithm [305]. Logic [734, 996, 1039, 1371, 391, 197, 778, 1028, 655, 1041, 551, 144, 1061, 589, 657, 146, 101, 147, 999, 1272, 528, 28, 785, 336, 788, 890, 982, 39, 1199, 1267, 337, 954, 1019, 956, 438, 878, 998, 186, 278, 476, 687, 986, 627, 949, 692, 1366, 102, 995, 340, 960, 507, 961, 1178, 84, 1058, 1194, 21, 292, 642, 1297, 150, 268, 426, 151, 602, 27, 1055, 966, 1378, 449, 76, 895, 896, 970, 251, 992, 249]. Logic-Based [890, 1194, 992]. Logical [851, 474, 481, 482, 529, 1195, 1184, 508, 966, 247, 266]. Logics [342, 951, 1373, 145, 68, 248, 786, 1047, 789, 983, 244, 544, 635, 932, 1377, 1310, 41, 80, 148, 448, 149, 32, 1365, 641, 516, 222, 1370, 895, 896]. Logistic [760]. LoLA [281]. Long [1070, 672]. Long-running [1070]. Look [594]. Look-Up [594]. Lookup [320]. Looping [675]. Lossless [978, 1251]. Low [735, 462, 1357, 1105]. Low-complexity [735]. Low-cost [1105]. LR [559]. LSB [1361]. LTF [270]. LTF-C [270]. Łukasiewicz [972, 602, 222].

Machine [479, 502, 501, 520]. Machines [160, 835, 9, 1238, 1228, 903, 1229, 840, 759, 280, 763, 1231, 1298, 178, 841, 596, 839, 419, 1234, 1207, 445, 199]. Macro [1181]. Made [1184]. Magnetic [264]. Magnitude [589]. Maintain [382]. Making [259, 927, 1104]. Managed [1337]. Management [120, 880, 46]. Many [481, 482, 1159, 510]. Mapping [179, 1307, 775]. Mappings [517, 731]. Maps [386, 420]. Marcus [503, 510]. Marginalia [75]. Markings [678]. Markov [185, 1094, 879, 1340, 987]. Markovian [1139]. Markup [991]. Martin [851]. Martin-Löf [851]. MAS [687, 649]. Masking [359]. Match [470]. Matching [1030, 740, 1361, 311, 466, 705, 318, 116, 530, 1035, 671]. Matchmaking [1370]. Mateescu [715]. Mathematicae [2]. Mathematical [486, 954, 956, 3, 1209, 1217]. Mathematics [800, 1214]. Matrices [1046]. Matrix [179, 406, 1152, 517, 731, 1222]. Maurer [940]. max [526, 462]. MAX-BISECTION [462]. max-plus [526]. Maximal [1397, 1398]. Maximisation [1037]. Maximum [1303, 1035]. Mazurkiewicz [923, 934, 924, 932, 1083, 936]. MC [201]. McDermott [1365]. MCMAS [914]. MCMAS-X [914]. MDatalog [108]. Mealy [419]. Meaning [819, 407, 568]. Means [919, 1144]. Measure [1363, 1152]. Measurement [1034]. Measures [302, 336, 31, 430, 803, 65, 1090]. Measuring [1358, 1380]. Mechanism [958]. Mechanisms [164, 859, 581, 195, 631]. Mechanist [958]. Medical [388, 287]. Member [979]. Membrane [1119, 1120, 1280, 819, 491, 822, 167, 172, 159, 176, 13, 728, 537]. Membranes [352, 764, 161, 162, 327, 173, 706]. Memory [776, 304, 56, 1077, 1031]. Memory-Based [1077]. Mereology [792, 50, 96, 269]. Mereotopology [864, 98]. Merging [737, 1257]. Message [934, 493, 1388, 14, 195, 252, 679]. Messages [672]. Meta [629]. Meta-heuristic [629]. Metadata [120]. Metalanguage [18]. Metalinear [829]. Method [835, 1094, 436, 978, 1097, 559, 1099, 25, 183, 1369, 619, 1143, 217, 443, 833]. Methodology [1058, 457]. Methods [1060, 558, 467, 468, 384, 555, 1306, 1307, 49, 1144]. Metric [1216, 326]. Metric-Based [1216]. Metrics [863, 93]. Micro [1098]. Microbiological [675]. Microprocessor [1368]. Mild [726]. Miller [1004]. Minimal [1226, 58, 943, 11, 935, 1296, 623]. Minimax [1399]. Minimisation [1263]. Minimization [416]. Minimizing [217]. Minimum [639, 29]. Mining [1171, 1359, 633, 550, 1172, 1173, 544, 1151, 943, 408, 546, 878, 381, 1393, 884, 1221, 543, 545, 834]. Mirkin [79]. Missing [572]. Mobile [677, 219, 404, 681, 744, 279, 1333]. Modal [197, 951, 1028, 1272, 244, 635, 337, 1019, 343, 344, 814, 32, 1365, 960, 21, 108, 292, 892, 788]. Modal-Like [814]. Mode [824]. Model [1112, 423, 564, 1167, 740, 1383, 1345, 277, 1381, 113, 405, 1376, 1070, 744, 1340, 1332, 1268, 380, 168, 182, 530, 476, 688, 1078, 910, 410, 1295, 1322, 1081, 1298, 595, 1306, 212, 298, 1085, 916, 1348, 629, 841, 708, 1370, 698, 1270, 1275, 839, 473, 469, 897, 1154, 1406, 1208, 1211]. Model-Based [708]. Modeled [145]. Modeling [676, 791, 1066, 1216, 570, 211, 1103, 542, 1111, 484, 1404]. Modelled [913, 574]. Modelling [1279, 675, 1162, 1238, 15, 209, 296, 681, 824, 907, 33, 1330, 411, 573, 151, 828, 577, 1088, 593, 939]. Models [836, 36, 136, 1094, 625, 821, 1180, 107, 403, 879, 957, 628, 381, 1388, 21, 297, 431, 413, 132, 1274, 654, 53, 420, 22, 1223, 540, 623]. Modification [209]. Modified [594, 1153, 1144]. Modular [1343, 861]. Modularity [753]. Module [30]. Modulus [1099]. Moisil [972]. Molecular [564, 661]. Moment [736]. Moments [1383, 242, 1153]. Monadic [1267, 438, 27]. Monitoring [345, 346]. Monocultures [827]. Monoidal [391]. Monoids [727]. Monotonic [1019, 32]. Monotonicity [1319]. Morphologicae [2]. Morphological [9]. Morphology [3, 5]. Most [1094]. Motifs [379]. MP3 [1385]. mpleteness [451]. MPMA [1312]. MR [1152, 1392]. MRI [1144]. MSC [570, 911]. MSC-specified [570, 911]. MSO [1047]. Mt. [1134]. mu [438]. mu-Calculus [438]. Muller [427]. Multi [474, 1171, 1059, 656, 1192, 1345, 538, 1097, 1047, 788, 1173, 39, 405, 791, 1201, 1175, 365, 278, 187, 910, 410, 648, 1355, 45, 130, 298, 1392, 1348, 1403, 1349, 287, 473]. Multi-Agent [474, 1345, 278, 910, 45, 1348, 473, 648, 298]. Multi-bit-reference [538]. Multi-Criteria [1097]. Multi-Dimensional [1173, 1201, 365]. Multi-facet [1059]. Multi-Level [656, 405, 130]. Multi-modal [788]. Multi-objective [1355, 1403]. Multi-party [1192]. Multi-region [39]. Multi-Relational [1171, 1175]. Multi-Robot [791]. Multi-seed [1392]. Multi-swarm [1355]. Multi-Swarms [1349]. Multi-Tasking [187]. Multi-Valued [1047, 410].

Multiactions [1139]. Multiagent [789]. Multidimensional [1249, 883]. Multilateral [477, 576]. Multilayer [1054]. Multimedia [120]. Multimodal [589, 68, 1310]. Multipattern [312]. Multiple [136, 625, 1064, 736, 1398, 504, 1101, 612, 828, 701, 22]. Multiple-Clause [612]. Multiprocessor [196]. Multirepeats [1398]. Multiscale [25, 5]. Multisequential [139]. Multiset [780, 171, 1293, 1294]. Multiset-Based [780]. Multisplitting [114]. Multistage [107]. Multistrategy [621]. Multithreaded [1000, 196]. Multivalued [1262]. Muscle [1383]. Music [886]. Musical [121]. Mutation [1352]. Mutually [1345]. MV [972, 615]. MV-Algebras [972, 615]. my [1302].

n [323]. Name [681, 1161]. Narrowing [133]. Nash [343]. Natural [496, 705, 956, 149, 1218]. Near [799, 1358]. Nearness [916, 799, 921, 1155]. Necessity [31]. Necklaces [1010]. Negative [642]. Negligibility [589]. Negotiation [576, 761, 581]. Negotiations [397, 477, 576]. Neighborhoods [874]. Nested [1374, 351, 45, 130, 211, 630]. Nesting [855]. Net [925, 1162, 1328, 1167, 274, 1285, 634, 534, 39, 127, 404, 567, 1065, 41, 279, 1160, 911, 1333, 297, 194, 1161, 133, 653, 207, 1343]. Net-based [1328]. Nets [1326, 1281, 37, 1166, 1342, 1329, 927, 1157, 62, 678, 1062, 125, 275, 566, 901, 1063, 1327, 552, 902, 209, 296, 276, 405, 567, 681, 682, 406, 1159, 17, 1076, 1138, 1108, 187, 933, 1079, 1292, 690, 909, 1387, 571, 1158, 1110, 45, 130, 211, 1389, 1390, 131, 412, 695, 132, 416, 574, 573, 51, 652, 919, 1088, 1301, 443, 218, 702, 660, 906, 1072, 1163, 1290, 294, 762]. Network [1247, 180, 88, 685, 1330, 1034, 1178, 1213]. Networks [1225, 1121, 29, 107, 86, 103, 1242, 355, 913, 900, 188, 237, 573, 542, 1308]. Neural [1121, 784, 86, 103, 1008, 662, 1123, 1213, 270, 1395, 1126]. Neuro [1132]. Neuro-Genetic [1132]. Neuronmuscular [1383]. Newton [963]. NFA [490]. Nice [714]. NIL [28]. NN [619]. No [710]. Nodes [1225]. Noise [1356]. Non [205, 273, 1018, 1011, 1019, 1321, 32, 855, 798, 299, 707, 215, 155, 576, 156, 1087, 240, 1233, 443, 252, 123, 249, 330]. Non-Clausal [156]. non-commutative [249]. Non-Deterministic [707, 155, 330]. Non-dominated [1018]. Non-extensive [1087]. Non-functional [215]. non-Instantaneous [123]. Non-Interference [205, 273]. Non-Isolated [576]. Non-linear [1011]. Non-local [855]. Non-Monotonic [1019, 32]. Non-Numerical [1321]. Non-Optimum-Time [1233]. Non-Reachability [299, 443]. Non-Regular [240]. Non-Symbolic [798]. non-uniform [252]. Nonassociative [307]. Nondeterminism [971, 230, 1150, 366]. Nondeterministic [1237, 28, 785, 1203, 561, 12, 730]. Nonmodularity [89]. Nonmonotonic [374, 142]. Nonparametric [1015]. Nonstandard [957, 250]. Nonterminal [1006]. Nonunique [392]. norm [391]. Normal [1374, 333, 159, 84, 141, 1265]. Normalisation [76]. Normalization [1041, 854]. Normalized [1305, 1102, 65]. Normalized-scale [1305]. Normative [152]. Notation [202]. Note [887, 1202, 426, 104, 733]. Notion [205, 273, 247, 329, 357]. Novel [658, 843, 640, 1049, 669, 1205]. NP [189].

Number [229, 811, 233, 369, 217]. Number-Conserving [233, 229]. Numbers [952, 769, 956, 898]. Numerical [930, 1321, 728]. Numerous [1116]. Nuovo [1190].

Object [258, 563, 902, 127, 209, 296, 404, 405, 681, 682, 1067, 1292, 279, 909, 1220, 1354, 202, 358, 671]. Object-based [358]. Object-Net [404]. Object-Oriented [671]. objective [1355, 1403]. Objects [55, 331, 904, 211, 174, 184, 514, 916, 799, 154, 580, 141, 617]. Obligations [1055]. Observable [1386]. Observation [767, 905]. Observational [674, 420]. Observations [717]. Observers [1168]. Occurrence [311, 41, 1387, 1152, 919]. Occurrences [715]. Octahedra [857]. off [1274]. Offline [1250]. Often [1174]. OI [113]. OI-implication [113]. Omega [417, 575, 697]. Omega-product [697]. Omega-Products [417, 575]. On-line [705]. Once [1174]. One [1191, 1236, 369, 227, 510, 246, 773, 876, 537, 231, 1245]. One-Bit [876]. One-Clock [1191]. One-dimensional [1236, 231]. One-Turn [227]. One-variable [773]. One-Visit [246]. One-Way [369]. Online [1250]. Online/Offline [1250]. Ontologically [1380]. Ontologies [338, 1286, 992]. Ontology [1288, 69, 213]. OO [1082]. Open [400, 453, 592, 597, 420, 523]. Opening [4]. Operational [398, 15]. Operations [1006, 1009, 1203, 375, 618, 428, 1026, 1336]. Operator [1352, 986, 1312]. Operators [85, 353, 657, 972, 621, 814, 531, 463, 965, 1090]. Optimal [718, 181, 114, 193, 316, 464, 1133, 601, 894, 1038, 1308]. Optimality [806]. Optimisation [815]. Optimistic [890]. Optimization [1353, 590, 193, 1074, 1354, 844, 430, 693, 1357, 1349, 283]. Optimizations [563]. Optimizer [1382]. Optimum [1129, 1233]. Oracles [141]. Order [223, 1060, 551, 589, 924, 276, 871, 1019, 607, 438, 1152, 1042, 48, 638, 433, 733, 420, 587, 565, 1267, 32, 692, 1312]. Ordered [316, 642]. Orders [262, 703]. Ordinal [1319, 112]. Organisations [910]. Organized [145]. Orientation [242]. Oriented [258, 1338, 46, 671, 1221]. Other [486]. Otsu [1253]. our [69]. ’Out-Gödel [958]. Outcomes [761]. Outfix [974]. Outfix-Free [974]. Outline [1288]. Output [1169, 13, 178]. Output-Determinacy [1169]. Overview [550, 92, 487, 509, 653]. OverWords [1047]. Ownership [669].

PAC [9]. package [314]. Packet [107]. Page [816]. Pair [1205]. Pairing [69]. Pairs [851]. Pairwise [1321]. Paper [1197, 1302, 1058]. Papers [847, 746, 524]. Paraconsistent [668, 1404]. Paradigm [643, 269]. Paradigms [1214]. Paradiso [1190]. Parallel [302, 295, 614, 1303, 403, 1286, 870, 1268, 1140, 1256, 47, 1185, 1348]. Parallelism [1226]. Parallelization [1306, 1307, 8]. Parameter [1094, 193]. Parameterizations [531]. Parameterized [552]. Parameters [1060, 1282]. Parametric [1239, 857, 31]. Parasitic [1023]. Pareto [858]. Parikh [179, 1046, 517, 731]. Parity [1385]. Parity-based [1385]. Parsable [942].

Parser [917]. Parsimoniously [1227]. Parsing [559, 499, 917, 1086, 1300]. Part [1389, 1390, 592, 481, 482, 467, 468, 1306, 1307, 556, 605, 1270, 1275, 597, 895, 896, 1394, 1405]. Partial [1226, 1166, 1342, 1095, 79, 584, 796, 1269, 1314, 48, 420, 483]. Partially [1386, 1140, 184, 920]. Particle [1353, 1354, 1357, 815]. Particles [1349, 362]. Partition [180, 979, 886]. Partitioning [300, 1307]. Partitions [1226]. Parts [131]. party [1192]. Passing [934, 14, 679]. Password [604, 1145]. Past [1163]. Path [981, 1182, 908, 47, 1308]. PathLog [184]. Patient [52]. Pattern [1030, 1132, 676, 779, 942, 318, 1273, 368, 873, 834]. Patterns [633, 646, 418, 577, 1205]. Pawlak [776, 778, 1028, 793, 794, 777, 800]. Payload [1148]. PCAM [1307]. PDL [78, 322]. Peak [886]. Pearson [1212]. Pebble [232, 1181]. Perceptron [1054]. Perceptual [1358]. Perfect [633]. Performance [988, 1341, 414, 196]. perICS [1092]. Perimeters [617]. Periodic [1010, 363]. Periodicity [488]. Permitting [981]. Permutation [1036]. Permutations [891]. Perseus [1283]. Persistency [1281]. Personal [581]. Persuasion [1061, 1283]. Petri [1326, 842, 294, 37, 1166, 1342, 1329, 927, 1162, 1167, 1157, 274, 207, 552, 634, 534, 39, 127, 209, 296, 276, 405, 567, 682, 1065, 406, 1159, 17, 1076, 1138, 1108, 187, 933, 911, 571, 1110, 45, 130, 211, 1139, 1343, 1389, 1390, 131, 297, 132, 416, 574, 573, 1161, 133, 653, 652, 919, 817, 1301, 443, 702, 1072, 1163, 1290]. Petrinets [299, 215]. Phase [978, 1322]. Pheromone [1273]. Philosophers [937]. Philosophy [274, 949]. Phones [693]. Photosynthesis [176]. Phylogenies [863]. Physical [238]. Pick [648]. Pick-up [648]. Picture [1237, 327, 794, 330]. Piecewise [439]. Piecewise-Quartics [439]. Pieri [955]. Pipeline [193]. Pixel [632, 1254, 1022, 1113, 1056, 1205]. Pixel-Value [1022]. Pixel-Values [632]. Pixels [1361]. Place [1281, 1157, 62, 125, 207]. Place/Transition [1281, 62, 125, 207]. Places [1159]. Plan [890]. Planar [1186]. Plane [424, 324, 1313]. Planning [345, 1068]. Platek [61, 60]. Play [475]. Plotkin [1271]. plus [526]. PMA [58]. Point [557, 531, 1369, 98, 144]. Point-Free [98, 557]. Points [858, 814, 98, 1038]. Polarizationless [1124]. Polarized [1041]. Policies [258, 807, 1327, 1194]. Polygonal [1248]. Polymorphic [206, 671]. Polynomial [1037, 455, 257]. Polynomial-Time [1037, 455, 257]. polynomials [259, 354]. Polyphonic [1400]. POMDP [1048]. Pomykała [945]. Pose [1320]. Poset [663]. Posets [262]. Positions [152]. Positive [21]. Possessing [1116]. Possibilistic [336, 102, 898]. Possibility [787]. Possibility-Theoretic [787]. Possible [683, 825, 182, 155]. Potential [962]. Power [460, 721, 169, 1244, 174]. Practical [1030, 323, 946, 1333]. Pragmatic [792]. Pragmatics [1039, 1055]. PRAM [304]. Pre [1123]. Pre-computed [1123]. Prebases [79]. Precursor [794]. Predator [1023]. Predator-Prey [1023]. Predicate [1088, 895, 896]. Predicate/Transition [1088]. Predicate/Transition-Nets [1088]. Predication [1040]. Prediction [754, 1180, 866, 1137, 112]. Predictive [865]. Preemptible [187]. Preface [471, 645, 818, 948, 1156, 1224, 1235, 35, 122, 271, 293, 396, 348, 204, 143, 335, 91, 111, 70, 1, 190, 452, 429, 447, 377, 1209, 1217]. Preference [539]. Preferences [147, 347]. Preferential [1377, 1182, 1034, 34]. Prefix [718, 1203]. Prefix-Free [718, 1203]. Prefixes [1343]. Pregroups [425]. Prelanguages [109]. Preprocessing [1060, 619]. Preprocessor [1282]. Prescribed [1014]. Presentability [894]. Presentation [1394, 1405]. Preservation [1247]. Preserving [925, 736, 885]. Pretopologies [385]. Preventing [1259]. Prevention [1387, 154]. Previous [966]. Prey [1023]. Pricing [810]. Prime [952, 974]. Primes [372]. Primitive [917]. Principal [72, 528]. Principle [1302]. Principles [953, 1302, 647, 309]. Prioritized [737, 1257]. Priority [443]. Privacy [880, 885]. Probabilistic [564, 339, 1069, 260, 548, 506, 1366, 1141, 729]. Probabilistically [533, 553]. Probability [666, 889, 1114, 80, 344, 878, 571, 1366]. Probably [1284]. Problem [352, 1227, 821, 479, 751, 126, 1267, 720, 722, 233, 530, 909, 648, 320, 305, 975, 1382, 601, 518, 1025, 1104, 1233, 389, 1125, 251, 261]. Problems [486, 902, 715, 1240, 1048, 59, 1355, 5, 153, 1357, 1031, 239, 876, 105, 1265, 671, 523, 257, 314]. Procedure [489, 642, 151]. Procedures [248, 268, 155]. Process [993, 435, 1283, 713, 38, 125, 879, 88, 1289, 1138, 186, 1341, 250, 761, 1270, 1275, 990, 938, 1331]. Process-algebraic [938]. Process-Specific [761]. Processes [185, 373, 274, 534, 404, 567, 1268, 860, 626, 933, 138, 987, 414, 415, 427, 697, 1087, 733, 593, 702]. Processing [400, 614, 182, 3, 798]. Processor [304, 811]. Processors [1225, 47, 1185]. Product [307, 64, 697]. Production [15]. Products [206, 982, 727, 417, 575]. Professor [776]. Profile [879]. Program [610, 666, 609, 557, 559, 608, 611, 556, 605, 1245]. Program-Size [666]. Programmable [1306]. Programming [734, 673, 899, 1057, 1278, 9, 565, 588, 562, 657, 538, 1375, 403, 680, 825, 347, 1058, 642, 202, 968, 592, 597, 839, 895, 478, 1331]. Programs [402, 1374, 930, 403, 1286, 558, 770, 855, 561, 995, 961, 84, 21, 292, 707, 1299, 1378, 14, 1092]. Progress [1171]. Projection [1083]. Promela [1084]. Proof [1000, 106, 1004, 849, 374, 558, 891, 20, 642, 894, 1155, 1271]. Proof-Carrying [849]. Proofs [854, 850]. propagating [707]. Properties [10, 677, 925, 926, 1167, 930, 704, 717, 973, 1241, 1013, 723, 1079, 690, 1311, 171, 1294, 1110, 961, 915, 48, 298, 215, 1016, 522, 1349, 733, 1265, 301, 906, 469]. Property [809, 888, 595, 729]. Propositional [891, 968]. Propositions [997]. Propositions-as-Types [997]. Protection [669]. Protein [881]. Proteins [706]. Protocol [1162, 688, 582, 604, 914]. Protocols [209, 446, 907, 1291, 913, 1295, 1081]. Prototyping [414]. Prove [299, 443]. Prover [1297]. Proving [489, 275, 566, 1347, 592, 597]. Provision [1337]. Proximal [1005]. Proximity [755, 756, 1155]. Proxy [323, 253, 622]. Pruning [657, 1356]. PSO [1356]. PSO-SVM [1356]. PSPACE [352, 28, 932]. PSPACE-Complete [352, 28]. Psychologism [970]. Publication [1328]. Pure [820, 245]. Purse [837]. Push [81]. Pushdown [971, 554, 1293, 1294, 227]. Putnam [888]. Puzzle [1227].

QoS [810, 1339, 1337]. QoS-Driven [1339]. Qualitative [97, 589, 92, 94, 50, 542]. Quandary [837]. Quantales [1266]. Quantification [891]. Quantifier [692]. Quantifier-free [692]. Quantifying [1289]. Quantitative [1339]. Quantization [658, 808, 1385, 866, 1137, 669, 1255, 470]. Quantization-Based [1255]. Quantum [1236, 1241, 691, 1349, 1088, 1125]. Quantum-Inspired [1125]. Quartics [439]. Quasi [454, 336, 1038, 526]. Quasi-indicators [1038]. quasi-interpretations [526]. Quasi-Possibilistic [336]. Quasi-static [454]. Queries [331, 1182, 157]. Query [614, 461, 108, 184]. Querying [1244]. Queues [1245]. Quotient [389, 1308].

R. [542]. Races [401]. Radial [1130, 873]. RAM [841]. Random [941, 1362, 981, 1261, 1118, 760, 603, 1089]. Randomized [1219]. Randomness [1003, 223]. Range [933]. Ranges [221]. Rank [853, 496, 1097, 460, 1018, 975, 772]. Ranking [614, 1321, 397, 477]. Rational [474, 38, 1096, 769, 590, 720, 985, 831, 730]. Rationalism [100]. Rationality [539, 970]. RBFFCA [873]. RCC [540]. RCC8 [308]. Reachability [126, 909, 300, 299, 239, 762, 443, 218, 301, 422, 583, 922]. Reaction [790, 241]. Reaction-Diffusion [241]. Reactive [453, 484]. read [1336]. Real [1248, 765, 219, 717, 448, 1078, 1341, 536, 1056, 240, 372]. Real- [1341]. Real-Coded [1056]. Real-Time [1078, 240, 372]. Realizability [1042, 1044]. Reals [1017]. Reason [213]. Reasoning [996, 1061, 124, 589, 147, 999, 92, 786, 406, 94, 1195, 867, 355, 84, 50, 386, 156, 580, 1404]. Rebeca [484]. receptivity [252]. Recognition [1132, 599, 232, 942, 1177, 1020, 19, 693, 708, 441, 886, 236]. Recognizability [1240]. Recognizable [1237, 1346, 831]. Recognizer [1124]. Recognizing [352, 710]. Reconfiguration [999, 42]. Reconstruction [1320, 654]. Record [52]. Recording [207]. Records [529]. Rectangle [1129]. Rectangular [1186]. Recurrent [973]. Recursion [61, 1043, 587]. Recursive [162, 765, 853, 1386, 130, 333, 536, 917, 894, 1245]. Recursive-Descent [917]. Recursively [498, 511]. Reduced [1116, 183]. Reduces [1227]. Reducible [900]. Reducing [51, 359]. Reduction [106, 1137, 59, 1109, 227, 434, 598, 1093]. Reductions [48, 105, 1265]. Reducts [796, 265, 434, 135]. Redundant [606]. reference [538]. Refined [436]. Refinement [836, 113, 1159, 501, 457]. Refinements [837, 243]. Refining [55]. Regexpcount [314]. Region [951, 99, 755, 756, 100, 623, 39]. Region-Based [99, 951, 755, 756, 100]. Regions [93, 95]. Register [424]. Registers [732]. Regression [112]. Regular [144, 1196, 1363, 497, 981, 1008, 1182, 719, 813, 974, 1203, 616, 462, 1256, 19, 1103, 1016, 104, 240, 317, 288, 314]. regularities [1032]. Regulated [227]. Regulatory [1247, 542]. Rehashing [669]. Reinforcement [1098, 664]. Reiter [1045]. Related [231, 303, 481, 482, 983, 733]. Relating [373]. Relation [757, 1287, 1051, 69, 738, 1027]. Relation-Algebraic [1027]. Relational [1171, 639, 331, 621, 1173, 547, 585, 944, 1177, 1305, 918, 1175].

Relations [589, 1172, 533, 553, 393, 904, 31, 572, 1305, 20, 26, 729, 516, 155, 730, 282]. Relationship [1383, 504, 992]. Relative [785, 1137]. Relativizing [58]. Relaxed [180]. Relevance [161]. Reliability [451]. Reliable [29]. Remarks [871, 415]. Remote [1054, 1056]. Reorganization [193]. Replaceability [1328]. Replacement [747]. Replicated [160]. Replication [1161]. Repository [120]. Representability [69]. Representation [734, 1003, 92, 663, 406, 1036, 1083, 153, 26, 1367, 152]. Representations [85, 1342, 498]. Representative [384]. Representing [994]. Requirements [1257]. Research [954, 428, 923]. Resemblances [1358]. Residual [224]. Residuals [1212]. Residuated [620]. Resistant [446]. Resolution [668, 1109, 292, 642, 832]. Resolving [119]. Resonance [264]. Resource [810, 660, 294]. Resource-Constrained [660]. Resources [590, 1123, 1380, 1087]. Respect [1003, 302]. Respiratory [1068]. Restarted [1176]. Restarting [503]. Restoration [1207]. Restricted [352, 492, 983]. Restriction [979, 579]. Restriction-Based [579]. Restrictive [626]. Result [687]. Resulted [1382]. Results [1258, 490, 854, 505, 89, 256, 22, 523, 90]. Retaining [555]. Retracts [797]. Retrenching [837]. Retrieval [118, 617, 833]. Retrieving [66]. Revealed [539]. Reversal [1009]. Reversible [754, 771, 988, 1254, 259]. Revision [463, 259]. Revisited [1165, 821, 757, 731]. Revisiting [274]. Rewrite [610, 1350]. Rewriting [160, 767, 680, 998, 505, 105, 1265, 483, 518]. RFCM [947]. Rialto [625]. Rich [593]. Rings [392]. RIONA [225]. Risk [357, 376]. Risk-type [376]. Rlationship [438]. Robinson [969]. Robot [791, 578]. Robots [219, 828]. Robust [394, 1264, 1356, 708, 665, 1145]. Roles [145]. ROSECON [413]. Rotation [1227]. Rough [1028, 655, 378, 676, 779, 1316, 782, 783, 1317, 1198, 620, 1260, 465, 68, 757, 244, 210, 683, 1288, 1118, 380, 254, 944, 1311, 1323, 618, 1053, 947, 262, 535, 1391, 384, 255, 12, 387, 297, 49, 132, 50, 96, 263, 269, 426, 415, 516, 650, 801, 816, 222, 282, 519, 264, 1325, 651, 700, 433, 918, 578, 579, 701, 341, 1404, 287, 1406, 390]. Route [83]. Routes [727]. RRIA [390]. RST [1067]. RTCP [762]. RTCP-nets [762]. RTPA [1215]. Rule [882, 1226, 991, 1319, 680, 225, 1053, 285, 935, 612, 1378, 264, 651, 700, 883, 142, 390]. Rule-Based [680, 700]. RuleML [994]. Rules [764, 991, 322, 993, 610, 888, 116, 169, 1322, 796, 1269, 1314, 1296, 399, 432, 426, 214, 433, 1105, 229]. running [1070]. Runtimes [574].

Safe [927, 567, 1079]. Safety [247]. Sandpiles [812]. SAT [1242, 688, 456, 627, 172, 1085, 301, 422, 583, 922, 1091]. SAT-Based [922, 1085, 422, 583]. Satellite [1113]. Satisfaction [42]. Satisfiability [1227, 932, 568, 1067]. Saturation [1332]. SBL [1202, 1187]. Scalable [320]. Scale [1359, 6, 1305]. Scale-free [1359]. Scaling [1249, 883]. Scan [1186]. Scattered [82, 288]. Scenario [453, 1330]. Scenario-Based [453]. Scenarios [1291]. Scenery [445]. Schedulability [1110]. Schedules [574].

Scheduling [810, 454, 647, 1355, 131]. Schemaless [184]. Scheme [538, 458, 658, 632, 843, 1147, 1197, 1190, 1250, 736, 640, 865, 323, 1049, 253, 945, 1388, 600, 669, 1021, 946, 1255, 622, 775, 665, 1154, 1276, 1145]. Schemes [633, 714, 305, 643, 1245, 1205]. Science [800, 428, 967]. Scoped [610]. Scouts [18]. SE [104]. SE-Systems [104]. Search [1352, 590, 973, 544, 984, 636, 312, 643, 630]. Searching [808, 809, 659, 846, 881]. Second [1267, 871, 438, 1152, 1042, 587]. Second-Order [1042, 1267]. Secondary [881]. Secret [830, 1362, 714]. Secretly [1116]. SecSpaces [741]. Secure [402, 1190, 277, 741, 1330, 604, 1154, 672]. Security [905, 1069, 1070, 1289, 547, 907, 1291, 913, 1295, 945, 1081]. seed [1392]. Segment [358]. Segmentation [1152, 1367, 1392, 1144, 445, 1220]. Selected [847, 1188, 746, 524]. Selection [1064, 1339, 117]. Selective [86, 103]. Selector [504]. Selectors [139]. Self [780, 149, 199, 1306, 1307, 671, 644]. Self-Adaptive [1306, 1307]. Self-Assembly [780]. Self-describing [199]. Self-dual [149]. Self-Stabilization [644]. Selfish [985]. Semantic [991, 993, 819, 849, 995, 1055, 1380, 1370, 883]. Semantics [882, 1329, 402, 1373, 740, 62, 848, 634, 243, 584, 585, 182, 1138, 1108, 1215, 995, 535, 641, 1391, 961, 398, 292, 1043, 1103, 1270, 1275, 733, 895, 896, 480, 15, 461]. Semi [711, 1323, 1054, 1369, 1234, 582]. Semi-GAPS [1369]. Semi-local [582]. Semi-Simple [711]. Semi-supervised [1323, 1054, 1369]. Semi-Weakly [1234]. Semigroup [887]. Semilattices [3]. Semilinear [912]. Semirings [1095, 912, 428]. Semistructured [115]. Sense [1252, 520]. Sensing [1054, 1056]. Sensitive [1183]. Sensitivity [86, 103, 726]. Sensor [132]. Sentences [148]. Separation [950, 682]. seq [608]. Sequence [1173, 116, 240]. Sequences [675, 1360, 1094, 379, 661, 1398, 802, 881, 157, 543]. Sequent [868]. Sequential [302, 824, 1268, 721, 138, 415, 834, 286]. Sequentiality [532]. Series [720, 1256, 1015, 261]. Series-Parallel [1256]. Service [926, 1193, 1192, 1338, 446, 1339, 742]. Services [1328, 1370]. Session [745]. Set [378, 676, 779, 1372, 1316, 1198, 1375, 465, 244, 460, 1288, 1332, 347, 618, 995, 1391, 384, 255, 395, 387, 297, 49, 263, 269, 415, 516, 650, 1358, 1325, 433, 578, 579, 701, 189, 839, 217, 283, 287, 1406, 390, 816]. Set-based [1372]. Set-Theoretic [395]. Sets [639, 766, 782, 303, 1317, 953, 1302, 1260, 68, 568, 1230, 1118, 148, 380, 31, 1311, 831, 912, 947, 464, 935, 513, 799, 200, 415, 650, 801, 282, 519, 264, 651, 700, 918, 53, 341, 1404, 7, 189]. Setting [333, 450]. Seventh [1334]. Shadowed [303]. Shape [93]. Shapely [351]. Shapes [1135]. Shapley [263]. Shared [743, 832]. Sharing [830, 1116, 1362, 56, 714]. Shi [253]. Shift [1142]. Shifts [1013]. shop [1355]. Short [1256, 520, 1208]. Shortest [722, 47, 83]. Shuffle [725, 774, 977]. Shuffles [1246]. Side [470]. Side-Match [470]. Signal [599, 455, 1107, 40, 1098, 17, 1109, 51, 862, 195]. Signal-Event [40, 51]. Signals [1383]. Signature [1250, 323, 253, 305, 622, 1154, 1208]. Signatures [1043]. Signed [1003]. Signed-Digit [1003]. Signers [323, 622]. Silent [1258]. SIMD [1368].

Similar [1147, 316]. Similarities [883]. Similarity [353, 124, 1363, 208, 496, 786, 1199, 16, 615, 904, 1391, 643, 516, 249]. Similarity-Based [786, 1391, 16]. Simple [106, 1004, 711, 1122, 242, 1261, 684, 770, 724, 320, 47, 743, 596, 330]. Simplification [667]. Simulate [771]. Simulating [175, 178]. Simulation [1012, 1291, 1034, 841, 1088]. Simulations [365, 237, 176, 364]. Simulator [163]. Simultaneous [51]. Since [966]. Single [940, 1320]. Single-Thread [940]. Singular [542]. Situated [613, 473]. Six [1316]. Sixth [1164]. Size [666, 1230]. Skeletal [1383]. Skeptical [84]. Skewed [1175]. Skolem [1229]. SLD [292]. SLD-Resolution [292]. Slicing [1189, 409, 686]. Small [1231, 1233, 1234]. Smaller [1126]. Smith [1368]. Smoothness [717]. SMVQ [754]. Social [1242, 576]. Sofic [1051]. Soft [1381]. Software [1188, 739, 164, 163, 1257, 651, 745, 1283]. Solid [704, 1100]. Solution [1025, 1233]. Solutions [472, 876]. Solve [1048, 153]. Solving [352, 821, 479, 1123, 1357, 389]. Some [1248, 1046, 1006, 715, 983, 276, 210, 148, 957, 415, 1086, 1204, 1104, 876, 733, 22, 142, 482, 247]. Sort [684]. Sorting [1121, 1018, 1200]. Sound [678, 121, 886]. Sources [1024]. Space [951, 191, 1167, 755, 756, 1174, 318, 1071, 183, 366, 321, 916, 100, 576, 828, 1038, 389, 1308, 370]. Space- [366]. Space-Efficient [321]. Spaces [1028, 681, 407, 568, 683, 904, 61, 60, 367, 1142, 385, 664, 133, 699, 326, 921, 1090, 1155, 356]. Spatial [632, 92, 928, 50, 833]. Spatial-Histogram [833]. Spatio [1201, 580]. Spatio-Temporal [1201, 580]. Spatiotemporal [381]. sPBC [1139]. SPCF [855]. Special [856, 877, 1164, 1106, 1119, 1120, 1334, 1396, 746, 624, 201, 799]. Specific [544, 1151, 761]. Specification [673, 899, 1057, 1278, 562, 502, 45, 1082, 592, 597]. Specifications [272, 453, 30, 1257, 64, 861, 14]. Specified [846, 570, 911]. Spectral [1344, 875]. Spectrum [1245]. SpecVer [1058]. Speech [599, 693, 708]. Speed [1151]. Speeding [315]. Speeding-up [315]. SPICE [878]. Spiking [1121, 784, 1008, 662, 1123, 1395, 1126]. Splicing [711, 537]. Split [1011]. SQEMA [1272]. Squad [233, 775, 1233]. Stability [1384, 234, 1143]. Stabilization [644]. Stable [674, 961, 53]. Stage [193, 990, 287, 1113]. Standard [1162, 1154, 1208]. Standpoint [1067]. Star [1009, 569, 689, 696]. Star-Connected [569, 689]. Star-free [696]. Start [823]. State [1112, 835, 1167, 224, 929, 1107, 1009, 840, 1340, 1203, 501, 1077, 1033, 456, 987, 1298, 299, 133, 839, 1143, 775, 876, 1233, 644, 523, 841]. State-Efficient [775, 876]. State-Estimation [1077]. Stateless [1195]. Statements [650]. States [1094, 542, 217, 593]. Static [1168, 304, 928, 154, 1170, 454]. Statistical [710, 1212]. Steel [1133]. Steganographic [458]. Steganography [979, 1385, 1022]. Steiner [1038]. Step [1327, 1385, 194, 603]. STG [456, 627, 1109]. Stochastic [1345, 1096, 646, 749, 1341, 1348, 817, 236]. Stochastic-Time [1341]. Stop [823]. Stopwatches [1389, 1390]. Storm [216]. Straight [1186]. Straight-Line [1186]. Strategies [656, 1147, 647, 8]. Strategy [538, 811, 1176]. Stream [160, 502]. Streams [1151]. Strictly [704]. String [767, 784, 1046, 646, 466, 705, 1031]. String-Rewriting [767]. Stringology [1396]. Strings [735, 226, 1032]. Strong [1041, 854, 1181, 687, 44, 76, 604, 1124]. Strong-Password [604]. Structural [677, 925, 1062, 40, 1110, 398, 192]. Structure [97, 661, 254, 411, 874, 881, 1104, 388, 286]. Structured [678, 549, 636, 355, 1387, 548]. Structures [525, 674, 1282, 303, 481, 482, 620, 928, 275, 1285, 714, 752, 39, 209, 591, 1128, 167, 955, 670, 738, 34, 1170, 593, 921]. Stubborn [53, 217, 283]. Study [210, 371, 222, 651, 598, 289]. Stuttering [986]. Style [272, 888]. Subcontexts [1206]. Subgraphs [546]. Submonoids [772, 109]. Subregular [494]. Subset [176, 1123]. Subsets [1364]. Subshifts [231, 1051, 234]. Substitution [538]. Substitutions [1272, 723]. Substring [311]. Subsumption [565, 1176]. Subtree [550]. Subtrees [548]. Subtypes [206]. Subword [715, 802]. Successor [217]. Succinctness [27]. Suffix [661, 321]. Sum [893, 1123]. Summarizing [331]. Sunspot [694]. Sup [85]. Sup-Compact [85]. Super [690, 535]. Super-Dual [690]. Superclasses [1080]. superpixels [1220]. Superstring [722]. supervised [1323, 1054, 1369]. Support [1207, 445]. Supporting [136, 613, 741]. Surface [704]. Surfer [603]. Surjective [851]. Survey [332, 520]. SVM [1356]. Swarm [1351, 1353, 1354, 1357, 1358, 1355]. Swarms [815, 1349]. Switching [751, 350]. Symbol [806]. Symbolic [926, 930, 768, 798, 314]. Symbols [820]. Symmetric [854, 812]. Symmetry [1056, 1369, 1204, 1393]. Symport [514]. Symport/Antiport [514]. Synchronicity [1108]. Synchronisation [405]. Synchronization [563, 491, 233, 775, 1129, 1233]. Synchronized [725, 64, 1111, 1246]. Synchronous [552, 451, 861, 330]. Synergy [14]. Syntactic [1037, 942, 883]. Syntax [1218]. Synthesis [526, 1166, 1342, 453, 207, 1063, 1285, 1327, 859, 627, 1169, 1160, 987, 297, 96, 419, 192]. Synthesising [194]. System [1000, 856, 1164, 1106, 1334, 322, 15, 146, 163, 166, 225, 905, 1202, 1386, 891, 624, 1387, 911, 20, 46, 176, 868, 413, 202, 828, 579, 653, 654, 1187, 457, 358, 617, 1337]. System-level [457]. Systems [160, 474, 352, 764, 400, 818, 996, 161, 197, 778, 1132, 564, 675, 1280, 1059, 819, 820, 925, 488, 781, 1238, 453, 821, 1345, 657, 481, 482, 491, 767, 327, 1121, 784, 712, 1122, 857, 57, 713, 822, 980, 277, 1064, 768, 590, 790, 747, 404, 87, 1240, 40, 165, 870, 824, 1008, 137, 94, 374, 825, 1287, 647, 1098, 1048, 749, 380, 721, 662, 409, 793, 375, 823, 827, 187, 279, 33, 1160, 169, 505, 170, 570, 691, 1366, 771, 1123, 1324, 45, 130, 211, 411, 508, 509, 1341, 159, 174, 900]. Systems [1333, 285, 291, 935, 1296, 175, 514, 728, 706, 298, 177, 194, 50, 200, 214, 178, 1161, 894, 155, 1348, 484, 732, 997, 577, 1087, 1232, 829, 701, 920, 104, 330, 1350, 537, 241, 473, 593, 939, 1394, 1405, 644, 1124, 1395, 1126, 142, 196, 356, 1125]. Szymanski [315].

t [323, 1377]. Table [538, 978, 308, 1212]. Tableau [1297]. Tableaux [342, 1310]. Tables [594, 541, 431, 413, 65, 654]. Tabular [585]. TADDs [1299]. Tagging [1177]. Tantrix^{T M} [1227]. Targets [1381]. Tasking [187]. Taverna [1270, 1275]. TBase2 [52]. TCP [926]. TCTL [897]. Team [1246]. Teams [474, 57]. Teamwork [328]. Technique [594, 1148, 1255, 707, 1056, 1392, 359, 360, 192]. Techniques [1252, 688, 1387, 1178, 556, 605]. Technology [1315]. Telemetric [1201]. Teleportation [1088]. Temperature [444]. Temporal [10, 677, 982, 983, 932, 1201, 186, 476, 986, 966, 580]. Tense [972]. Tensor [1393]. Term [734, 1342, 1041, 1350, 1265, 483]. Terms [634, 793]. TESLA [914]. Testers [1241]. Testing [36, 502, 501, 933, 1176, 1215, 1347, 14, 885]. Text [321]. Texts [466]. Textual [136]. Texture [1344, 408, 1153]. Their [621, 585, 723, 1305, 227, 240, 1209, 1217, 1131, 1063, 404, 247, 14, 195]. Theorem [1146, 1004, 766, 893, 75, 592, 597, 958]. Theorems [608]. Theorems-Based [608]. Theoretic [465, 248, 787, 1013, 395, 1043, 1267]. Theoretical [486, 289, 556, 605, 1223]. Theories [951, 1248, 825, 333, 1178, 100, 156, 90]. Theory [709, 1127, 1040, 24, 1192, 952, 1260, 755, 756, 748, 752, 113, 244, 437, 1288, 368, 1391, 595, 868, 799, 269, 650, 803, 152, 1325, 592, 597, 189, 450, 1222, 287, 1266, 1027, 389, 483]. There [560]. Thermodynamics [1087]. Thesis [962]. Thomas [542]. Thoughtful [759]. Thread [656, 940]. Three [978, 786, 1254, 514, 602, 241]. Three-Circular-Pixel [1254]. Three-dimensional [241]. Three-Phase [978]. Three-Valued [786, 602]. Threshold [830, 323, 1253, 622]. Thresholding [803]. Thue [64]. Tightly [997]. Tiles [1114]. Tiling [1237, 1240]. Time [1167, 1157, 1037, 219, 491, 1062, 455, 318, 840, 349, 1340, 448, 1078, 183, 366, 509, 1341, 451, 150, 48, 416, 574, 573, 201, 1015, 698, 817, 256, 450, 240, 372, 1129, 1233, 1301, 257, 1389, 1390]. Time-Bounded [1340, 366]. Time-Optimum [1129]. Time-Varying [509, 256]. Time/Space [318]. Timed [1191, 36, 674, 123, 205, 273, 1309, 1165, 1167, 1258, 459, 208, 1136, 1287, 1289, 907, 409, 686, 908, 1073, 81, 1295, 43, 129, 1366, 628, 989, 1084, 1085, 300, 299, 203, 449, 652, 420, 301, 897, 422, 583, 922, 480]. Times [941]. Timestamps [131]. Timing [435, 250]. Tissue [764, 327, 824]. Tissue-like [327]. Token [842, 1329, 274, 1337, 1072]. Token-Managed [1337]. Tokens [1079, 1158]. Tolerance [1356, 1358]. Tolerant [807]. Tool [94, 414]. Tools [1384, 370]. Topological [10, 486, 167, 415, 1016]. Topologies [1038]. Total [492]. Trace [1076, 696, 695]. Trace-Expressible [1076]. Traces [924, 932, 1083, 936, 575]. Tracing [533, 553]. Tracking [1354]. Trade [1274]. Trade-off [1274]. Training [270]. Trajectories [709, 716, 1135, 513]. Trajectory [393, 439]. Trajectory-Based [393]. Transaction [884]. Transactions [1070, 1290]. Transducer [77]. Transducers [1181]. Transductions [929]. Transfer [926]. Transform [1402, 8, 1325, 1153]. Transformation [610, 557, 93, 747, 748, 349, 749, 351, 611, 1177, 556, 605, 707, 586]. Transformation-based [1177]. Transformations [609, 608, 129, 561].

Transforming [490, 682, 176]. Transforms [5]. Transition [1281, 1157, 62, 125, 207, 1285, 455, 1107, 1287, 628, 177, 194, 862]. Transition-Nets [1088]. Transitions [927, 1258]. Transitive [1185]. Translation [493, 1073, 911, 1084, 1091]. Translator [1299]. Transport [1280, 648, 159]. Transversal [943]. Traversal [633]. Treatment [1068]. Tree [77, 1172, 1117, 981, 982, 983, 1346, 1181, 549, 548, 636, 661, 375, 307, 183, 246, 1103, 1186, 1105, 390, 288]. Tree-based [1105]. Tree-Structured [636, 548]. Trees [198, 73, 1047, 973, 1267, 1399, 1151, 1263, 438, 316, 112, 11, 12, 19, 188, 285, 430, 774, 977, 804, 1325, 4, 543, 364, 545]. Triangle [808, 350]. Triangle-Free [350]. Triangular [915]. Truly [62]. Truncation [736, 640]. Truth [655]. TSPN [1112]. Tuning [479]. Tupling [606]. Turing [1228, 903, 199, 1231, 178, 1234]. Turn [227]. Tutorial [849]. Twenty [726]. Two [1225, 1236, 808, 929, 1009, 825, 59, 503, 1322, 1113, 990, 775]. Two-Bounds [808]. Two-Dimensional [775]. Two-Phase [1322]. Two-Stage [990, 1113]. Two-Way [503, 929]. Type [1040, 24, 72, 1025, 592, 597, 921, 257, 376, 141]. Type-1 [141]. Type-2 [141]. Typed [847, 529, 748, 524, 1043]. Typeful [609]. Types [1225, 1040, 1131, 1149, 464, 154, 592, 597, 997, 745, 671, 587, 527]. Typing [528, 848, 745].

Ultraproducts [306]. UMFFs [1397]. UML [1298]. Unambiguous [1237]. Unary [1237, 1363, 872]. Unavoidable [958]. Unbounded [476, 1085]. Uncertain [648]. Uncertainty [994]. Undecidabilities [1184]. Undecidability [1258, 719, 278, 989]. Undecidable [1240, 1313]. Understanding [450, 309]. Unequal [718]. Unfolding [1329, 126, 561]. Unfoldings [128, 260, 456, 627, 1109, 1343, 1333]. Unification [295, 782, 181, 16, 464, 976]. Uniform [932, 1034, 82, 513, 252]. Unifying [882, 928, 269]. Unique [1397, 251]. Unit [764]. Units [1268]. Universal [1236, 166, 367, 1243, 1231, 514, 48, 212, 202, 1015, 1234, 582, 1126]. Universality [1191, 1228, 232, 768, 763, 159, 989, 773]. Unordered [545]. Unorthodox [322]. Unranked [1263]. Unsoundness [958]. Unsupervised [1252, 1113, 1367, 1144]. Until [986]. Untraceable [305]. Untyped [851]. Update [180, 58, 463, 1312, 698]. Updating [1185]. Upper [618]. UPSILON [202]. Use [1193, 1174, 17, 408, 884]. Used [216]. Useful [1167]. User [1180, 119, 879, 1145]. User-Level [1180]. User/Process [879]. Uses [957]. Using [1344, 136, 538, 1317, 1361, 594, 632, 808, 857, 1046, 1135, 1286, 681, 406, 193, 640, 865, 1137, 25, 456, 118, 112, 947, 873, 1339, 1333, 18, 131, 153, 300, 299, 612, 1392, 1369, 133, 651, 579, 1403, 1143, 1088, 582, 360, 445, 1249, 811, 345, 984, 1332, 1098, 467, 468, 1297, 1103, 803, 484, 883, 745, 1331]. Utilities [581].

Vague [650, 519]. Valences [87]. Validating [998]. Validation [637, 1330]. Value [1022, 263, 707]. Value-propagating [707]. Valued [481, 482, 786, 1047, 1296, 222, 1406, 410, 426, 602]. Values [632, 410, 572]. vanishing [362]. Variable [770, 773]. Variables [56, 760, 729, 1378]. Variances [898]. Variant [969]. Variants [137, 247, 376]. Varieties [983]. Various [1053, 430, 257]. Varying [509, 256]. Vector [658, 808, 866, 1137, 691, 669, 442, 326, 470, 1207, 445]. VerICS [1078]. Verification [857, 208, 626, 628, 45, 914, 1333, 1082, 215, 281, 484, 449, 264, 938, 1092]. Verified [1190]. Verifying [677, 913, 1347, 298]. Versus [302, 173, 109, 822]. Vertex [639]. Vertexlet [1359]. Very [596]. via [1114, 606, 275, 1010, 1220, 235, 298, 96, 301]. Video [1103, 1220]. View [378, 787, 1130, 758]. Viewpoint [598]. Views [704]. Viral [722]. Visit [246]. Visual [1259, 1362, 358]. Visualization [883]. Voices [732]. Volcano [1134]. Volumetric [216]. VQ [436, 754, 843, 809, 978, 865, 1253, 541, 1143]. VQ-Based [436]. vs [1161].

w.r.t [966]. Wajsberg [481]. Walks [941]. Wang [1114]. Watchman [83]. Waterman [1368]. Watermarking [1264, 1205, 1276]. Watermarks [360, 470]. Watershed [8]. Wave [973]. Wavelet [121]. Way [503, 369, 929]. Ways [207]. Weak [1262, 438]. Weaker [221, 969]. Weakly [846, 1234, 356]. Web [991, 993, 1193, 1328, 1180, 995, 816, 52, 1370]. Weight [581]. Weight-Based [581]. Weighted [1359, 717, 1263, 661, 1020, 1152]. Weights [1314]. Well [141, 461]. Well-Behaved [141]. well-founded [461]. Wet [1197]. Whether [720, 288]. Which [460, 1158]. Widespread [482]. Windows [1301]. Winning [554]. without [979, 986]. Word [79, 1020, 19, 1382, 520]. Words [710, 1007, 802, 774, 977, 522, 1211, 317, 314]. Work [777, 1044]. Workbench [1270, 1275]. Workflow [1062, 660, 906]. Working [824, 1163]. Workshop [1029]. World [791, 1066]. Worm [174]. Write [1336]. Write-after-read [1336].

X [914]. XCCS [1279]. XOR [1242]. XOR-SAT [1242].

Years [726]. Yu [1004].

Z. [794]. Zdzisław [776, 777, 800]. Zernike [1153]. Zero [567, 1079, 1245, 359]. Zero-one [1245]. Zero-Safe [567, 1079].

References

Goutsias:2000:P

[1] John Goutsias and Henk Heijmans. Preface. Fundamenta Informaticae, 41(1–2):??, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Goutsias:2000:FMM

[2] John Goutsias and Henk J. A. M. Heijmans. Fundamenta morphologicae mathematicae. Fundamenta Informaticae, 41(1–2):1–31, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keshet:2000:MMC

[3] Renato Keshet (Kresch). Mathematical morphology on complete semilattices and its applications to image processing. Fundamenta Informaticae, 41(1–2):33–56, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vincent:2000:GOT

[4] Luc Vincent. Granulometries and opening trees. Fundamenta Informaticae, 41(1–2):57–90, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maragos:2000:CED

[5] Petros Maragos and Muhammad Akmal Butt. Curve evolution, differential morphology, and distance transforms applied to multiscale and eikonal problems. Fundamenta Informaticae, 41(1–2):91–129, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Soille:2000:BGS

[6] Pierre Soille. From binary to grey scale convex hulls. Fundamenta Informaticae, 41(1–2):131–146, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Serra:2000:CSF

[7] Jean Serra. Connections for sets and functions. Fundamenta Informaticae, 41(1–2):147–186, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Roerdink:2000:WTD

[8] Jos B. T. M. Roerdink and Arnold Meijster. The watershed transform: Definitions, algorithms and parallelization strategies. Fundamenta Informaticae, 41(1–2):187–228, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barrera:2000:APM

[9] J. Barrera, R. Terada, R. Hirata, Jr., and N. S. T. Hirata. Automatic programming of morphological machines by PAC learning. Fundamenta Informaticae, 41(1–2):229–258, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baier:2000:THT

[10] Christel Baier and Marta Kwiatkowska. On topological hierarchies of temporal properties. Fundamenta Informaticae, 41(3):259–294, February 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2000:ACD

[11] Mikhail Moshkov and Igor Chikalov. On algorithm for constructing of decision trees with minimal depth. Fundamenta Informaticae, 41(3): 295–299, February 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2000:DND

[12] Mikhail Moshkov. Deterministic and nondeterministic decision trees for rough computing. Fundamenta Informaticae, 41(3):301–311, February 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2000:MCE

[13] Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Membrane computing with external output. Fundamenta Informaticae, 41(3):313– 340, February 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsiatsoulis:2000:TDM

[14] Z. Tsiatsoulis and J. Y. Cotronis. Testing and debugging message passing programs in synergy with their specifications. Fundamenta Informaticae, 41(3):341–366, February 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bos:2000:FPS

[15] V. Bos and J. J. T. Kleijn. Formalisation of a production system modelling language: The operational semantics of χ core. Fundamenta Informaticae, 41(4):367–392, March 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Formato:2000:SBU

[16] Ferrante Formato, Giangiacomo Gerla, and Maria I. Sessa. Similaritybased unification. Fundamenta Informaticae, 41(4):393–414, March 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hanisch:2000:SEP

[17] H.-M. Hanisch and A. Lüder. A signal extension for Petri nets and its use in controller design. Fundamenta Informaticae, 41(4):415–431, March 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalski:2000:BKS

[18] Ryszard S. Michalski and Kenneth A. Kaufman. Building knowledge scouts using KGL metalanguage. Fundamenta Informaticae, 41(4):433– 447, March 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2000:DTR

[19] Mikhail Moshkov. Decision trees for regular language word recognition. Fundamenta Informaticae, 41(4):449–461, March 2000. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

MacCaull:2000:PSD

[20] Wendy MacCaull. A proof system for dependencies for information relations. Fundamenta Informaticae, 42(1):1–27, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2000:CLM

[21] Linh Anh Nguyen. Constructing the least models for positive modal logic programs. Fundamenta Informaticae, 42(1):29–60, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2000:MDC

[22] Yongge Wang, Yvo Desmedt, and Mike Burmester. Models for dependable computation with multiple inputs and some hardness results. Fundamenta Informaticae, 42(1):61–73, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2000:ADF

[23] Józef Winkowski. An algebra of data flows. Fundamenta Informaticae, 42 (1):75–104, January 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boldini:2000:FCI

[24] Pascal Boldini. Formalizing context in intuitionistic type theory. Fundamenta Informaticae, 42(2):105–127, May 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Inkielman:2000:MHD

[25] Marcin Inkielman. A method for hierarchical description of images using multiscale analysis. Fundamenta Informaticae, 42(2):129–148, May 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paz:2000:RIR

[26] Azaria Paz, Robert Y. Geva, and Milan Studený. Representation of irrelevance relations by annotated graphs. Fundamenta Informaticae, 42 (2):149–199, May 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rabinovich:2000:SGB

[27] A. Rabinovich. Succinctness gap between monadic logic and duration calculus. Fundamenta Informaticae, 42(2):201–210, May 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demri:2000:NIL

[28] Stéphane Demri. The nondeterministic information logic NIL is PSPACE-complete. Fundamenta Informaticae, 42(3–4):211–234, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Denejko:2000:RMF

[29] Piotr Denejko, Krzysztof Diks, Andrzej Pelc, and Marek Piotrów. Reliable minimum finding comparator networks. Fundamenta Informaticae, 42(3–4):235–249, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engel:2000:TAC

[30] Marcin Engel, Artur Kret, and Janina Mincer-Daszkiewicz. Towards automated consistency checking of module interface specifications. Fundamenta Informaticae, 42(3–4):251–278, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Inuiguchi:2000:NMP

[31] Masahiro Inuiguchi and Tetsuzo Tanino. Necessity measures and parametric inclusion relations of fuzzy sets. Fundamenta Informaticae, 42 (3–4):279–302, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:2000:FON

[32] Michael Kaminski and Guy Rey. First-order non-monotonic modal logics. Fundamenta Informaticae, 42(3–4):303–333, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohn:2000:FAS

[33] Markus Kohn and Hartmut Schmeck. Formal asynchronous systems modelling. Fundamenta Informaticae, 42(3–4):335–389, July 2000. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schlechta:2000:NAP

[34] Karl Schlechta, Laurent Gourmelen, Stéphanie Motré, Olivier Rolland, and Bensalah Tahar. A new approach to preferential structures. Fundamenta Informaticae, 42(3–4):391–410, July 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2000:P

[35] Hans-Dieter Burkhard, Peter Starke, and Ludwik Czaja. Preface. Fundamenta Informaticae, 43(1–4):??, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andreeva:2000:ATC

[36] Maria V. Andreeva, Elena N. Bozhenkova, and Irina B. Virbitskaite. Analysis of timed concurrent models based on testing equivalence. Fundamenta Informaticae, 43(1–4):1–20, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bednarczyk:2000:FCC

[37] Marek A. Bednarczyk, Andrzej M. Borzyszkowski, and Rafał Somla. Finite completeness of categories of Petri nets. Fundamenta Informaticae, 43(1–4):21–48, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2000:RLA

[38] Ludwik Czaja and Manfred Kudlek. Rational, linear and algebraic process languages and iteration lemmata. Fundamenta Informaticae, 43 (1–4):49–60, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2000:MRL

[39] Berndt Farwer. A multi-region linear logic based calculus for dynamic Petri net structures. Fundamenta Informaticae, 43(1–4):61–79, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Foremniak:2000:SAS

[40] Adrianna Foremniak and Peter H. Starke. Structural analysis of signalevent systems. Fundamenta Informaticae, 43(1–4):81–104, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haar:2000:ONL

[41] Stefan Haar. Occurrence net logics. Fundamenta Informaticae, 43(1– 4):105–127, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hannebauer:2000:FAD

[42] Markus Hannebauer. A formalization of autonomous dynamic reconfiguration in distributed constraint satisfaction. Fundamenta Informaticae, 43(1–4):129–151, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lanotte:2000:TCA

[43] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Adriano Peron. Timed cooperating automata. Fundamenta Informaticae, 43(1–4):153–173, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Latvala:2000:CSF

[44] Timo Latvala and Keijo Heljanko. Coping with strong fairness. Fundamenta Informaticae, 43(1–4):175–193, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2000:NPN

[45] Irina A. Lomazova. Nested Petri nets — a formalism for specification and verification of multi-agent distributed systems. Fundamenta Informaticae, 43(1–4):195–214, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Munch:2000:CCA

[46] Ines Münch and Gabriela Lindemann-v.Trzebiatowski. ChariTime — concepts of analysis and design of an agent-oriented system for appointment management. Fundamenta Informaticae, 43(1–4):215–226, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nepomniaschaya:2000:SID

[47] Anna S. Nepomniaschaya and Maria A. Dvoskina. A simple implementation of Dijkstra’s shortest path algorithm on associative parallel processors. Fundamenta Informaticae, 43(1–4):227–243, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2000:IPO

[48] Wojciech Penczek, Maciej Szreter, Rob Gerth, and Ruurd Kuiper. Improving partial order reductions for universal branching time properties. Fundamenta Informaticae, 43(1–4):245–267, August 2000. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2000:ARS

[49] James F. Peters, Andrzej Skowron, and Zbigniew Suraj. An application of rough set methods in control design. Fundamenta Informaticae, 43(1– 4):269–290, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2000:RMI

[50] Lech Polkowski and Andrzej Skowron. Rough mereology in information systems with applications to qualitative spatial reasoning. Fundamenta Informaticae, 43(1–4):291–320, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Roch:2000:ARS

[51] Stephan Roch. Analyzing and reducing simultaneous firing in signalevent nets. Fundamenta Informaticae, 43(1–4):321–330, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schroter:2000:TWB

[52] Kay Schröter, Gabriela Lindemann-v.Trzebiatowski, and Lutz Fritsche. TBase2 — a Web–based electronic patient record. Fundamenta Informaticae, 43(1–4):343–353, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Varpaaniemi:2000:SMS

[53] Kimmo Varpaaniemi. Stable models for stubborn sets. Fundamenta Informaticae, 43(1–4):355–375, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aho:2000:IK

[54] Isto Aho. Interactive knapsacks. Fundamenta Informaticae, 44(1–2):1– 23, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buchi:2000:FRC

[55] Martin Büchi and Emil Sekerinski. A foundation for refining concurrent objects. Fundamenta Informaticae, 44(1–2):25–61, August 2000. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chrobot:2000:SVD

[56] Stanisław Chrobot. Sharing variables in distributed memory. Fundamenta Informaticae, 44(1–2):63–81, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2000:DTE

[57] Erzsébet Csuhaj-Varjú and Victor Mitrana. Dynamical teams in ecogrammar systems. Fundamenta Informaticae, 44(1–2):83–94, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:2000:PRM

[58] Patrick Doherty, Witold Łukaszewicz, and Ewa Madalińska-Bugaj. The PMA and relativizing minimal change for action update. Fundamenta Informaticae, 44(1–2):95–131, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Intrigila:2000:TPR

[59] Benedetto Intrigila and Anna R. Laurenzi. Two problems on reduction graphs in Lambda calculus. Fundamenta Informaticae, 44(1–2):133–144, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ivanov:2000:PS

[60] Lyubomir Ivanov. Platek spaces. Fundamenta Informaticae, 44(1–2): 145–181, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ivanov:2000:BRP

[61] Lyubomir Ivanov. Boldface recursion on Platek spaces. Fundamenta Informaticae, 44(1–2):183–208, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Busi:2000:CTC

[62] Nadia Busi and G. Michele Pinna. Comparing truly concurrent semantics for contextual place/transition nets. Fundamenta Informaticae, 44(3): 209–244, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martinek:2000:CCF

[63] Pavel Martinek. On a construction of context-free grammars. Fundamenta Informaticae, 44(3):245–264, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Payet:2000:TSI

[64] Étienne Payet. Thue specifications, infinite graphs and synchronized product. Fundamenta Informaticae, 44(3):265–290, August 2000. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slezak:2000:NDF

[65] Dominik Ślezak. Normalized decision functions and measures for inconsistent decision tables analysis. Fundamenta Informaticae, 44(3):291– 319, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Analyti:2000:DRC

[66] Anastasia Analyti, Nicolas Spyratos, and Panos Constantopoulos. Deriving and retrieving contextual categorical information through instance inheritance. Fundamenta Informaticae, 44(4):321–351, December 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2000:CC

[67] Jürgen Dassow, Carlos Martín-Vide, Gheorghe Păun, and Alfonso Rodríguez-Patón. Conditional concatenation. Fundamenta Informaticae, 44(4):353–372, December 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demri:2000:CCM

[68] Stéphane Demri and Jarosław Stepaniuk. Computational complexity of multimodal logics based on rough sets. Fundamenta Informaticae, 44(4):373–396, December 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurucz:2000:RPR

[69] Agnes Kurucz and István Németi. Representability of pairing relation algebras representability of pairing relation algebras depends on our ontology. Fundamenta Informaticae, 44(4):397–420, December 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Girard:2001:P

[70] Jean-Yves Girard and Paweł Urzyczyn. Preface. Fundamenta Informaticae, 45(1–2):??, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baillot:2001:ECG

[71] Patrick Baillot and Marco Pedicini. Elementary complexity and geometry of interaction. Fundamenta Informaticae, 45(1–2):1–31, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Broda:2001:CTP

[72] Sabine Broda and Luís Damas. Counting a type’s (principal) inhabitants. Fundamenta Informaticae, 45(1–2):33–51, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

David:2001:CBT

[73] René David. Computing with Böhm trees. Fundamenta Informaticae, 45 (1–2):53–77, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sato:2001:EE

[74] Masahiko Sato, Takafumi Sakurai, and Rod Burstall. Explicit environments. Fundamenta Informaticae, 45(1–2):79–115, January 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Statman:2001:MTJ

[75] Rick Statman. Marginalia to a theorem of Jacopini. Fundamenta Informaticae, 45(1–2):117–121, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urban:2001:SNC

[76] C. Urban and G. M. Bierman. Strong normalisation of cut-elimination in classical logic. Fundamenta Informaticae, 45(1–2):123–155, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apro:2001:GTT

[77] János Apró and Sándor Vágvölgyi. The ground tree transducer game with identical tree automata. Fundamenta Informaticae, 45(3):157–172, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2001:IFP

[78] Philippe Balbiani and Dimiter Vakarelov. Iteration-free PDL with intersection: a complete axiomatization. Fundamenta Informaticae, 45(3): 173–194, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2001:MPA

[79] Jean-Marc Champarnaud and Djelloul Ziadi. From Mirkin’s prebases to Antimirov’s word partial derivatives. Fundamenta Informaticae, 45(3): 195–205, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hajek:2001:CFP

[80] Petr Hájek and Sauro Tulipani. Complexity of fuzzy probability logics. Fundamenta Informaticae, 45(3):207–213, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krishnan:2001:DTP

[81] Padmanabhan Krishnan. Decomposing timed push down automata. Fundamenta Informaticae, 45(3):215–229, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meduna:2001:UGL

[82] Alexander Meduna. Uniform generation of languages by scattered context grammars. Fundamenta Informaticae, 45(3):231–235, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nilsson:2001:ASW

[83] Bengt J. Nilsson. Approximating a shortest Watchman route. Fundamenta Informaticae, 45(3):235–281, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Milnikel:2001:SRF

[84] Robert Saxon Milnikel. Skeptical reasoning in FC-normal logic programs is π_{1}^{1}-complete. Fundamenta Informaticae, 45(3):237–252, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barrera:2001:SCI

[85] Junior Barrera and Ronaldo Fumio Hasimoto. Sup-compact and infcompact representations of W -operators. Fundamenta Informaticae, 45 (4):283–294, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engelbrecht:2001:SASa

[86] Andries P. Engelbrecht. Sensitivity analysis for selective learning by feedforward neural networks. Fundamenta Informaticae, 45(4):295–328, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2001:VLS

[87] Henning Fernau and Ralf Stiebe. Valences in Lindenmayer systems. Fundamenta Informaticae, 45(4):329–358, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2001:PAN

[88] Damas P. Gruska and Andrea Maggiolo-Schettini. Process algebras for network communication. Fundamenta Informaticae, 45(4):359–378, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salibra:2001:NRL

[89] Antonino Salibra. Nonmodularity results for Lambda calculus. Fundamenta Informaticae, 45(4):379–392, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhao:2001:CRD

[90] Xishun Zhao and Decheng Ding. Complexity results for 2CNF default theories. Fundamenta Informaticae, 45(4):393–404, January 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2001:P

[91] Ivo Düntsch. Preface. Fundamenta Informaticae, 46(1–2):??, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cohn:2001:QSR

[92] A. G. Cohn and S. M. Hazarika. Qualitative spatial representation and reasoning: An overview. Fundamenta Informaticae, 46(1–2):1–29, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Davis:2001:CST

[93] Ernest Davis. Continuous shape transformation and metrics on regions. Fundamenta Informaticae, 46(1–2):31–54, January 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Galton:2001:DDT

[94] Antony Galton. Dominance diagrams: A tool for qualitative reasoning about continuous systems. Fundamenta Informaticae, 46(1–2):55–70, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2001:AAR

[95] Ivo Düntsch, Ewa Orłowska, and Hui Wang. Algebras of approximating regions. Fundamenta Informaticae, 46(1–2):71–82, January 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2001:CSR

[96] Lech Polkowski. On connection synthesis via rough mereology. Fundamenta Informaticae, 46(1–2):83–96, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bittner:2001:QSB

[97] Thomas Bittner. The qualitative structure of built environments. Fundamenta Informaticae, 46(1–2):97–128, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schoop:2001:PPF

[98] Dominik J. Schoop. Points in point-free mereotopology. Fundamenta Informaticae, 46(1–2):129–143, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bennett:2001:CAR

[99] Brandon Bennett. A categorical axiomatisation of region-based geometry. Fundamenta Informaticae, 46(1–2):145–158, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pratt-Hartmann:2001:ERR

[100] Ian Pratt-Hartmann. Empiricism and rationalism in region-based theories of space. Fundamenta Informaticae, 46(1–2):159–186, January 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Celani:2001:FID

[101] Sergio A. Celani. A fragment of intuitionistic dynamic logic. Fundamenta Informaticae, 46(3):187–197, August 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liau:2001:PDL

[102] Churn-Jung Liau and Duen-Ren Liu. A possibilistic decision logic with applications. Fundamenta Informaticae, 46(3):199–217, August 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engelbrecht:2001:SASb

[103] Andries P. Engelbrecht. Sensitivity analysis for selective learning by feedforward neural networks. Fundamenta Informaticae, 46(3):219–252, August 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tiplea:2001:NSS

[104] Ferucio Laurenţiu Ţiplea and Erkki Mäkinen. A note on SE-systems and regular canonical systems. Fundamenta Informaticae, 46(3):253–256, August 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Verma:2001:ARR

[105] Rakesh M. Verma, Michael Rusinowitch, and Denis Lugiez. Algorithms and reductions for rewriting problems. Fundamenta Informaticae, 46 (3):257–276, August 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Belohlavek:2001:RSP

[106] Radim Bělohlávek. Reduction and a simple proof of characterization of fuzzy concept lattices. Fundamenta Informaticae, 46(4):277–285, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dietzfelbinger:2001:DMP

[107] Martin Dietzfelbinger, Anna Gambin, and Sławomir Lasota. On different models for packet flow in multistage interconnection networks. Fundamenta Informaticae, 46(4):287–314, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2001:MQL

[108] Linh Anh Nguyen. The modal query language MDatalog. Fundamenta Informaticae, 46(4):315–342, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ostravsky:2001:PVS

[109] Jan Ostravský. Prelanguages versus submonoids. Fundamenta Informaticae, 46(4):343–348, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ronse:2001:GCL

[110] Christian Ronse and Jean Serra. Geodesy and connectivity in lattices. Fundamenta Informaticae, 46(4):349–395, December 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

El-Kwae:2001:P

[111] E. El-Kwae and Z. Raś. Preface. Fundamenta Informaticae, 47(1–2):??, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kramer:2001:POC

[112] Stefan Kramer, Gerhard Widmer, Bernhard Pfahringer, and Michael De Groeve. Prediction of ordinal classes using regression trees. Fundamenta Informaticae, 47(1–2):1–13, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esposito:2001:GMB

[113] Floriana Esposito, Nicola Fanizzi, Stefano Ferilli, and Giovanni Semeraro. A generalization model based on OI-implication for ideal theory refinement. Fundamenta Informaticae, 47(1–2):15–33, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Elomaa:2001:CCO

[114] Tapio Elomaa and Juho Rousu. On the computational complexity of optimal multisplitting. Fundamenta Informaticae, 47(1–2):35–52, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hacid:2001:CBA

[115] Mohand-Saïd Hacid, Farouk Toumani, and Ahmed K. Elmagarmid. Constraint-based approach to semistructured data. Fundamenta Informaticae, 47(1–2):53–73, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hyun:2001:DME

[116] Park Sang Hyun and Wesley W. Chu. Discovering and matching elastic rules from sequence databases. Fundamenta Informaticae, 47(1–2):75–90, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Puuronen:2001:LFS

[117] Seppo Puuronen and Alexey Tsymbal. Local feature selection with dynamic integration of classifiers. Fundamenta Informaticae, 47(1–2):91– 117, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kim:2001:CBR

[118] Minkoo Kim, Jitender S. Deogun, and Vijay V. Raghavan. Concept based retrieval using generalized retrieval functions. Fundamenta Informaticae, 47(1–2):119–135, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernandes:2001:DRU

[119] Chris Fernandes and Lawrence Henschen. Discovering and resolving user intent in heterogeneous databases. Fundamenta Informaticae, 47(1–2): 137–154, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertino:2001:MII

[120] Elisa Bertino, Barbara Catania, and Gian Piero Zarri. Metadata, intelligent indexing, and repository management for multimedia digital libraries. Fundamenta Informaticae, 47(1–2):155–173, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wieczorkowska:2001:MSC

[121] Alicja Wieczorkowska. Musical sound classification based on wavelet analysis. Fundamenta Informaticae, 47(1–2):175–188, September 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2001:P

[122] Hans-Dieter Burkhard, Peter Starke, and Ludwik Czaja. Preface. Fundamenta Informaticae, 47(3–4):??, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2001:TAN

[123] Roberto Barbuti, Nicoletta De Francesco, and Luca Tesei. Timed automata with non-instantaneous actions. Fundamenta Informaticae, 47(3– 4):189–200, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2001:SDC

[124] Hans-Dieter Burkhard. Similarity and distance in case based reasoning. Fundamenta Informaticae, 47(3–4):201–215, October 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2001:PLP

[125] Ludwik Czaja and Manfred Kudlek. ω-process languages for place/ transition nets. Fundamenta Informaticae, 47(3–4):217–229, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:2001:UBA

[126] Javier Esparza and Claus Schröter. Unfolding based algorithms for the reachability problem. Fundamenta Informaticae, 47(3–4):231–245, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2001:CCO

[127] Bernd Farwer. Comparing concepts of object Petri net formalisms. Fundamenta Informaticae, 47(3–4):247–258, October 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haar:2001:CCU

[128] Stefan Haar. Clusters, confusion and unfoldings. Fundamenta Informaticae, 47(3–4):259–270, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lanotte:2001:TTC

[129] Ruggero Lanotte, Andrea Maggiolo-Schettini, Simone Tini, and Adriano Peron. Transformations of timed cooperating automata. Fundamenta Informaticae, 47(3–4):271–282, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2001:NPN

[130] Irina A. Lomazova. Nested Petri nets: Multi-level and recursive systems. Fundamenta Informaticae, 47(3–4):283–293, October 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Neuendorf:2001:DSP

[131] Klaus-Peter Neuendorf, Dong-Ho Lee, Dimitris Kiritsis, and Paul Xirouchakis. Disassembly scheduling with parts commonality using Petri nets with timestamps. Fundamenta Informaticae, 47(3–4):295–306, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2001:SFF

[132] James F. Peters, S. Ramanna, M. Borkowski, Andrzej Skowron, and Zbigniew Suraj. Sensor, filter, and fusion models with rough Petri nets. Fundamenta Informaticae, 47(3–4):307–323, October 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schmidt:2001:NPN

[133] Karsten Schmidt. Narrowing Petri net state spaces using the state equation. Fundamenta Informaticae, 47(3–4):325–335, October 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2001:IGD

[134] Andrzej Skowron and Jarosław Stepaniuk. Information granule decomposition. Fundamenta Informaticae, 47(3–4):337–350, October 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wroblewski:2001:ECB

[135] Jakub Wróblewski. Ensembles of classifiers based on approximate reducts. Fundamenta Informaticae, 47(3–4):351–360, October 2001. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bach:2001:STI

[136] Joscha Bach. Supporting textual input by using multiple entropy models. Fundamenta Informaticae, 48(1):1–7, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frisco:2001:VCD

[137] Pierluigi Frisco and Claudio Zandron. On variants of communicating distributed H systems. Fundamenta Informaticae, 48(1):9–20, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:2001:BAC

[138] Maciej Koutny and Giuseppe Pappalardo. Behaviour abstraction for communicating sequential processes. Fundamenta Informaticae, 48(1): 21–54, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meduna:2001:MGH

[139] Aleksander Meduna and Petr Vurm. Multisequential grammars with homogeneous selectors. Fundamenta Informaticae, 48(1):55–60, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2001:EHC

[140] Hung Son Nguyen. On efficient handling of continuous attributes in large data bases. Fundamenta Informaticae, 48(1):61–81, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tourlakis:2001:CTO

[141] George Tourlakis. Computability in Type-2 objects with well-behaved Type-1 oracles is p-normal. Fundamenta Informaticae, 48(1):83–91, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhao:2001:SAE

[142] Xishun Zhao and Decheng Ding. Some algorithms for extension computation of nonmonotonic rule systems. Fundamenta Informaticae, 48 (1):93–106, October 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demolombe:2001:P

[143] Robert Demolombe and Risto Hilpinen. Preface. Fundamenta Informaticae, 48(2–3):??, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Broersen:2001:FPC

[144] Jan Broersen, Roel Wieringa, and John-Jules Meyer. A fixed-point characterization of a Deontic logic of regular action. Fundamenta Informaticae, 48(2–3):107–128, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carmo:2001:DAL

[145] José Carmo and Olga Pacheco. Deontic and action logics for organized collective agency, modeled through institutionalized agents and roles. Fundamenta Informaticae, 48(2–3):129–163, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carmo:2001:ADL

[146] Jose Carmo, Robert Demolombe, and Andrew J. I. Jones. An application of Deontic logic to information system constraints. Fundamenta Informaticae, 48(2–3):165–181, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cholvy:2001:AAL

[147] Laurence Cholvy and Christophe Garion. An attempt to adapt a logic of conditional preferences for reasoning with contrary-to-duties. Fundamenta Informaticae, 48(2–3):183–204, November 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hansen:2001:SSS

[148] Jörg Hansen. Sets, sentences, and some logics about imperatives. Fundamenta Informaticae, 48(2–3):205–226, November 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jennings:2001:NFS

[149] Ray E. Jennings. Natural frames and self-dual logics. Fundamenta Informaticae, 48(2–3):227–236, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Parent:2001:CIT

[150] Xavier Parent. Cumulativity, identity and time in Deontic logic. Fundamenta Informaticae, 48(2–3):237–252, November 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Prakken:2001:MDL

[151] Henry Prakken. Modelling defeasibility in law: Logic or procedure? Fundamenta Informaticae, 48(2–3):253–271, November 2001. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sergot:2001:RAA

[152] Marek Sergot and Fiona Richards. On the representation of action and agency in the theory of normative positions. Fundamenta Informaticae, 48(2–3):273–293, November 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2001:UDF

[153] Ngoc Thanh Nguyen. Using distance functions to solve representation choice problems. Fundamenta Informaticae, 48(4):295–314, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Puntigam:2001:TAO

[154] Franz Puntigam and Christof Peter. Types for active objects with static deadlock prevention. Fundamenta Informaticae, 48(4):315–341, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sakai:2001:EPH

[155] Hiroshi Sakai. Effective procedures for handling possible equivalence relations in non-deterministic information systems. Fundamenta Informaticae, 48(4):343–362, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stachniak:2001:NCR

[156] Zbigniew Stachniak. Non-clausal reasoning with definite theories. Fundamenta Informaticae, 48(4):363–388, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tyszkiewicz:2001:CSQ

[157] Jerzy Tyszkiewicz. Computability by sequences of queries. Fundamenta Informaticae, 48(4):389–414, December 2001. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin-Vide:2002:F

[158] Carlos Martín-Vide and Gheorghe Păun. Foreword. Fundamenta Informaticae, 49(1–3):??, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin-Vide:2002:MSC

[159] Carlos Martín-Vide, Andrei Păun, Gheorghe Păun, and Grzegorz Rozenberg. Membrane systems with coupled transport: Universality and normal forms. Fundamenta Informaticae, 49(1–3):1–15, January 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aguado:2002:SRR

[160] Joaquin Aguado, Tudor Balanescu, Tony Cowling, Marian Gheorghe, Mike Holcombe, and Florentin Ipate. P systems with replicated rewriting and stream X-machines (Eilenberg machines). Fundamenta Informaticae, 49(1–3):17–33, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ardelean:2002:RCM

[161] Ioan I. Ardelean. The relevance of cell membranes for P systems. General aspects. Fundamenta Informaticae, 49(1–3):35–43, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Atanasiu:2002:RCM

[162] Adrian Atanasiu and Carlos Martín-Vide. Recursive calculus with membranes. Fundamenta Informaticae, 49(1–3):45–59, January 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2002:SSS

[163] Gabriel Ciobanu and Dorin Paraschiv. P system software simulator. Fundamenta Informaticae, 49(1–3):61–66, January 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2002:GES

[164] Gabriel Ciobanu and Bogdan Tanasă. Gene expression by software mechanisms. Fundamenta Informaticae, 49(1–3):67–80, January 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Freund:2002:GSF

[165] Rudolf Freund and Marion Oswald. GP systems with forbidding context. Fundamenta Informaticae, 49(1–3):81–102, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frisco:2002:DCU

[166] Pierluigi Frisco, Hendrik Jan Hoogeboom, and Paul Sant. A direct construction of a universal P system. Fundamenta Informaticae, 49(1–3): 103–122, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giavitto:2002:TSM

[167] Jean-Louis Giavitto and Olivier Michel. The topological structures of membrane computing. Fundamenta Informaticae, 49(1–3):123–145, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ji:2002:BIE

[168] Sungchul Ji. The Bhopalator: An information/energy dual model of the living cell (II). Fundamenta Informaticae, 49(1–3):147–165, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krishna:2002:PSC

[169] S. N. Krishna, K. Lakshmanan, and R. Rama. On the power of P systems with contextual rules. Fundamenta Informaticae, 49(1–3):167–178, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krithivasan:2002:CS

[170] Kamala Krithivasan and Mutyam Madhu. Contextual P systems. Fundamenta Informaticae, 49(1–3):179–189, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2002:CPM

[171] Manfred Kudlek and Victor Mitrana. Closure properties of multiset language families. Fundamenta Informaticae, 49(1–3):191–203, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manca:2002:DMA

[172] Vincenzo Manca. DNA and membrane algorithms for SAT. Fundamenta Informaticae, 49(1–3):205–221, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcus:2002:MVD

[173] Salomon Marcus. Membranes versus DNA. Fundamenta Informaticae, 49(1–3):223–227, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mate:2002:PSD

[174] José L. Maté, A. Rodríguez-Patón, and Andrés Silva. On the power of P systems with DNA-worm-objects. Fundamenta Informaticae, 49(1–3): 229–239, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nicolau:2002:CLS

[175] Dan V. Nicolau, Jr., Gerardin Solana, Florin Fulga, and Dan V. Nicolau. A C library for simulating P systems. Fundamenta Informaticae, 49(1– 3):241–248, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nishida:2002:SPS

[176] Taishin Yasunobu Nishida. Simulations of photosynthesis by a K-subset transforming system with membrane. Fundamenta Informaticae, 49(1– 3):249–259, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez-Jimenez:2002:FTS

[177] Mario J. Pérez-Jiménez and Fernando Sancho-Caparrini. A formalization of transition P systems. Fundamenta Informaticae, 49(1–3):261–271, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romero-Jimenez:2002:STM

[178] Álvaro Romero-Jiménez and Mario J. Pérez-Jiménez. Simulating Turing machines by P systems with external output. Fundamenta Informaticae, 49(1–3):273–287, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Atanasiu:2002:IPM

[179] Adrian Atanasiu, Carlos Martín-Vide, and Alexandru Mateescu. On the injectivity of the Parikh matrix mapping. Fundamenta Informaticae, 49 (4):289–299, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bodlaender:2002:RUP

[180] Hans L. Bodlaender, Michael J. Dinneen, and Bakhadyr Khoussainov. Relaxed update and partition network games. Fundamenta Informaticae, 49(4):301–312, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dziemidowicz:2002:OUL

[181] Barbara Dziemidowicz. Optimal unification and learning algorithms for categorial grammars. Fundamenta Informaticae, 49(4):313–324, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Johnson:2002:PID

[182] C. A. Johnson. Processing indefinite deductive databases under the possible model semantics. Fundamenta Informaticae, 49(4):325–347, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klopotek:2002:NBT

[183] Mieczysław Alojzy Kłopotek. A new Bayesian tree learning method with reduced time and space complexity. Fundamenta Informaticae, 49(4): 349–367, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pankowski:2002:PQL

[184] Tadeusz Pankowski. PathLog: a query language for schemaless databases of partially labeled objects. Fundamenta Informaticae, 49(4):369–395, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beauquier:2002:MDP

[185] Danièle Beauquier. Markov decision processes and deterministic Büchi automata. Fundamenta Informaticae, 50(1):1–13, January 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2002:CAP

[186] Magdalena Kacprzak. A complete axiomatization of process temporal logic. Fundamenta Informaticae, 50(1):15–31, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klaudel:2002:CCP

[187] Hanna Klaudel and Franck Pommereau. A class of composable and preemptible high-level Petri nets with an application to multi-tasking systems. Fundamenta Informaticae, 50(1):33–55, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2002:DTB

[188] Mikhail Ju. Moshkov. On decision trees for (1, 2)-Bayesian networks. Fundamenta Informaticae, 50(1):57–76, January 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sureson:2002:ASD

[189] Claude Sureson. Analytic sets in descriptive set theory and NP sets in complexity theory. Fundamenta Informaticae, 50(1):77–110, January 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2002:P

[190] Ryszard Janicki. Preface. Fundamenta Informaticae, 50(2):??, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baldamus:2002:BSC

[191] Michael Baldamus and Klaus Schneider. The BDD space complexity of different forms of concurrency. Fundamenta Informaticae, 50(2):111–133, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carmona:2002:SET

[192] Josep Carmona, Jordi Cortadella, and Enric Pastor. A structural encoding technique for the synthesis of asynchronous circuits. Fundamenta Informaticae, 50(2):135–154, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garnic:2002:OAD

[193] O. Garnic, J. Lanchares, and R. Hermida. Optimization of asynchronous delay-insensitive pipeline latency using stage reorganization and optimal stage parameter estimation. Fundamenta Informaticae, 50(2):155–174, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pietkiewicz-Koutny:2002:SEN

[194] Marta Pietkiewicz-Koutny. Synthesising elementary net systems with inhibitor arcs from step transition systems. Fundamenta Informaticae, 50 (2):175–203, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xia:2002:ASM

[195] Fei Xia and Ian Clark. Algorithms for signal and message asynchronous communication mechanisms and their analysis. Fundamenta Informaticae, 50(2):205–222, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zuberek:2002:APB

[196] W. M. Zuberek. Analysis of performance bottlenecks in multithreaded multiprocessor systems. Fundamenta Informaticae, 50(2):223–241, February 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2002:MLI

[197] Philippe Balbiani. A modal logic for indiscernibility and complementarity in information systems. Fundamenta Informaticae, 50(3–4):243–263, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chikalov:2002:ADD

[198] Igor Chikalov. On average depth of decision trees implementing Boolean functions. Fundamenta Informaticae, 50(3–4):265–284, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2002:SDT

[199] Maurice Margenstern and Yurii Rogozhin. Self-describing Turing machines. Fundamenta Informaticae, 50(3–4):285–303, August 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2002:FDI

[200] Lech Polkowski. On fractal dimension in information systems. Toward exact sets in infinite information systems. Fundamenta Informaticae, 50(3–4):305–314, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Poranen:2002:LTS

[201] Timo Poranen and Jyrki Nummenmaa. A linear time special case for MC games. Fundamenta Informaticae, 50(3–4):315–324, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Postow:2002:UUP

[202] Brian Postow, Kenneth Regan, and Carl H. Smith. UPSILON: Universal programming system with incomplete lazy object notation. Fundamenta Informaticae, 50(3–4):325–359, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Reniers:2002:CT

[203] Michel Reniers, J. F. Groote, M. B. van der Zwaag, and J. van Wamel. Completeness of timed μCRL. Fundamenta Informaticae, 50(3–4):361– 402, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2002:P

[204] Ludwik Czaja, Hans-Dieter Burkhard, and Peter Starke. Preface. Fundamenta Informaticae, 51(1–2):??, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2002:NNI

[205] Roberto Barbuti, Nocoletta De Francesco, Antonella Santone, and Luca Tesei. A notion of non-interference for timed automata. Fundamenta Informaticae, 51(1–2):1–11, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bono:2002:PPS

[206] Viviana Bono and Jerzy Tiuryn. Products and polymorphic subtypes. Fundamenta Informaticae, 51(1–2):13–41, March 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2002:PTP

[207] Ludwik Czaja. Place/transition Petri net evolutions: Recording ways, analysis and synthesis. Fundamenta Informaticae, 51(1–2):43–58, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dembinski:2002:VTA

[208] Piotr Dembiński and Agata Półrola. Verification of timed automata based on similarity. Fundamenta Informaticae, 51(1–2):59–89, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2002:DMO

[209] Bernard Farwer. Dynamic modification of object Petri nets. An application to modelling protocols with fork-join structures. Fundamenta Informaticae, 51(1–2):91–101, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2002:CSS

[210] Anna Gomolińska. A comparative study of some generalized rough approximations. Fundamenta Informaticae, 51(1–2):103–119, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2002:MDO

[211] Irina A. Lomazova. Modeling dynamic objects in distributed systems with nested Petri nets. Fundamenta Informaticae, 51(1–2):121–133, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2002:BMC

[212] Wojciech Penczek, Bożena Woźna, and Andrzej Zbrzezny. Bounded model checking for the universal fragment of CTL. Fundamenta Informaticae, 51(1–2):135–156, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2002:TOA

[213] James F. Peters, Andrzej Skowron, Jarosław Stepaniuk, and Sheela Ramanna. Towards an ontology of approximate reason. Fundamenta Informaticae, 51(1–2):157–173, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ras:2002:GAR

[214] Zbigniew W. Raś and Shishir Gupta. Global action rules in distributed knowledge systems. Fundamenta Informaticae, 51(1–2):175–184, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Richling:2002:VNF

[215] Jan Richling, Louchka Popova-Zeugmann, and Matthias Werner. Verification of non-functional properties of a composable architecture with Petrinets. Fundamenta Informaticae, 51(1–2):185–200, March 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2002:CDD

[216] Zbigniew Suraj, J. F. Peters, and W. Rzasa. A comparison of different decision algorithms used in volumetric storm cells classification. Fundamenta Informaticae, 51(1–2):201–214, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Varpaaniemi:2002:MNS

[217] Kimmo Varpaaniemi. Minimizing the number of successor states in the stubborn set method. Fundamenta Informaticae, 51(1–2):215–234, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2002:RCN

[218] Józef Winkowski. Reachability in contextual nets. Fundamenta Informaticae, 51(1–2):235–250, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2002:RTC

[219] Hans-Dieter Burkhard. Real time control for autonomous mobile robots. Fundamenta Informaticae, 51(3):251–270, March 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dunin-Keplicz:2002:CI

[220] Barbara Dunin-Kȩplicz and Rineke Verbrugge. Collective intentions. Fundamenta Informaticae, 51(3):271–295, March 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutras:2002:WAM

[221] Costas D. Koutras and Pavlos Pappas. Weaker axioms, more ranges. Fundamenta Informaticae, 51(3):297–310, March 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sen:2002:SIB

[222] Jayanta Sen and Mihir K. Chakraborty. A study of interconnections between rough and 3-valued Łukasiewicz logics. Fundamenta Informaticae, 51(3):311–324, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Becher:2002:AEH

[223] Verónica Becher and Gregory Chaitin. Another example of higher order randomness. Fundamenta Informaticae, 51(4):325–338, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Denis:2002:RFS

[224] François Denis, Aurélien Lemay, and Alain Terlutte. Residual finite state automata. Fundamenta Informaticae, 51(4):339–368, March 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gora:2002:RNC

[225] Grzegorz Góra and Arkadiusz Wojna. RIONA: A new classification system combining rule induction and instance-based learning. Fundamenta Informaticae, 51(4):369–390, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kopecek:2002:EIS

[226] Ivan Kopeček and Miroslav Novotný. On equations including strings. Fundamenta Informaticae, 51(4):391–398, March 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meduna:2002:OTR

[227] Alexander Meduna and Dušan Kolář. One-turn regulated pushdown automata and their reduction. Fundamenta Informaticae, 51(4):399–405, March 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2002:I

[228] Gianpiero Cattaneo, Bruno Durand, Giancarlo Mauri, and Thomas Worsch. Introduction. Fundamenta Informaticae, 52(1–3):??, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boccara:2002:NCC

[229] Nino Boccara and Henryk Fukś. Number-conserving cellular automaton rules. Fundamenta Informaticae, 52(1–3):1–13, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buchholz:2002:IAL

[230] Thomas Buchholz, Andreas Klein, and Martin Kutrib. On interacting automata with limited nondeterminism. Fundamenta Informaticae, 52 (1–3):15–38, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2002:CSR

[231] Gianpiero Cattaneo, Alberto Dennunzio, and Luciano Margara. Chaotic subshifts and related languages applications to one-dimensional cellular automata. Fundamenta Informaticae, 52(1–3):39–80, April 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delorme:2002:PAF

[232] Marianne Delorme and Jacques Mazoyer. Pebble automata. Figures families recognition and universality. Fundamenta Informaticae, 52(1–3):81– 132, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Imai:2002:FSS

[233] Katsunobu Imai, Kenichi Morita, and Kenji Sako. Firing squad synchronization problem in number-conserving cellular automata. Fundamenta Informaticae, 52(1–3):133–141, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurka:2002:SSC

[234] Petr Kůrka and Alejandro Maass. Stability of subshifts in cellular automata. Fundamenta Informaticae, 52(1–3):143–155, April 2002. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin:2002:GHG

[235] Bruno Martin. A geometrical hierarchy on graphs via cellular automata. Fundamenta Informaticae, 52(1–3):157–181, April 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Merkle:2002:FLR

[236] Daniel Merkle and Thomas Worsch. Formal language recognition by stochastic cellular automata. Fundamenta Informaticae, 52(1–3):183– 201, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nichitiu:2002:ESH

[237] Codrin M. Nichitiu and Eric Rémila. Effective simulations on hyperbolic networks. Fundamenta Informaticae, 52(1–3):203–231, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sanders:2002:CAE

[238] Peter Sanders, Roland Vollmar, and Thomas Worsch. Cellular automata: Energy consumption and physical feasibility. Fundamenta Informaticae, 52(1–3):233–248, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sutner:2002:CAI

[239] K. Sutner. Cellular automata and intermediate reachability problems. Fundamenta Informaticae, 52(1–3):249–256, April 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umeo:2002:DRT

[240] Hiroshi Umeo and Naoki Kamikawa. A design of real-time non-regular sequence generation algorithms and their implementations on cellular automata with 1-bit inter-cell communications. Fundamenta Informaticae, 52(1–3):257–275, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Weimar:2002:TDC

[241] Jörg R. Weimar. Three-dimensional cellular automata for reactiondiffusion systems. Fundamenta Informaticae, 52(1–3):277–284, April 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dey:2002:SAC

[242] Sabyasachi Dey, Bhargab B. Bhattacharya, Malay K. Kundu, and Tinku Acharya Acharya. A simple architecture for computing moments and orientation of an image. Fundamenta Informaticae, 52(4):285–295, December 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eshuis:2002:CRF

[243] Rik Eshuis and Maarten M. Fokkinga. Comparing refinements for failure and bisimulation semantics. Fundamenta Informaticae, 52(4):297–321, December 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fan:2002:MFD

[244] Tuan-Fang Fan, Churn-Jung Liau, and Yiyu Yao. On modal and fuzzy decision logics based on rough set theory. Fundamenta Informaticae, 52(4):323–344, December 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:2002:CPG

[245] Miroslav Novotný. Construction of pure grammars. Fundamenta Informaticae, 52(4):345–360, December 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okhotin:2002:OVC

[246] Alexander Okhotin, Kai Salomaa, and Michael Domaratzki. One-visit caterpillar tree automata. Fundamenta Informaticae, 52(4):361–375, December 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2002:LACa

[247] Tomasz Terlikowski. A logical approach to control. Some new logical concepts and their application to the notion of safety for control variants. Fundamenta Informaticae, 52(4):377–394, December 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demri:2002:ATD

[248] Stéphane Demri and Ulrike Sattler. Automata-theoretic decision procedures for information logics. Fundamenta Informaticae, 53(1):1–22, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Georgescu:2002:CLS

[249] George Georgescu and Andrei Popescu. Concept lattices and similarity in non-commutative fuzzy logic. Fundamenta Informaticae, 53(1):23– 54, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Middelburg:2002:PAN

[250] Kees A. Middelburg. Process algebra with nonstandard timing. Fundamenta Informaticae, 53(1):55–77, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhao:2002:CUE

[251] Xishun Zhao and Paolo Liberatore. Complexity of the unique extension problem in default logic. Fundamenta Informaticae, 53(1):79–104, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amadio:2002:MDN

[252] Roberto Amadio, Gerard Boudol, and Cedric Lhoussaine. On message deliverability and non-uniform receptivity. Fundamenta Informaticae, 53(2):105–129, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hwang:2002:CHS

[253] Min-Shiang Hwang, Cheng-Ch Lee, and Shin-Jia Hwang. Cryptanalysis of the Hwang-shi proxy signature scheme. Fundamenta Informaticae, 53 (2):131–134, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2002:SRA

[254] Jouni Järvinen. On the structure of rough approximations. Fundamenta Informaticae, 53(2):135–153, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Midelfart:2002:LRS

[255] Herman Midelfart, Jan Komorowski, Kristin Nørsett, Fekadu Yadetie, Arne K. Sandovik, and Astrid Lægreid. Learning rough set classifiers from gene expressions and clinical data. Fundamenta Informaticae, 53 (2):155–183, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tiplea:2002:CRT

[256] Ferucio Ţiplea, Erkki Mäkinen, Dragos Trincă, and Costel Enea. Characterization results for time-varying codes. Fundamenta Informaticae, 53(2):185–198, May 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aho:2002:NPT

[257] Isto Aho. New polynomial-time instances to various knapsack-type problems. Fundamenta Informaticae, 53(3–4):199–228, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bai:2002:OOD

[258] Yun Bai and Vijay Varadharajan. Object oriented database with authorization policies. Fundamenta Informaticae, 53(3–4):229–250, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benferhat:2002:MRR

[259] Salem Benferhat, Didier Dubois, Sylvain Lagrue, and Odile Papini. Making revision reversible: an approach based on polynomials. Fundamenta Informaticae, 53(3–4):251–280, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haar:2002:PCU

[260] Stefan Haar. Probabilistic cluster unfoldings. Fundamenta Informaticae, 53(3–4):281–314, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:2002:NCA

[261] Juha Honkala. A new class of algebraic series having a decidable equivalence problem. Fundamenta Informaticae, 53(3–4):315–320, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mani:2002:REP

[262] A. Mani. Rough equalities from posets and rough difference orders. Fundamenta Informaticae, 53(3–4):321–333, August 2002. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2002:RSA

[263] Lech Polkowski and Bolesław Araszkiewicz. A rough set approach to estimating the game value and the Shapley value from data. Fundamenta Informaticae, 53(3–4):335–343, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slowinski:2002:ARI

[264] Krzysztof Słowiński, Jerzy Stefanowski, and Dariusz Siwiński. Application of rule induction and rough sets to verification of magnetic resonance diagnosis. Fundamenta Informaticae, 53(3–4):345–363, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slezak:2002:AER

[265] Dominik Ślezak. Approximate entropy reducts. Fundamenta Informaticae, 53(3–4):365–390, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2002:LACb

[266] Tomasz Terlikowski. A logical approach to control. The concept of inner implication. Fundamenta Informaticae, 53(3–4):391–397, August 2002. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Latkowski:2003:DID

[267] Rafał Latkowski. On decomposition for incomplete data. Fundamenta Informaticae, 54(1):1–16, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peltier:2003:CDP

[268] Nicolas Peltier. Constructing decision procedures in equational clausal logic. Fundamenta Informaticae, 54(1):17–65, June 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2003:RMR

[269] Lech Polkowski. Rough mereology: A rough set paradigm for unifying rough set theory and fuzzy set theory. Fundamenta Informaticae, 54(1): 67–88, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wojnarski:2003:LCA

[270] Marcin Wojnarski. LTF-C: Architecture, training algorithm and applications of new neural classifier. Fundamenta Informaticae, 54(1):89–105, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2003:Pa

[271] Hans-Dieter Burkhard, Gabriela Lindemann, Peter Starke, and Ludwik Czaja. Preface. Fundamenta Informaticae, 54(2–3):??, June 2003. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abraham:2003:ESA

[272] Uri Abraham and Tamar Pinhas. Exercises in style (alpha specifications). Fundamenta Informaticae, 54(2–3):107–135, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2003:DNT

[273] Roberto Barbuti and Luca Tesei. A decidable notion of timed noninterference. Fundamenta Informaticae, 54(2–3):137–150, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Coja-Oghlan:2003:RAP

[274] Amin Coja-Oghlan and Mark-Olivier Stehr. Revisiting the algebra of Petri net processes under the collective token philosophy. Fundamenta Informaticae, 54(2–3):151–164, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2003:PNC

[275] Ludwik Czaja. Proving nets correct via cause-effect structures (an experiment). Fundamenta Informaticae, 54(2–3):165–183, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2003:SCH

[276] Berndt Farwer, Manfred Kudlek, and Kundan Misra. Some considerations on higher-order Petri nets. Fundamenta Informaticae, 54(2–3): 185–193, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeFrancesco:2003:AIM

[277] Nicolette De Francesco, Antonella Santone, and Luca Tesei. Abstract interpretation and model checking for checking secure information flow in concurrent systems. Fundamenta Informaticae, 54(2–3):195–211, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2003:UMA

[278] Magdalena Kacprzak. Undecidability of a multi-agent logic. Fundamenta Informaticae, 54(2–3):213–220, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler:2003:CMO

[279] Michael Köhler and Heiko Rölke. Concurrency in mobile object net systems. Fundamenta Informaticae, 54(2–3):221–235, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lanotte:2003:DHM

[280] Ruggero Lanotte, Andrea Maggiolo-Schettini, Adriano Peron, and Simone Tini. Dynamic hierarchical machines. Fundamenta Informaticae, 54(2–3):237–252, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schmidt:2003:DVL

[281] Karsten Schmidt. Distributed verification with LoLA. Fundamenta Informaticae, 54(2–3):253–262, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2003:RSI

[282] Andrzej Skowron, Jarosław Stepaniuk, and James F. Peters. Rough sets and infomorphisms: Towards approximation of relations in distributed environments. Fundamenta Informaticae, 54(2–3):263–277, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Varpaaniemi:2003:TAA

[283] Kimmo Varpaaniemi. Towards ambitious approximation algorithms in stubborn set optimization. Fundamenta Informaticae, 54(2–3):279–294, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2003:ABC

[284] Raymond Devillers, Hanna Klaudel, Maciej Koutny, and Franck Pommereau. Asynchronous Box calculus. Fundamenta Informaticae, 54(4): 295–344, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2003:CIIa

[285] Mikhail Ju. Moshkov. Classification of infinite information systems depending on complexity of decision trees and decision rule systems. Fundamenta Informaticae, 54(4):345–368, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2003:DIF

[286] Tomasz Terlikowski. Descriptional independence and the formal definition of sequential control structure. Fundamenta Informaticae, 54(4): 369–385, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wakulicz-Deja:2003:ARS

[287] Alicja Wakulicz-Deja and Piotr Peszek. Applying rough set theory to multi stage medical diagnosing. Fundamenta Informaticae, 54(4):387– 408, June 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bloom:2003:DWF

[288] Stephen L. Bloom and Zoltán Ésik. Deciding whether the frontier of a regular tree is scattered. Fundamenta Informaticae, 55(1):1–21, July 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2003:TSI

[289] Jean-Marc Champarnaud and Fabien Coulon. Theoretical study and implementation of the canonical automaton. Fundamenta Informaticae, 55(1):23–38, July 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2003:CAH

[290] Liviu P. Dinu. On the classification and aggregation of hierarchies with different constitutive elements. Fundamenta Informaticae, 55(1):39–50, July 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2003:CIIb

[291] Mikhail Ju. Moshkov. Compressible infinite information systems. Fundamenta Informaticae, 55(1):51–61, July 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2003:FSS

[292] Linh Anh Nguyen. A fixpoint semantics and an SLD-resolution calculus for modal logic programs. Fundamenta Informaticae, 55(1):63–100, July 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2003:Pb

[293] Hans-Dieter Burkhard, Gabriela Lindemann, Peter Starke, and Ludwik Czaja. Preface. Fundamenta Informaticae, 55(2):??, August 2003. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bashkin:2003:PNR

[294] Vladimir A. Bashkin and Irina A. Lomazova. Petri nets and resource bisimulation. Fundamenta Informaticae, 55(2):101–114, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2003:APU

[295] Marco Bellia and Eugenia M. Occhiuto. N-axioms parallel unification. Fundamenta Informaticae, 55(2):115–128, August 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2003:MHO

[296] Berndt Farwer and Kundan Misra. Modelling with hierarchical object Petri nets. Fundamenta Informaticae, 55(2):129–147, August 2003. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pancerz:2003:SPN

[297] Krzysztof Pancerz and Zbigniew Suraj. Synthesis of Petri net models: A rough set approach. Fundamenta Informaticae, 55(2):149–165, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2003:VEP

[298] Wojciech Penczek and Alessio Lomuscio. Verifying epistemic properties of multi-agent systems via bounded model checking. Fundamenta Informaticae, 55(2):167–185, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popova-Zeugmann:2003:USE

[299] Louchka Popova-Zeugmann, Matthias Werner, and Jan Richling. Using state equation to prove non-reachability in timed Petrinets. Fundamenta Informaticae, 55(2):187–202, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polrola:2003:RAT

[300] Agata Półrola, Wojciech Penczek, and Maciej Szreter. Reachability analysis for timed automata using partitioning algorithms. Fundamenta Informaticae, 55(2):203–221, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna:2003:CRP

[301] Bożena Woźna, Andrzej Zbrzezny, and Wojciech Penczek. Checking reachability properties for timed automata via SAT. Fundamenta Informaticae, 55(2):223–241, August 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aydin:2003:SVP

[302] Suna Aydin and Henning Bordihn. Sequential versus parallel grammar formalisms with respect to measures of descriptional complexity. Fundamenta Informaticae, 55(3–4):243–254, September 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2003:SSR

[303] Gianpiero Cattaneo and Davide Ciucci. Shadowed sets and related algebraic structures. Fundamenta Informaticae, 55(3–4):255–284, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chlebus:2003:DCP

[304] Bogdan S. Chlebus, Leszek Gasieniec, and Andrzej Pelc. Deterministic computations on a PRAM with static processor and memory faults. Fundamenta Informaticae, 55(3–4):285–306, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2003:UBS

[305] Cheng-Chi Lee, Wei-Pang Yang, and Min-Shiang Hwang. Untraceable blind signature schemes based on discrete logarithm problem. Fundamenta Informaticae, 55(3–4):307–320, September 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diaconescu:2003:IIU

[306] Răzvan Diaconescu. Institution-independent ultraproducts. Fundamenta Informaticae, 55(3–4):321–348, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kandulski:2003:DTL

[307] Maciej Kandulski. Derived tree languages of nonassociative Lambek categorial grammars with product. Fundamenta Informaticae, 55(3–4):349– 362, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2003:ERC

[308] Sanjiang Li and Mingsheng Ying. Extensionality of the RCC8 composition table. Fundamenta Informaticae, 55(3–4):363–385, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Weicker:2003:BPU

[309] Karsten Weicker and Nicole Weicker. Basic principles for understanding evolutionary algorithms. Fundamenta Informaticae, 55(3–4):387–403, September 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Smyth:2003:F

[310] Bill Smyth. Foreword. Fundamenta Informaticae, 56(1–2):??, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crochemore:2003:OSH

[311] Maxime Crochemore, Costas S. Iliopoulos, Thierry Lecroq, Yoan J. Pinzon, Wojciech Plandowski, and Wojciech Rytter. Occurrence and substring heuristics for δ-matching. Fundamenta Informaticae, 56(1–2):1–21, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuri:2003:FMS

[312] Josué Kuri, Gonzalo Navarro, and Ludovic Mé. Fast multipattern search algorithms for intrusion detection. Fundamenta Informaticae, 56(1–2): 23–49, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhardt:2003:BFG

[313] Stefan Burkhardt and Juha Kärkkäinen. Better filtering with gapped qgrams. Fundamenta Informaticae, 56(1–2):51–70, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nicodeme:2003:RSP

[314] Pierre Nicodème. Regexpcount, a symbolic package for counting problems on regular expressions and words. Fundamenta Informaticae, 56 (1–2):71–88, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crochemore:2003:SHH

[315] Maxime Crochemore, Costas S. Iliopoulous, and Yoan J. Pinzon. Speeding-up Hirschberg and Hunt–Szymanski LCS algorithms. Fundamenta Informaticae, 56(1–2):89–103, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jansson:2003:FAO

[316] Jesper Jansson and Andrzej Lingas. A fast algorithm for optimal alignment between similar ordered trees. Fundamenta Informaticae, 56(1–2): 105–120, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beal:2003:CFW

[317] Marie-Pierre Béal, Maxime Crochemore, Filippo Mignosi, Antonio Restivo, and Marinella Sciortino. Computing forbidden words of regular languages. Fundamenta Informaticae, 56(1–2):121–135, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasieniec:2003:TSE

[318] Leszek Gasieniec and Igor Potapov. Time/space efficient compressed pattern matching. Fundamenta Informaticae, 56(1–2):137–154, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Afrati:2003:DC

[319] Foto Afrati, Hans Leiß, and Michel de Rougemont. Definability and compression. Fundamenta Informaticae, 56(1–2):155–180, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2003:SSA

[320] Inbok Lee, Kunsoo Park, Yanghee Choi, and Sung Kwon Chung. A simple and scalable algorithm for the IP address lookup problem. Fundamenta Informaticae, 56(1–2):181–190, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makinen:2003:CSA

[321] Veli Mäkinen. Compact suffix array — a space-efficient full-text index. Fundamenta Informaticae, 56(1–2):191–210, January 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2003:EUD

[322] Philippe Balbiani. Eliminating unorthodox derivation rules in an axiom system for iteration-free PDL with intersection. Fundamenta Informaticae, 56(3):211–242, November 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huang:2003:EPT

[323] Hui-Feng Huang and Chin-Chen Chang. An efficient and practical (t, n) threshold proxy signature scheme with known signers. Fundamenta Informaticae, 56(3):243–253, November 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2003:IHP

[324] Maurice Margenstern. On the infinigons of the hyperbolic plane, a combinatorial approach. Fundamenta Informaticae, 56(3):255–272, November 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ord:2003:ENF

[325] Toby Ord and Tien D. Kieu. On the existence of a new family of Diophantine equations for Ω. Fundamenta Informaticae, 56(3):273–284, November 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wojna:2003:CBI

[326] Arkadiusz Wojna. Center-based indexing in vector and metric spaces. Fundamenta Informaticae, 56(3):285–310, November 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ceterchi:2003:TLS

[327] Rodica Ceterchi, Radu Gramatovici, Nataša Jonoska, and K. G. Subramanian. Tissue-like P systems with active membranes for picture generation. Fundamenta Informaticae, 56(4):311–328, December 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dunin-Keplicz:2003:ECC

[328] Barbara Dunin-Kȩplicz and Rineke Verbrugge. Evolution of collective commitment during teamwork. Fundamenta Informaticae, 56(4):329– 371, December 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2003:NED

[329] Tomasz Terlikowski. On the notion of elementary description. Fundamenta Informaticae, 56(4):373–387, December 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Truthe:2003:FPL

[330] Bianca Truthe. On the finiteness of picture languages of synchronous, simple non-deterministic chain code picture systems. Fundamenta Informaticae, 56(4):389–409, December 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demolombe:2003:ERA

[331] Robert Demolombe. An extended relational algebra on abstract objects for summarizing answers to queries. Fundamenta Informaticae, 57(1): 1–15, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jain:2003:ICL

[332] Sanjay Jain. The intrinsic complexity of learning: A survey. Fundamenta Informaticae, 57(1):17–37, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Malerba:2003:LRT

[333] Donato Malerba. Learning recursive theories in the normal ILP setting. Fundamenta Informaticae, 57(1):39–77, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2003:EDF

[334] Shu-Ching Wang and Chieng-Fu Cheng. Eventually dual failure agreement. Fundamenta Informaticae, 57(1):79–99, January 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:2003:P

[335] Patrick Doherty, Andrzej Skowron, Witold Łukaszewicz, and Andrzej Szałas. Preface. Fundamenta Informaticae, 57(2–4):??, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dubois:2003:QPL

[336] Didier Dubois, Sébastien Konieczny, and Henri Prade. Quasi-possibilistic logic and its measures of information and conflict. Fundamenta Informaticae, 57(2–4):101–125, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Godo:2003:FML

[337] Lluís Godo, Petr Hájek, and Francesc Esteva. A fuzzy modal logic for belief functions. Fundamenta Informaticae, 57(2–4):127–146, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:2003:TFA

[338] Patrick Doherty, Michał Grabowski, Witold Łukaszewicz, and Andrzej Szałas. Towards a framework for approximate ontologies. Fundamenta Informaticae, 57(2–4):147–165, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grosskreutz:2003:PCA

[339] Henrik Grosskreutz and Gerhard Lakemeyer. Probabilistic complex actions in GOLOG. Fundamenta Informaticae, 57(2–4):167–192, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Madalinska-Bugaj:2003:FDL

[340] Ewa Madalińska-Bugaj and Witold Łukaszewicz. Formalizing defeasible logic in CAKE. Fundamenta Informaticae, 57(2–4):193–213, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vitoria:2003:RSR

[341] Aida Vitória, Carlos Viegas Damásio, and Jan Małuszyński. From rough sets to rough knowledge bases. Fundamenta Informaticae, 57(2–4):215– 246, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baader:2003:TAD

[342] Franz Baader, Jan Hladik, Carsten Lutz, and Frank Wolter. From tableaux to automata for description logics. Fundamenta Informaticae, 57(2–4):247–279, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Harrenstein:2003:MCN

[343] Paul Harrenstein, Wiebe van der Hoek, John-Jules Meyer, and Cees Witteveen. A modal characterization of Nash equilibrium. Fundamenta Informaticae, 57(2–4):281–321, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herzig:2003:MPB

[344] Andreas Herzig. Modal probability, belief, and actions. Fundamenta Informaticae, 57(2–4):323–344, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dix:2003:MAU

[345] Jürgen Dix, Thomas Eiter, Michael Fink, Axel Polleres, and Yingqian Zhang. Monitoring agents using declarative planning. Fundamenta Informaticae, 57(2–4):345–370, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fichtner:2003:IEM

[346] Matthias Fichtner, Axel Großmann, and Michael Thielscher. Intelligent execution monitoring in dynamic environments. Fundamenta Informaticae, 57(2–4):371–392, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konczak:2003:GCA

[347] Kathrin Konczak, Torsten Schaub, and Thomas Linke. Graphs and colorings for answer set programming with preferences. Fundamenta Informaticae, 57(2–4):393–421, February 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Corradini:2003:P

[348] Andrea Corradini and Hans-Jörg Kreowski. Preface. Fundamenta Informaticae, 58(1):??, March 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gyapay:2003:GTT

[349] Szilvia Gyapay, Dániel Varró, and Reiko Heckel. Graph transformation with time. Fundamenta Informaticae, 58(1):1–22, March 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hage:2003:EGT

[350] Jurriaan Hage, Tero Harju, and Emo Welzl. Euler graphs, triangle-free graphs and bipartite graphs in switching classes. Fundamenta Informaticae, 58(1):23–37, March 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoffmann:2003:ACS

[351] Berthold Hoffmann. Abstraction and control for shapely nested graph transformation. Fundamenta Informaticae, 58(1):39–65, March 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2003:SPC

[352] Artiom Alhazov, Carlos Martín-Vide, and Linqiang Pan. Solving a PSPACE-complete problem by recognizing P systems with restricted active membranes. Fundamenta Informaticae, 58(2):67–77, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Belohlavek:2003:FCO

[353] Radim Bělohlávek. Fuzzy closure operators induced by similarity. Fundamenta Informaticae, 58(2):79–91, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chan:2003:NCI

[354] Chao-Wen Chan and Chin-Chen Chang. A new channel for image hiding based on interpolating polynomials. Fundamenta Informaticae, 58(2):93– 104, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kllopotek:2003:RLE

[355] Mieczysław Kłlopotek. Reasoning and learning in extended structured Bayesian networks. Fundamenta Informaticae, 58(2):105–137, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kou:2003:WLA

[356] Hui Kou and Maokang Luo. On (weakly) local approximation spaces of information systems. Fundamenta Informaticae, 58(2):139–150, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2003:RCN

[357] Tomasz Terlikowski. Risk in control. On the notion of risk function. Fundamenta Informaticae, 58(2):151–165, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsai:2003:IIS

[358] Chwei-Shyong Tsai and Chin-Chen Chang. An improvement to image segment based on human visual system for object-based coding. Fundamenta Informaticae, 58(2):167–178, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsai:2003:DBD

[359] Chwei-Shyong Tsai and Chin-Chen Chang. A DWT-based deblocking technique based on edge detection and zero masking for reducing blocking effect. Fundamenta Informaticae, 58(2):179–188, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2003:HDW

[360] Hsien-Chu Wu and Chin-Chen Chang. Hiding digital watermarks using fractal compression technique. Fundamenta Informaticae, 58(2):189–202, April 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2003:F

[361] Martin Kutrib, Maurice Margenstern, and Hiroshi Umeo. Foreword. Fundamenta Informaticae, 58(3–4):??, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurka:2003:CAV

[362] Petr Kůrka. Cellular automata with vanishing particles. Fundamenta Informaticae, 58(3–4):203–221, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sutner:2003:APC

[363] Klaus Sutner. Almost periodic configurations on linear cellular automata. Fundamenta Informaticae, 58(3–4):223–240, May 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Worsch:2003:SBC

[364] Thomas Worsch. Simulations between cellular automata on trees extended by horizontal edges. Fundamenta Informaticae, 58(3–4):241–260, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Iwamoto:2003:SBM

[365] Chuzo Iwamoto, Katsuyuki Tateishi, Kenichi Morita, and Katsunobu Imai. Simulations between multi-dimensional deterministic and alternating cellular automata. Fundamenta Informaticae, 58(3–4):261–271, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2003:STB

[366] Martin Kutrib and Jan-Thomas Löwe. Space- and time-bounded nondeterminism for cellular automata. Fundamenta Informaticae, 58(3–4): 273–293, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2003:EUD

[367] Jia Lee, Susumu Adachi, Ferdinand Peper, and Kenichi Morita. Embedding universal delay-insensitive circuits in asynchronous cellular spaces. Fundamenta Informaticae, 58(3–4):295–320, May 2003. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2003:TAC

[368] Pradipta Maji, Chandrama Shaw, Niloy Ganguly, Biplab K. Sikdar, and P. Pal Chaudhuri. Theory and application of cellular automata for pattern classification. Fundamenta Informaticae, 58(3–4):321–354, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Malcher:2003:OWC

[369] Andreas Malcher. On one-way cellular automata with a fixed number of cells. Fundamenta Informaticae, 58(3–4):355–368, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2003:TDC

[370] Maurice Margenstern and Gencho Skordev. Tools for devising cellular automata in the hyperbolic 3D space. Fundamenta Informaticae, 58 (3–4):369–398, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nishio:2003:IDC

[371] Hidenosuke Nishio and Takashi Saito. Information dynamics of cellular automata I — an algebraic study. Fundamenta Informaticae, 58(3–4): 399–420, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umeo:2003:RTG

[372] Hiroshi Umeo and Naoki Kamikawa. Real-time generation of primes by a 1-bit-communication cellular automaton. Fundamenta Informaticae, 58 (3–4):421–435, May 2003. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burton:2004:RCP

[373] Jonathan Burton, Maciej Koutny, and Giuseppe Pappalardo. Relating communicating processes with different interfaces. Fundamenta Informaticae, 59(1):1–37, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghosh:2004:NPS

[374] Sujata Ghosh and Mihir Kr. Chakraborty. Nonmonotonic proof systems: Algebraic foundations. Fundamenta Informaticae, 59(1):39–65, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jonoska:2004:TOS

[375] Nataša Jonoska and Maurice Margenstern. Tree operations in P systems and λ-calculus. Fundamenta Informaticae, 59(1):67–90, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2004:RTD

[376] Tomasz Terlikowski. Risk-type description of control variants. Fundamenta Informaticae, 59(1):91–106, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2004:P

[377] Guoyin Wang, Qing Liu, Tsau Young Lin, Yi Yu Yao, and Lech Polkowski. Preface. Fundamenta Informaticae, 59(2–3):??, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2004:VRS

[378] Jan Bazan, Hung Son Nguyen, and Marcin Szczuka. A view on rough set concept approximations. Fundamenta Informaticae, 59(2–3):107–118, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guan:2004:DMD

[379] J. W. Guan, D. Y. Liu, and D. A. Bell. Discovering motifs in DNA sequences. Fundamenta Informaticae, 59(2–3):119–134, July 2004. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2004:NRS

[380] Xiaohua Hu, T. Y. Lin, and Jianchao Han. A new rough sets model based on database systems. Fundamenta Informaticae, 59(2–3):135–152, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2004:IMS

[381] Dan Li and Jitender S. Deogun. Interpolation models for spatiotemporal association mining. Fundamenta Informaticae, 59(2–3):153–172, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2004:DMD

[382] Hongsong Li, Houkuan Huang, and Youfang Lin. DSD: Maintain data cubes more efficiently. Fundamenta Informaticae, 59(2–3):173–190, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazlack:2004:IC

[383] Lawrence J. Mazlack. Imperfect causality. Fundamenta Informaticae, 59 (2–3):191–201, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Miao:2004:CRS

[384] Duoqian Miao and Lishan Hou. A comparison of rough set methods and representative inductive learning algorithms. Fundamenta Informaticae, 59(2–3):203–219, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pagliani:2004:PDS

[385] Piero Pagliani. Pretopologies and dynamic spaces. Fundamenta Informaticae, 59(2–3):221–239, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2004:RIM

[386] Andrzej Skowron and Piotr Synak. Reasoning in information maps. Fundamenta Informaticae, 59(2–3):241–259, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2004:RSA

[387] Tuan Trung Nguyen. Rough set approach to domain knowledge approximation. Fundamenta Informaticae, 59(2–3):261–270, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2004:ESM

[388] Shusaku Tsumoto. Extraction of structure of medical diagnosis from clinical data. Fundamenta Informaticae, 59(2–3):271–285, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2004:QST

[389] Ling Zhang and Bo Zhang. The quotient space theory of problem solving. Fundamenta Informaticae, 59(2–3):287–298, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zheng:2004:RRS

[390] Zheng Zheng and Guoyin Wang. RRIA: A rough set and rule tree based incremental knowledge acquisition algorithm. Fundamenta Informaticae, 59(2–3):299–313, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baaz:2004:ACM

[391] Matthias Baaz, Agata Ciabattoni, and Franco Montagna. Analytic calculi for monoidal T -norm based logic. Fundamenta Informaticae, 59(4): 315–332, December 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dobrev:2004:LER

[392] Stefan Dobrev and Andrzej Pelc. Leader election in rings with nonunique labels. Fundamenta Informaticae, 59(4):333–347, December 2004. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Domaratzki:2004:TBE

[393] Michael Domaratzki. Trajectory-based embedding relations. Fundamenta Informaticae, 59(4):349–363, December 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hartmann:2004:ERF

[394] Sven Hartmann, Sebastian Link, Klaus-Dieter Schewe, and Bernhard Thalheim. Error-robust functional dependencies. Fundamenta Informaticae, 59(4):365–372, December 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moschoyiannis:2004:STF

[395] Sotiris Moschoyiannis and Michael W. Shields. A set-theoretic framework for component composition. Fundamenta Informaticae, 59(4):373–396, December 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2004:P

[396] Hans-Dieter Burkhard, Gabriela Lindemann, Ludwik Czaja, and Zbigniew Suraj. Preface. Fundamenta Informaticae, 60(1–4):??, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2004:BRN

[397] Antoni Mazurkiewicz. Bilateral ranking negotiations. Fundamenta Informaticae, 60(1–4):1–16, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mosses:2004:ELS

[398] Peter D. Mosses. Exploiting labels in structural operational semantics. Fundamenta Informaticae, 60(1–4):17–31, September 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pawlak:2004:DRD

[399] Zdzisław Pawlak. Decision rules and dependencies. Fundamenta Informaticae, 60(1–4):33–39, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ambroszkiewicz:2004:ELD

[400] Stanisław Ambroszkiewicz. Entish: A language for describing data processing in open distributed systems. Fundamenta Informaticae, 60(1–4): 41–66, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2004:AIA

[401] Roberto Barbuti, Stefano Cataudella, and Luca Tesei. Abstract interpretation against races. Fundamenta Informaticae, 60(1–4):67–79, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardeschi:2004:CAS

[402] Cinzia Bernardeschi, Nicoletta De Francesco, and Giuseppe Lettieri. Concrete and abstract semantics to check secure information flow in concurrent programs. Fundamenta Informaticae, 60(1–4):81–98, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doroshenko:2004:MPP

[403] Analoliy Doroshenko and Georgy Tseitlin. Models and parallel programming abstractions to enhance concurrency of parallel programs. Fundamenta Informaticae, 60(1–4):99–111, September 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2004:MON

[404] Berndt Farwer and Michael Köhler. Mobile object-net systems and their processes. Fundamenta Informaticae, 60(1–4):113–129, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2004:NSM

[405] Berndt Farwer and Manfred Kudlek. A new synchronisation model for multi-level object Petri nets. Fundamenta Informaticae, 60(1–4):131– 142, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fryc:2004:FRU

[406] B. Fryc, K. Pancerz, J. F. Peters, and Z. Suraj. On fuzzy reasoning using matrix representation of extended fuzzy Petri nets. Fundamenta Informaticae, 60(1–4):143–157, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2004:GMF

[407] Anna Gomolińska. A graded meaning of formulas in approximation spaces. Fundamenta Informaticae, 60(1–4):159–172, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heckel:2004:DMU

[408] Martin Heckel and Jaroslav Zendulka. Data mining and its use in texture analysis. Fundamenta Informaticae, 60(1–4):173–186, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janowska:2004:STS

[409] Agata Janowska and Paweł Janowski. Slicing timed systems. Fundamenta Informaticae, 60(1–4):187–210, September 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konikowka:2004:DVM

[410] Beata Konikowka and Wojciech Penczek. On designated values in multivalued CTL^{∗} model checking. Fundamenta Informaticae, 60(1–4):211– 224, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazowa:2004:CIA

[411] Irina A. Lomazowa. Communities of interacting automata for modelling distributed systems with dynamic structure. Fundamenta Informaticae, 60(1–4):225–235, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ochmanski:2004:BFH

[412] Edward Ochmański. Best fairness hierarchy in elementary nets. Fundamenta Informaticae, 60(1–4):237–250, September 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pancerz:2004:DCM

[413] Krzysztof Pancerz and Zbigniew Suraj. Discovering concurrent models from data tables with the ROSECON system. Fundamenta Informaticae, 60(1–4):251–268, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polak:2004:PET

[414] Michał Polak, Paweł Majdzik, Zbigniew Banaszak, and Robert Wójcik. The performance evaluation tool for automated prototyping of concurrent cyclic processes. Fundamenta Informaticae, 60(1–4):269–289, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2004:SRS

[415] Lech Polkowski and Maria Semeniuk-Polkowska. Some remarks on sets of communicating sequential processes in topological rough set framework. Fundamenta Informaticae, 60(1–4):291–305, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polrola:2004:MAT

[416] Agata Półrola and Wojciech Penczek. Minimization algorithms for time Petri nets. Fundamenta Informaticae, 60(1–4):307–331, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2004:AOP

[417] Roman R. Redziejowski. On associative Omega-products. Fundamenta Informaticae, 60(1–4):333–350, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2004:CP

[418] Andrzej Skowron and Piotr Synak. Complex patterns. Fundamenta Informaticae, 60(1–4):351–366, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tini:2004:CSG

[419] Simone Tini and Andrea Maggiolo-Schettini. Compositional synthesis of generalized Mealy machines. Fundamenta Informaticae, 60(1–4):367– 382, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Virbitskaite:2004:OMO

[420] Irina B. Virbitskaite and Natalya S. Gribovskaya. Open maps and observational equivalences for timed partial order models. Fundamenta Informaticae, 60(1–4):383–399, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2004:GCD

[421] Marcin Wolski. Galois connections and data analysis. Fundamenta Informaticae, 60(1–4):401–415, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2004:ISB

[422] Andrzej Zbrzezny. Improvements in SAT-based reachability analysis for timed automata. Fundamenta Informaticae, 60(1–4):417–434, September 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bandini:2004:IFI

[423] Stefania Bandini and Carla Simone. Integrating forms of interaction in a distributed model. Fundamenta Informaticae, 61(1):1–17, October 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grigorieff:2004:RCA

[424] Serge Grigorieff and Maurice Margenstern. Register cellular automata in the hyperbolic plane. Fundamenta Informaticae, 61(1):19–27, October 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kislak-Malinowska:2004:CP

[425] Aleksandra Kiślak-Malinowska. Conjoinability in pregroups. Fundamenta Informaticae, 61(1):29–36, October 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2004:NVR

[426] Lech Polkowski. A note on 3-valued rough logic accepting decision rules. Fundamenta Informaticae, 61(1):37–45, October 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2004:ACC

[427] Roman R. Redziejowski. Asynchronous circuits, communicating processes, and Muller automaton. Fundamenta Informaticae, 61(1):47–59, October 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudeanu:2004:SOR

[428] Sergiu Rudeanu and Dragoş Vaida. Semirings in operations research and computer science: More algebra. Fundamenta Informaticae, 61(1):61–85, October 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2004:P

[429] Zbigniew Suraj. Preface. Fundamenta Informaticae, 61(2):??, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshov:2004:COD

[430] Mikhail Ju. Moshov and Igor V. Chikalov. Consecutive optimization of decision trees concerning various complexity measures. Fundamenta Informaticae, 61(2):87–96, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pancerz:2004:DAC

[431] Krzysztof Pancerz and Zbigniew Suraj. Discovery of asynchronous concurrent models from experimental tables. Fundamenta Informaticae, 61 (2):97–116, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pindur:2004:HAD

[432] Roman Pindur, Robert Susmaga, and Jerzy Stefanowski. Hyperplane aggregation of dominance decision rules. Fundamenta Informaticae, 61(2):

117–137, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:2004:LFO

[433] Jarosław Stepaniuk and Piotr Hońko. Learning first-order rules: A rough set approach. Fundamenta Informaticae, 61(2):139–157, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Susmaga:2004:RCA

[434] Robert Susmaga. Reducts and constructs in attribute reduction. Fundamenta Informaticae, 61(2):159–181, November 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2004:LAP

[435] Jan A. Bergstra and Cornelis A. Middelburg. Located actions in process algebra with timing. Fundamenta Informaticae, 61(3–4):183–211, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chan:2004:RVB

[436] Yung-Kuan Chan, Her-Fa Wang, and Chin-Fan Lee. A refined VQ-based image compression method. Fundamenta Informaticae, 61(3–4):213–221, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Georgescu:2004:CCF

[437] Irina Georgescu. Consistency conditions in fuzzy consumers theory. Fundamenta Informaticae, 61(3–4):223–245, August 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janin:2004:RBM

[438] David Janin and Giacomo Lenzi. On the rlationship between monadic and weak monadic second order logic on arbitrary trees, with applications to the mu-calculus. Fundamenta Informaticae, 61(3–4):247–265, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kozera:2004:ALT

[439] Ryszard Kozera. Asymptotics for length and trajectory from cumulative chord piecewise-quartics. Fundamenta Informaticae, 61(3–4):267–283, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kozera:2004:ICC

[440] Ryszard Kozera and Lyle Noakes. C^{1} interpolation with cumulative chord cubics. Fundamenta Informaticae, 61(3–4):285–301, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:2004:DLF

[441] Władysław Skarbek, Krzysztof Kucharski, and Mirosław Bober. Dual LDA for face recognition. Fundamenta Informaticae, 61(3–4):303–334, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:2004:FIF

[442] Władysław Skarbek, Thomas Sikora, and Grzegorz Galinski. Fast index filtering in vector approximation file. Fundamenta Informaticae, 61(3– 4):335–349, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Werner:2004:MPN

[443] Matthias Werner, Louchka Popova-Zeugmann, and Jan Richling. A method to prove non-reachability in priority duration Petri nets. Fundamenta Informaticae, 61(3–4):351–368, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wnukowicz:2004:IID

[444] Karol Wnukowicz. Image indexing by distributed color temperature descriptions. Fundamenta Informaticae, 61(3–4):369–378, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2004:SIS

[445] Yuan-Hui Yu and Chin-Chen Chang. Scenery image segmentation using support vector machines. Fundamenta Informaticae, 61(3–4):379–388, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hwang:2004:EKA

[446] Min-Shiang Hwang, Jung-Wen Lo, and Chia-Hsin Liu. Enhanced of key agreement protocols resistant to a denial-of-service attack. Fundamenta Informaticae, 61(3–4):389–398, August 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trakhtenbrot:2004:P

[447] Boris A. Trakhtenbrot. Preface. Fundamenta Informaticae, 62(1):??, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hirshfeld:2004:LRT

[448] Yoram Hirshfeld and Alexander Rabinovich. Logics for real time: Decidability and complexity. Fundamenta Informaticae, 62(1):1–28, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slissenko:2004:LFV

[449] Anatol Slissenko. A logic framework for verification of timed algorithms. Fundamenta Informaticae, 62(1):29–67, January 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trakhtenbrot:2004:UBA

[450] B. A. Trakhtenbrot. Understanding basic automata theory in the continuous time setting. Fundamenta Informaticae, 62(1):69–121, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pardo:2004:SCC

[451] D. Pardo, A. Rabinovich, and B. A. Trakhtenbrot. Synchronous circuits over continuous time: Feedback reliability and mpleteness. Fundamenta Informaticae, 62(1):123–137, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2004:P

[452] Ryszard Janicki. Preface. Fundamenta Informaticae, 62(2):??, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bontemps:2004:SOR

[453] Yves Bontemps, Pierre-Yves Schobbens, and Christof Löding. Synthesis of open reactive systems from scenario-based specifications. Fundamenta Informaticae, 62(2):139–169, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cortadella:2004:QSS

[454] Jordi Cortadella, Alex Kondratyev, Luciano Lavagno, Alexander Taubin, and Yosinori Watanabe. Quasi-static scheduling for concurrent architectures. Fundamenta Informaticae, 62(2):171–196, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:2004:PTA

[455] Javier Esparza. A polynomial-time algorithm for checking consistency of free-choice signal transition graphs. Fundamenta Informaticae, 62(2): 197–220, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khomenko:2004:DSE

[456] Victor Khomenko, Maciej Koutny, and Alex Yakovlev. Detecting state encoding conflicts in STG unfoldings using SAT. Fundamenta Informaticae, 62(2):221–241, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Talpin:2004:FRC

[457] Jean-Pierre Talpin, Paul Le Guernic, Sandeep Kumar Shukla, Frédéric Doucet, and Rajesh Gupta. Formal refinement checking in a system-level design methodology. Fundamenta Informaticae, 62(2):243–273, February 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2004:ASS

[458] Chin-Chen Chang, Piyu Tsai, and Min-Hui Lin. An adaptive steganographic scheme for color images. Fundamenta Informaticae, 62(3–4):275– 289, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cassez:2004:TEA

[459] Franck Cassez, Claire Pagetti, and Olivier Roux. A timed extension for ALTARICA. Fundamenta Informaticae, 62(3–4):291–332, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2004:PFL

[460] Olivier Finkel. An ω-power of a finitary language which is a Borel set of infinite rank. Fundamenta Informaticae, 62(3–4):333–342, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Johnson:2004:QCU

[461] Chris A. Johnson. Query compilation under the disjunctive well-founded semantics. Fundamenta Informaticae, 62(3–4):343–368, March 2004. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karpinski:2004:AAM

[462] Marek Karpinski, Mirosław Kowaluk, and Andrzej Lingas. Approximation algorithms for MAX-BISECTION on low degree regular graphs. Fundamenta Informaticae, 62(3–4):369–375, March 2004. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liberatore:2004:CBR

[463] Paolo Liberatore and Marco Schaerf. The compactness of belief revision and update operators. Fundamenta Informaticae, 62(3–4):377–393, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marciniec:2004:OUI

[464] Jacek Marciniec. Optimal unification of infinite sets of types. Fundamenta Informaticae, 62(3–4):395–407, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

De:2004:RST

[465] Supriya Kumar De. A rough set theoretic approach to clustering. Fundamenta Informaticae, 62(3–4):409–417, March 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fredriksson:2004:ESM

[466] Kimmo Fredriksson and Jorma Tarhio. Efficient string matching in Huffman compressed texts. Fundamenta Informaticae, 63(1):1–16, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2004:KAUa

[467] Richard Holzer. Knowledge acquisition under incomplete knowledge using methods from formal concept analysis: Part I. Fundamenta Informaticae, 63(1):17–39, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2004:KAUb

[468] Richard Holzer. Knowledge acquisition under incomplete knowledge using methods from formal concept analysis: Part II. Fundamenta Informaticae, 63(1):41–63, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna:2004:APB

[469] Bożena Woźna. ACTLS properties and bounded model checking. Fundamenta Informaticae, 63(1):65–87, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2004:EIW

[470] Hsien-Chu Wu and Chin-Chen Chang. Embedding invisible watermarks into digital images based on side-match vector quantization. Fundamenta Informaticae, 63(1):89–106, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2004:P

[471] Anonymous. Preface. Fundamenta Informaticae, 63(2–3):??, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shehory:2004:CFT

[472] Onn Shehory. Coalition formation: Towards feasible solutions. Fundamenta Informaticae, 63(2–3):107–124, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Weyns:2004:FMS

[473] Danny Weyns and Tom Holvoet. A formal model for situated multiagent systems. Fundamenta Informaticae, 63(2–3):125–158, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aldewereld:2004:RTL

[474] Huib Aldewereld, Wiebe van der Hoek, and John-Jules Meyer. Rational teams: Logical aspects of multi-agent systems. Fundamenta Informaticae, 63(2–3):159–183, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jamroga:2004:AKH

[475] Wojciech Jamroga and Wiebe van der Hoek. Agents that know how to play. Fundamenta Informaticae, 63(2–3):185–219, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2004:BUM

[476] M. Kacprzak, A. Lomuscio, and W. Penczek. From bounded to unbounded model checking for temporal epistemic logic. Fundamenta Informaticae, 63(2–3):221–240, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2004:MRN

[477] Antoni Mazurkiewicz. Multilateral ranking negotiations. Fundamenta Informaticae, 63(2–3):241–258, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deHaan:2004:KBA

[478] Hendrik Wietze de Haan, Wim H. Hesselink, and Gerard R. Renardel de Lavalette. Knowledge-based asynchronous programming. Fundamenta Informaticae, 63(2–3):259–281, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dunin-Keplicz:2004:TMC

[479] Barbara Dunin-Kȩplicz and Rineke Verbrugge. A tuning machine for cooperative problem solving. Fundamenta Informaticae, 63(2–3):283– 307, May 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deBoer:2004:TLL

[480] Frank S. de Boer, Maurizio Gabbrielli, and Maria C. Meo. A timed Linda language and its denotational semantics. Fundamenta Informaticae, 63 (4):309–330, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2004:ASRa

[481] Gianpiero Cattaneo, Davide Ciucci, Roberto Giuntini, and Martin Konig. Algebraic structures related to many valued logical systems. Part I: Heyting–Wajsberg algebras. Fundamenta Informaticae, 63(4):331–355, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2004:ASRb

[482] Gianpiero Cattaneo, Davide Ciucci, Roberto Giuntini, and Martin Konig. Algebraic structures related to many valued logical systems. Part II: Equivalence among some widespread structures. Fundamenta Informaticae, 63(4):357–373, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dojer:2004:ATR

[483] Norbert Dojer. Applying term rewriting to partial algebra theory. Fundamenta Informaticae, 63(4):375–384, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sirjani:2004:MVR

[484] Marjan Sirjani, Ali Movaghar, Amin Shali, and Frank S. de Boer. Modeling and verification of reactive systems using Rebeca. Fundamenta Informaticae, 63(4):385–410, June 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:2005:CC

[485] C. Calude, Gh. Păun, and G. Rozenberg. Contagious creativity. Fundamenta Informaticae, 64(1–4):??, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balaban:2005:MCC

[486] Alexandru T. Balaban. Mathematical chemistry: (3, g)-cages with girth g, topological indices, and other graph-theoretical problems. Fundamenta Informaticae, 64(1–4):1–16, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Enguix:2005:BO

[487] Gemma Bel Enguix and Maria Dolores Jiménez López. Biosyntax. An overview. Fundamenta Informaticae, 64(1–4):17–28, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardini:2005:SAP

[488] Francesco Bernardini, Marian Gheorghe, and Vincenzo Manca. On P systems and almost periodicity. Fundamenta Informaticae, 64(1–4):29– 42, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:2005:PCP

[489] Cristian S. Calude and Sergiu Rudeanu. Proving as a computable procedure. Fundamenta Informaticae, 64(1–4):43–52, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Campeanu:2005:RTN

[490] Cezar Câmpeanu, Lila Kari , and Andrei Păun. Results on transforming NFA into DFCA. Fundamenta Informaticae, 64(1–4):53–63, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cavaliere:2005:TSM

[491] Matteo Cavaliere and Dragoş Sburlan. Time and synchronization in membrane systems. Fundamenta Informaticae, 64(1–4):65–77, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ceterchi:2005:TCG

[492] Rodica Ceterchi and Radu Gramatovici. On total contextual grammars with restricted choice. Fundamenta Informaticae, 64(1–4):79–91, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2005:ADG

[493] Gabriel Ciobanu and Mihai Gontineac. An automata description of the genetic message translation. Fundamenta Informaticae, 64(1–4):93–107, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2005:CGS

[494] Jürgen Dassow. Contextual grammars with subregular choice. Fundamenta Informaticae, 64(1–4):109–118, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Davidson:2005:FDC

[495] Michael Davidson and Lucian Ilie. Fast data compression with antidictionaries. Fundamenta Informaticae, 64(1–4):119–134, July 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2005:RDA

[496] Liviu P. Dinu. Rank distance with applications in similarity of natural languages. Fundamenta Informaticae, 64(1–4):135–149, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Domosi:2005:NIL

[497] Pál Dömösi and Manfred Kudlek. New iteration lemmata for regular languages. Fundamenta Informaticae, 64(1–4):151–157, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2005:RRE

[498] Henning Fernau, Rudolf Freund, and Markus Holzer. Representations of recursively enumerable array languages by contextual array grammars. Fundamenta Informaticae, 64(1–4):159–170, July 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gramatovici:2005:PLI

[499] Radu Gramatovici and Florin Manea. Parsing local internal contextual languages with context-free choice. Fundamenta Informaticae, 64(1–4):

171–183, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Head:2005:DLD

[500] Tom Head. Does light direct life toward cosmic awareness? Fundamenta Informaticae, 64(1–4):185–189, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ipate:2005:RFS

[501] Florentin Ipate and Tudor Bălănescu. Refinement in finite state machine testing. Fundamenta Informaticae, 64(1–4):191–203, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ipate:2005:CTS

[502] Florentin Ipate and Mike Holcombe. Complete testing from a stream X-machine specification. Fundamenta Informaticae, 64(1–4):205–216, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurdzinski:2005:DTW

[503] Tomasz Jurdziński, Friedrich Otto, František Mráz, and Martin Plátek. Deterministic two-way restarting automata and Marcus contextual grammars. Fundamenta Informaticae, 64(1–4):217–228, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kappes:2005:RBM

[504] Martin Kappes. On the relationship between multiple brackets and selector languages in bracketed contextual grammars. Fundamenta Informaticae, 64(1–4):229–239, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krishna:2005:FRC

[505] Shankara Narayanan Krishna, R. Rama, and H. Ramesh. Further results on contextual and rewriting P systems. Fundamenta Informaticae, 64 (1–4):241–253, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2005:PCG

[506] Manfred Kudlek. On probabilistic contextual grammars. Fundamenta Informaticae, 64(1–4):255–260, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manca:2005:LGB

[507] Vincenzo Manca. On the logic and geometry of bilinear forms. Fundamenta Informaticae, 64(1–4):261–273, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mardare:2005:LAB

[508] Radu Mardare and Corrado Priami. Logical analysis of biological systems. Fundamenta Informaticae, 64(1–4):275–289, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2005:TVD

[509] Maurice Margenstern, Sergey Verlan, and Yurii Rogozhin. Time-varying distributed H systems: An overview. Fundamenta Informaticae, 64(1– 4):291–306, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitrana:2005:MEC

[510] Victor Mitrana. Marcus external contextual grammars: From one to many dimensions. Fundamenta Informaticae, 64(1–4):307–316, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mutyam:2005:CRE

[511] Madhu Mutyam, Kamala Krithivasan, and A. Siddhartha Reddy. On characterizing recursively enumerable languages by insertion grammars. Fundamenta Informaticae, 64(1–4):317–324, July 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:2005:CCG

[512] Miroslav Novotný. Construction of contextual grammars. Fundamenta Informaticae, 64(1–4):325–340, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okhotin:2005:CGU

[513] Alexander Okhotin and Kai Salomaa. Contextual grammars with uniform sets of trajectories. Fundamenta Informaticae, 64(1–4):341–351, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2005:SAS

[514] Gheorghe Păun, Mario J. Pérez-Jiménez, Juan Pazos, and Alfonso Rodríguez-Patón. Symport/antiport P systems with three objects are universal. Fundamenta Informaticae, 64(1–4):353–367, July 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pawlak:2005:FGI

[515] Zdzisław Pawlak. Flow graphs and intelligent data analysis. Fundamenta Informaticae, 64(1–4):369–377, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2005:RSL

[516] Lech Polkowski and Maria Semeniuk-Polkowska. On rough set logics based on similarity relations. Fundamenta Informaticae, 64(1–4):379– 390, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salomaa:2005:IPM

[517] Arto Salomaa. On the injectivity of Parikh matrix mappings. Fundamenta Informaticae, 64(1–4):391–404, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Scollo:2005:RCP

[518] Giuseppe Scollo. ω-rewriting the Collatz problem. Fundamenta Informaticae, 64(1–4):405–416, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2005:RSV

[519] Andrzej Skowron. Rough sets and vague concepts. Fundamenta Informaticae, 64(1–4):417–431, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tatar:2005:WSD

[520] Doina Tatar. Word sense disambiguation by machine learning approach: A short survey. Fundamenta Informaticae, 64(1–4):433–442, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Titchener:2005:DCE

[521] Mark R. Titchener, Radu Nicolescu, Ludwig Staiger, Aaron Gulliver, and Ulrich Speidel. Deterministic complexity and entropy. Fundamenta Informaticae, 64(1–4):443–461, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tomescu:2005:APF

[522] Ioan Tomescu. Asymptotic properties of the factors of words over a finite alphabet. Fundamenta Informaticae, 64(1–4):463–470, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2005:SCR

[523] Sheng Yu. State complexity: Recent results and open problems. Fundamenta Informaticae, 64(1–4):471–480, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hofmann:2005:TLC

[524] Martin Hofmann and Paweł Urzyczyn. Typed Lambda calculi and applications 2003, selected papers. Fundamenta Informaticae, 65(1–2):??, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abbott:2005:DDD

[525] Michael Abbott, Thorsten Altenkirch, Conor McBride, and Neil Ghani. ∂ for data: Differentiating data structures. Fundamenta Informaticae, 65 (1–2):1–28, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amadio:2005:SMP

[526] Roberto M. Amadio. Synthesis of max-plus quasi-interpretations. Fundamenta Informaticae, 65(1–2):29–60, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blanqui:2005:ITC

[527] Frédéric Blanqui. Inductive types in the calculus of algebraic constructions. Fundamenta Informaticae, 65(1–2):61–86, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Coppola:2005:PTL

[528] Paolo Coppola and Simona Ronchi della Rocca. Principal typing for lambda calculus in elementary affine logic. Fundamenta Informaticae, 65 (1–2):87–112, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Coquand:2005:LFD

[529] Thierry Coquand, Randy Pollack, and Makoto Takeyama. A logical framework with dependently typed records. Fundamenta Informaticae, 65(1–2):113–134, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Joly:2005:DFM

[530] Thierry Joly. On λ-definability I: the fixed model problem and generalizations of the matching problem. Fundamenta Informaticae, 65(1–2): 135–151, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kakutani:2005:PFP

[531] Yoshihiko Kakutani and Masahito Hasegawa. Parameterizations and fixed-point operators on control categories. Fundamenta Informaticae, 65(1–2):153–172, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laird:2005:SBB

[532] Jim Laird. Sequentiality in bounded biorders. Fundamenta Informaticae, 65(1–2):173–191, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doberkat:2005:TRPa

[533] Ernst-Erich Doberkat. Tracing relations probabilistically. Fundamenta Informaticae, 65(3):193–209, September 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrig:2005:BIH

[534] Hartmut Ehrig. Behaviour and instantiation of high-level Petri net processes. Fundamenta Informaticae, 65(3):211–247, September 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mani:2005:SRS

[535] A. Mani. Super rough semantics. Fundamenta Informaticae, 65(3):249– 261, September 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mycka:2005:RRF

[536] Jerzy Mycka. Real recursive functions and Baire classes. Fundamenta Informaticae, 65(3):263–278, September 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Verlan:2005:ASS

[537] Sergey Verlan and Maurice Margenstern. About splicing P systems with one membrane. Fundamenta Informaticae, 65(3):279–290, September 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chan:2005:IHS

[538] Chi-Shiang Chan and Chin-Chen Chang. An image hiding scheme based on multi-bit-reference substitution table using dynamic programming strategy. Fundamenta Informaticae, 65(4):291–305, October 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Georgescu:2005:RPC

[539] Irina Georgescu. Revealed preference, congruence and rationality: A fuzzy approach. Fundamenta Informaticae, 65(4):307–328, October 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2005:CRM

[540] Sanjiang Li, Mingsheng Ying, and Yongming Li. On countable RCC models. Fundamenta Informaticae, 65(4):329–351, October 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lu:2005:DCV

[541] Tzu-Chuen Lu and Chin-Chen Chang. Data compressor for VQ index tables. Fundamenta Informaticae, 65(4):353–371, October 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Richard:2005:RTM

[542] Adrien Richard, Jean-Paul Comet, and Gilles Bernot. R. Thomas’ modeling of biological regulatory networks: Introduction of singular states in the qualitative dynamics. Fundamenta Informaticae, 65(4):373–392, October 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Washio:2005:AMG

[543] Takashi Washio, Luc De Raedt, and Joost N. Kok. Advances in mining graphs, trees and sequences. Fundamenta Informaticae, 66(1–2):??, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferre:2005:DSA

[544] Sébastien Ferré and Ross D. King. A dichotomic search algorithm for mining and learning in domain-specific logics. Fundamenta Informaticae, 66(1–2):1–32, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zaki:2005:EMF

[545] Mohammed J. Zaki. Efficiently mining frequent embedded unordered trees. Fundamenta Informaticae, 66(1–2):33–52, November 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Inokuchi:2005:GFM

[546] Akihiro Inokuchi, Takashi Washio, and Hiroshi Motoda. A general framework for mining frequent subgraphs from labeled graphs. Fundamenta Informaticae, 66(1–2):53–82, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holder:2005:GBR

[547] Lawrence Holder, Diane Cook, Jeff Coble, and Maitrayee Mukherjee. Graph-based relational learning with application to security. Fundamenta Informaticae, 66(1–2):83–101, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Habrard:2005:DIS

[548] Amaury Habrard, Marc Bernard, and Marc Sebban. Detecting irrelevant subtrees to improve probabilistic learning from tree-structured data. Fundamenta Informaticae, 66(1–2):103–130, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geamsakul:2005:CDT

[549] Warodom Geamsakul, Tetsuya Yoshida, Kouzou Ohara, Hiroshi Motoda, Hideto Yokoi, and Katsuhiko Takabayashi. Constructing a decision tree for graph-structured data and its applications. Fundamenta Informaticae, 66(1–2):131–160, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chi:2005:FSM

[550] Yun Chi, Richard R. Muntz, Siegfried Nijssen, and Joost N. Kok. Frequent subtree mining — an overview. Fundamenta Informaticae, 66(1–2): 161–198, November 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Borzyszkowski:2005:GIF

[551] Tomasz Borzyszkowski. Generalized interpolation in first order logic. Fundamenta Informaticae, 66(3):199–219, December 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2005:SAC

[552] Raymond Devillers and Hanna Klaudel. Synchronous and asynchronous communications in composable parameterized high-level Petri nets. Fundamenta Informaticae, 66(3):221–257, December 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doberkat:2005:TRPb

[553] Ernst-Erich Doberkat. Tracing relations probabilistically. Fundamenta Informaticae, 66(3):259–275, December 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2005:WCH

[554] Olivier Finkel. On winning conditions of high Borel complexity in pushdown games. Fundamenta Informaticae, 66(3):277–298, December 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkova:2005:EMD

[555] Albina Moshkova. On effective methods for diagnosis of retaining faults in circuits. Fundamenta Informaticae, 66(3):299–313, December 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pettorossi:2005:PTT

[556] Alberto Pettorossi and Maurizio Proietti. Program transformation: Theoretical foundations and basic techniques. Part 1. Fundamenta Informaticae, 66(4):??, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cunha:2005:PFP

[557] Alcino Cunha and Jorge Sousa Pinto. Point-free program transformation. Fundamenta Informaticae, 66(4):315–352, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gibbons:2005:PMC

[558] Jeremy Gibbons and Graham Hutton. Proof methods for corecursive programs. Fundamenta Informaticae, 66(4):353–366, January 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gluck:2005:MAP

[559] Robert Glück and Masahiko Kawabe. A method for automatic program inversion based on LR(0) parsing. Fundamenta Informaticae, 66(4):367– 395, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Danvy:2005:TBA

[560] Olivier Danvy and Mayer Goldberg. There and back again. Fundamenta Informaticae, 66(4):397–413, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lisper:2005:IUT

[561] Bjórn Lisper. Infinite unfolding and transformations of nondeterministic programs. Fundamenta Informaticae, 66(4):415–439, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2005:CSP

[562] Hans-Dieter Burkhard, Ludwik Czaja, Gabriela Lindemann, and Zbigniew Suraj. Concurrency specification and programming. Fundamenta Informaticae, 67(1–3):??, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2005:AIO

[563] Roberto Barbuti and Stefano Cataudella. Abstract interpretation of an object calculus for synchronization optimizations. Fundamenta Informaticae, 67(1–3):1–12, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2005:PMM

[564] Roberto Barbuti, Stefano Cataudella, Andrea Maggiolo-Schettini, Paolo Milazzo, and Angelo Troina. A probabilistic model for molecular systems. Fundamenta Informaticae, 67(1–3):13–27, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2005:HOP

[565] Marco Bellia and M. Eugenia Occhiuto. Higher order programming in Java: Introspection, subsumption and extraction. Fundamenta Informaticae, 67(1–3):29–44, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2005:AFP

[566] Ludwik Czaja. An axiomatic framework for proving correctness of nets. Fundamenta Informaticae, 67(1–3):45–64, January 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2005:PNP

[567] Berndt Farwer and Michael Köhler. Petri net processes for zero-safe nets. Fundamenta Informaticae, 67(1–3):65–75, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2005:SMF

[568] Anna Gomolińska. Satisfiability and meaning of formulas and sets of formulas in approximation spaces. Fundamenta Informaticae, 67(1–3): 77–92, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klunder:2005:SCF

[569] Barbara Klunder, Edward Ochmański, and Krystyna Stawikowska. On star-connected flat languages. Fundamenta Informaticae, 67(1–3):93– 105, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryvyy:2005:AMA

[570] Sergiy Kryvyy, Lyudmila Matvyeyeva, and Mariya Lopatina. Automatic modeling and analysis of MSC-specified systems. Fundamenta Informaticae, 67(1–3):107–120, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2005:PPN

[571] Manfred Kudlek. Probability in Petri nets. Fundamenta Informaticae, 67(1–3):121–130, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Latkowski:2005:FIR

[572] Rafał Latkowski. Flexible indiscernibility relations for missing attribute values. Fundamenta Informaticae, 67(1–3):131–147, January 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popova-Zeugmann:2005:TPN

[573] Louchka Popova-Zeugmann, Monika Heiner, and Ina Koch. Time Petri nets for modelling and analysis of biochemical networks. Fundamenta Informaticae, 67(1–3):149–162, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popova-Zeugmann:2005:ERS

[574] Louchka Popova-Zeugmann and Matthias Werner. Extreme runtimes of schedules modelled by time Petri nets. Fundamenta Informaticae, 67(1– 3):163–174, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2005:AOP

[575] Roman R. Redziejowski. Associative Omega-products of traces. Fundamenta Informaticae, 67(1–3):175–185, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schroter:2005:SFN

[576] Kay Schröter, Diemo Urbig, and Nora Hans. Social formation of negotiation space and groups for non-isolated multilateral negotiations. Fundamenta Informaticae, 67(1–3):187–201, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:2005:MCP

[577] Jarosław Stepaniuk, Jan G. Bazan, and Andrzej Skowron. Modelling complex patterns by information systems. Fundamenta Informaticae, 67(1–3):203–217, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2005:CDK

[578] Zbigniew Suraj, James F. Peters, and Piotr Grochowalski. A controller design for the Khepera robot: A rough set approach. Fundamenta Informaticae, 67(1–3):219–231, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2005:RBC

[579] Zbigniew Suraj and Krzysztof Pancerz. Restriction-based concurrent system design using the rough set formalism. Fundamenta Informaticae, 67(1–3):233–247, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Synak:2005:STA

[580] Piotr Synak, Jan G. Bazan, Andrzej Skowron, and James F. Peters. Spatio-temporal approximate reasoning over complex objects. Fundamenta Informaticae, 67(1–3):249–269, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urbig:2005:WBN

[581] Diemo Urbig. Weight-based negotiation mechanisms: Balancing personal utilities. Fundamenta Informaticae, 67(1–3):271–285, January 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wroblewski:2005:USL

[582] Dobiesław Wróblewski. Universal semi-local election protocol using forward links. Fundamenta Informaticae, 67(1–3):287–301, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2005:SBR

[583] Andrzej Zbrzezny. SAT-based reachability checking for timed automata with diagonal constraints. Fundamenta Informaticae, 67(1–3):303–322, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gire:2005:NPS

[584] Françoise Gire and Corinne Plourde. A new partial semantics for disjunctive deductive databases. Fundamenta Informaticae, 67(4):323–342, March 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2005:TET

[585] Ryszard Janicki and Alan Wassyng. Tabular expressions and their relational semantics. Fundamenta Informaticae, 67(4):343–370, March 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popa:2005:KTD

[586] Ionuţ Popa. KDS-transformation for data compression. Fundamenta Informaticae, 67(4):371–375, March 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xu:2005:FSO

[587] Jian Xu and Jeffery Zucker. First and second order recursion on abstract data types. Fundamenta Informaticae, 67(4):377–419, March 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boryczka:2005:EIA

[588] Mariusz Boryczka. Eliminating introns in ant colony programming. Fundamenta Informaticae, 68(1–2):1–19, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burrieza:2005:MLA

[589] Alfredo Burrieza and Manuel Ojeda-Aciego. A multimodal logic approach to order of magnitude qualitative reasoning with comparability and negligibility relations. Fundamenta Informaticae, 68(1–2):21–46, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eberbach:2005:CBR

[590] Eugene Eberbach. $-calculus of bounded rational agents: Flexible optimization as search under bounded resources in interactive systems. Fundamenta Informaticae, 68(1–2):47–102, January 2005. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fecher:2005:ESA

[591] Harald Fecher and Mila Majster-Cederbaum. Event structures for arbitrary disruption. Fundamenta Informaticae, 68(1–2):103–130, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stehr:2005:OCCa

[592] Mark-Oliver Stehr. The open calculus of constructions (part i): An equational type theory with dependent types for programming, specification, and interactive theorem proving. Fundamenta Informaticae, 68(1–2): 131–174, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2005:TFM

[593] Józef Winkowski. Towards a framework for modelling systems with rich internal structures of states and processes. Fundamenta Informaticae, 68(1–2):175–206, January 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2005:IHT

[594] Chin-Chen Chang and Chi-Shiang Chan. An inverse halftoning technique using modified look-up tables. Fundamenta Informaticae, 68(3): 207–220, May 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Palka:2005:FMP

[595] Ewa Palka. On finite model property of the equational theory of Kleene algebras. Fundamenta Informaticae, 68(3):221–230, May 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stay:2005:VSC

[596] Michael Stay. Very simple Chaitin machines for concrete AIT. Fundamenta Informaticae, 68(3):231–247, May 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stehr:2005:OCCb

[597] Mark-Oliver Stehr. The open calculus of constructions (Part II): An equational type theory with dependent types for programming, specification, and interactive theorem proving. Fundamenta Informaticae, 68 (3):249–288, May 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2005:CSA

[598] Guoyin Y. Wang, Jun Zhao, Jiujiang An, and Yu Wu. A comparative study of algebra viewpoint and information viewpoint in attribute reduction. Fundamenta Informaticae, 68(3):289–301, May 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bocharov:2005:IDC

[599] Igor Bocharov and Pavel Lukin. Information divergence criterion in speech signal recognition. Fundamenta Informaticae, 68(4):303–313, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2005:NDB

[600] Chih-Yang Lin, Chin-Chen Chang, and Chia-Chen Lin. A new densitybased scheme for clustering based on genetic algorithm. Fundamenta Informaticae, 68(4):315–331, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Poranen:2005:GHA

[601] Timo Poranen. A genetic hillclimbing algorithm for the optimal linear arrangement problem. Fundamenta Informaticae, 68(4):333–356, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qi:2005:ICL

[602] Guilin Qi, Peter Milligan, and Paul Sage. Incidence calculus on Łukasiewicz’s three-valued logic. Fundamenta Informaticae, 68(4):357– 378, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sydow:2005:RSB

[603] Marcin Sydow. Random surfer with back step. Fundamenta Informaticae, 68(4):379–398, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2005:SSP

[604] Hsien-Chu Wu, Min-Shiang Hwang, and Chia-Hsin Liu. A secure strongpassword authentication protocol. Fundamenta Informaticae, 68(4):399– 406, June 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pettorossi:2006:PTT

[605] Alberto Pettorossi and Maurizio Proietti. Program transformation: Theoretical foundations and basic techniques. Part 2. Fundamenta Informaticae, 69(1–2):??, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chin:2006:RCE

[606] Wei-Ngan Chin, Siau-Cheng Khoo, and Neil Jones. Redundant call elimination via tupling. Fundamenta Informaticae, 69(1–2):1–37, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamilton:2006:HOD

[607] Geoff W. Hamilton. Higher order deforestation. Fundamenta Informaticae, 69(1–2):39–61, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Johann:2006:ISF

[608] Patricia Johann and Janis Voigtländer. The impact of seq on free theorems-based program transformations. Fundamenta Informaticae, 69 (1–2):63–102, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2006:ITP

[609] Chiyan Chen, Rui Shi, and Hongwei Xi. Implementing typeful program transformations. Fundamenta Informaticae, 69(1–2):103–121, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bravenboer:2006:PTS

[610] Martin Bravenboer, Arthur van Dam, Karina Olmos, and Eelco Visser. Program transformation with scoped dynamic rewrite rules. Fundamenta Informaticae, 69(1–2):123–178, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

King:2006:CGP

[611] Andy King and Jonathan C. Martin. Control generation by program transformation. Fundamenta Informaticae, 69(1–2):179–218, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rosenblueth:2006:MCF

[612] David A. Rosenblueth. A multiple-clause folding rule using instantiation and generalization. Fundamenta Informaticae, 69(1–2):219–249, January 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bandini:2006:SAD

[613] Stefania Bandini, Giancarlo Mauri, and Giuseppe Vizzari. Supporting action-at-a-distance in situated cellular agents. Fundamenta Informaticae, 69(3):251–271, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dereniowski:2006:EPQ

[614] Dariusz Dereniowski and Marek Kubale. Efficient parallel query processing by graph ranking. Fundamenta Informaticae, 69(3):273–285, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gerla:2006:SMA

[615] B. Gerla and I. Leuştean. Similarity MV-algebras. Fundamenta Informaticae, 69(3):287–300, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:2006:REL

[616] Michael Kaminski and Tony Tan. Regular expressions for languages over infinite alphabets. Fundamenta Informaticae, 69(3):301–318, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2006:IRS

[617] Ching-Lin Wang, Ren-Hung Hwang, Yung-Kuan Chan, Chih-Ya Chen, and Chuan-Chung Cheng. An image retrieval system based on the color, areas, and perimeters of objects. Fundamenta Informaticae, 69(3):319– 330, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2006:ARS

[618] Gui-Long Liu. The axiomatization of the rough set upper approximation operations. Fundamenta Informaticae, 69(3):331–342, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2006:NMP

[619] Zbigniew Suraj and Pawel Delimata. On k-NN method with preprocessing. Fundamenta Informaticae, 69(3):343–358, February 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2006:ARS

[620] Davide Ciucci. On the axioms of residuated structures: Independence, dependencies and rough approximations. Fundamenta Informaticae, 69 (4):359–387, March 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esposito:2006:MOR

[621] Floriana Esposito, Nicola Fanizzi, Stefano Ferilli, Teresa M. A. Basile, and Nicola Di Mauro. Multistrategy operators for relational learning and their cooperation. Fundamenta Informaticae, 69(4):389–409, March 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pomykala:2006:ECB

[622] Jacek Pomykała and Sławomir Barabasz. Eliptic curve based threshold proxy signature scheme with known signers. Fundamenta Informaticae, 69(4):411–425, March 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See security analysis [945].

Xia:2006:MMR

[623] Lirong Xia and Sanjiang Li. On minimal models of the region connection calculus. Fundamenta Informaticae, 69(4):427–446, March 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2006:SIA

[624] Ryszard Janicki. Special issueon application of concurrency to system design. Fundamenta Informaticae, 70(1–2):??, April 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bjorklund:2006:RBE

[625] Dag Björklund and Johan Lilius. Rialto to B: An exercise in formal development of a language for multiple models of computation. Fundamenta Informaticae, 70(1–2):1–20, April 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kapoor:2006:VID

[626] Hemangee K. Kapoor, Mark B. Josephs, and Dennis P. Furey. Verification and implementation of delay-insensitive processes in restrictive environments. Fundamenta Informaticae, 70(1–2):21–48, April 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khomenko:2006:LSA

[627] Victor Khomenko, Maciej Koutny, and Alex Yakovlev. Logic synthesis for asynchronous circuits based on STG unfoldings and incremental SAT. Fundamenta Informaticae, 70(1–2):49–73, April 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lawford:2006:TIV

[628] Mark Lawford, Vera Pantelic, and Hong Zhang. Towards integrated verification of timed transition models. Fundamenta Informaticae, 70 (1–2):75–110, April 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seppi:2006:GMC

[629] Kevin Seppi, Michael Jones, and Peter Lamborn. Guided model checking with a Bayesian meta-heuristic. Fundamenta Informaticae, 70(1–2):111– 126, April 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tauriainen:2006:NES

[630] Heikki Tauriainen. Nested emptiness search for generalized Büchi automata. Fundamenta Informaticae, 70(1–2):127–154, April 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xia:2006:BAC

[631] Fei Xia, Fei Hao, Ian Clark, Alex Yakovlev, and E. Graeme Chester. Buffered asynchronous communication mechanisms. Fundamenta Informaticae, 70(1–2):155–170, April 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2006:SDI

[632] Chin-Chen Chang, Jun-Chou Chuang, and Yu-Chen Hu. Spatial domain image hiding scheme using pixel-values differencing. Fundamenta Informaticae, 70(3):171–184, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2006:PHS

[633] Chin-Chen Chang, Chih-Yang Lin, and Henry Chou. Perfect hashing schemes for mining traversal patterns. Fundamenta Informaticae, 70(3): 185–202, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2006:PNS

[634] Raymond Devillers, Hanna Klaudel, and Maciej Koutny. Petri net semantics of the finite π-calculus terms. Fundamenta Informaticae, 70(3): 203–226, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasquet:2006:ICM

[635] Olivier Gasquet. On the influence of confluence in modal logics. Fundamenta Informaticae, 70(3):227–250, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2006:ITS

[636] Yu-Chen Hu and Chin-Chen Chang. An improved tree-structured codebook search algorithm for grayscale image compression. Fundamenta Informaticae, 70(3):251–260, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jaworski:2006:BV

[637] Wojciech Jaworski. Bounds for validation. Fundamenta Informaticae, 70(3):261–275, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sali:2006:CFK

[638] Attila Sali and Klaus-Dieter Schewe. Counter-free keys and functional dependencies in higher-order datamodels. Fundamenta Informaticae, 70 (3):277–301, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berghammer:2006:ECM

[639] Rudolf Berghammer and Alexander Fronk. Exact computation of minimum feedback vertex sets with relational algebra. Fundamenta Informaticae, 70(4):301–316, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2006:NCI

[640] Yu-Chen Hu, Min-Hui Lin, and Ji-Han Jiang. A novel color image hiding scheme using block truncation coding. Fundamenta Informaticae, 70(4): 317–331, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mani:2006:ASL

[641] A. Mani. Di-algebraic semantics of logics. Fundamenta Informaticae, 70 (4):333–350, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2006:NOH

[642] Linh Anh Nguyen. Negative ordered hyper-resolution as a proof procedure for disjunctive logic programming. Fundamenta Informaticae, 70 (4):351–366, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pestov:2006:ISS

[643] Vladimir Pestov and Aleksandar Stojmirović. Indexing schemes for similarity search: an illustrated paradigm. Fundamenta Informaticae, 70 (4):367–385, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yen:2006:DAS

[644] Hsu-Chun Yen and Lien-Po Yu. Decidability analysis of self-stabilization for infinite-state systems. Fundamenta Informaticae, 70(4):387–402, June 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2006:P

[645] Anonymous. Preface. Fundamenta Informaticae, 71(1):??, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Flasinski:2006:AFS

[646] Mariusz Flasiński and Janusz Jurek. On the analysis of fuzzy string patterns with the help of extended and stochastic GDPLL(k) grammars. Fundamenta Informaticae, 71(1):1–14, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grochowski:2006:APS

[647] Marek Grochowski, Robert Schaefer, and Maciej Smołka. Architectural principles and scheduling strategies for computing agent systems. Fundamenta Informaticae, 71(1):15–26, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kozlak:2006:MAA

[648] Jarosław Koźlak, Jean-Charles Créput, Vincent Hilaire, and Abder Koukam. Multi-agent approach to dynamic pick-up and delivery problem with uncertain knowledge about future transport demands. Fundamenta Informaticae, 71(1):27–36, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nawarecki:2006:ADC

[649] Edward Nawarecki, Marek Kisiel-Dorohinicki, and Grzegorz Dobrowolski. Architecture for discovery of crises in MAS. Fundamenta Informaticae, 71(1):37–47, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2006:STR

[650] Lech Polkowski. A set theory for rough sets. Toward a formal calculus of vague statements. Fundamenta Informaticae, 71(1):49–61, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stefanowski:2006:ESU

[651] Jerzy Stefanowski. An empirical study of using rule induction and rough sets to software cost estimation. Fundamenta Informaticae, 71(1):63– 82, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2006:TAP

[652] Zbigniew Suraj and Barbara Fryc. Timed approximate Petri nets. Fundamenta Informaticae, 71(1):83–99, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2006:PNS

[653] Zbigniew Suraj, Barbara Fryc, Zofia Matusiewicz, and Krzysztof Pancerz. A Petri net system — an overview. Fundamenta Informaticae, 71(1):101–120, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2006:RCS

[654] Zbigniew Suraj and Krzysztof Pancerz. Reconstruction of concurrent system models described by decomposed data tables. Fundamenta Informaticae, 71(1):121–137, July 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Banerjee:2006:LRT

[655] Mohua Banerjee. Logic for rough truth. Fundamenta Informaticae, 71(2– 3):139–151, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2006:TAM

[656] J. A. Bergstra and C. A. Middelburg. Thread algebra with multi-level strategies. Fundamenta Informaticae, 71(2–3):153–182, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calimeri:2006:POD

[657] Francesco Calimeri, Wolfgang Faber, Gerald Pfeifer, and Nicola Leone. Pruning operators for disjunctive logic programming systems. Fundamenta Informaticae, 71(2–3):183–214, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2006:NIC

[658] Chin-Chen Chang, Guei-Mei Chen, and Yu-Chen Hu. A novel index coding scheme for vector quantization. Fundamenta Informaticae, 71(2– 3):215–227, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czyzowicz:2006:CSB

[659] Jurek Czyzowicz, Dariusz Kowalski, Euripides Markou, and Andrzej Pelc. Complexity of searching for a black hole. Fundamenta Informaticae, 71(2–3):229–242, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2006:RCW

[660] Kees van Hee, Natalia Sidorova, and Marc Voorhoeve. Resourceconstrained workflow nets. Fundamenta Informaticae, 71(2–3):243–257, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Iliopoulos:2006:WST

[661] Costas S. Iliopoulos, Christos Makris, Yannis Panagis, Katerina Perdikuri, Evangelos Theodoridis, and Athanasios Tsakalidis. The weighted suffix tree: An efficient data structure for handling molecular weighted sequences and its applications. Fundamenta Informaticae, 71(2–3):259–277, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ionescu:2006:SNS

[662] Mihai Ionescu, Gheorghe Păun, and Takashi Yokomori. Spiking neural P systems. Fundamenta Informaticae, 71(2–3):279–308, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeLoof:2006:ELI

[663] Karel De Loof, Hans De Meyer, and Bernard De Baets. Exploiting the lattice of ideals representation of a poset. Fundamenta Informaticae, 71(2–3):309–321, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2006:RLA

[664] James F. Peters and Christopher Henry. Reinforcement learning with approximation spaces. Fundamenta Informaticae, 71(2–3):323–349, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2006:RCB

[665] Ming-Ni Wu, Chia-Chen Lin, and Chin-Chen Chang. A robust contentbased copy detection scheme. Fundamenta Informaticae, 71(2–3):351– 366, August 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chaitin:2006:PPS

[666] Gregory Chaitin. Probability and program-size for functions. Fundamenta Informaticae, 71(4):367–370, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Christiansen:2006:SDI

[667] Henning Christiansen and Davide Martinenghi. On simplification of database integrity constraints. Fundamenta Informaticae, 71(4):371–417, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamide:2006:FPR

[668] Norihiro Kamide. Foundations of paraconsistent resolution. Fundamenta Informaticae, 71(4):419–441, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2006:NIO

[669] Chia-Chen Lin, Yu Chen Hu, and Chin-Chen Chang. A novel image ownership protection scheme based on rehashing concept and vector quantization. Fundamenta Informaticae, 71(4):443–451, September 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2006:FLA

[670] L. Kari, E. Losseva, S. Konstantinidis, P. Sosík, and G. Thierrin. A formal language analysis of DNA hairpin structures. Fundamenta Informaticae, 71(4):453–475, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winter:2006:PPO

[671] Michael Winter. On problems in polymorphic object-oriented languages with self types and matching. Fundamenta Informaticae, 71(4):477–491, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhong:2006:ESC

[672] Sheng Zhong. An efficient and secure cryptosystem for encrypting long messages. Fundamenta Informaticae, 71(4):493–497, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2006:CSP

[673] Anonymous. Concurrency specification and programming (CS&P). Fundamenta Informaticae, 72(1–3):??, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andreeva:2006:OET

[674] Maria V. Andreeva and Irina B. Virbitskaite. Observational equivalences for timed stable event structures. Fundamenta Informaticae, 72(1–3): 1–19, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2006:CLS

[675] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Angelo Troina. A calculus of looping sequences for modelling microbiological systems. Fundamenta Informaticae, 72(1–3):21–35, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2006:BPI

[676] Jan G. Bazan. Behavioral pattern identification through rough set modeling. Fundamenta Informaticae, 72(1–3):37–50, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bednarczyk:2006:EVT

[677] Marek A. Bednarczyk, Wojciech Jamroga, and Wiesław Pawłowski. Expressing and verifying temporal and structural properties of mobile agents. Fundamenta Informaticae, 72(1–3):51–63, September 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chrzastowski-Wachtel:2006:DSM

[678] Piotr Chrzastowski-Wachtel. Determining sound markings in structured nets. Fundamenta Informaticae, 72(1–3):65–79, September 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2006:EMP

[679] Ludwik Czaja. Equations for message passing. Fundamenta Informaticae, 72(1–3):81–93, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doroshenko:2006:RFR

[680] Anatoliy Doroshenko and Ruslan Shevchenko. A rewriting framework for rule-based programming dynamic applications. Fundamenta Informaticae, 72(1–3):95–108, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2006:MGL

[681] Berndt Farwer and Michael Köhler. Modelling global and local name spaces for mobile agents using object nets. Fundamenta Informaticae, 72(1–3):109–122, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2006:SCD

[682] Berndt Farwer and Mauricio Varea. Separation of control and data flow in high-level Petri nets: Transforming dual flow nets into object Petri nets. Fundamenta Informaticae, 72(1–3):123–137, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2006:PRI

[683] Anna Gomolińska. Possible rough ingredients of concepts in approximation spaces. Fundamenta Informaticae, 72(1–3):139–154, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2006:DBS

[684] Franciszek Grabowski and Dominik Strzalka. Dynamic behavior of simple insertion sort algorithm. Fundamenta Informaticae, 72(1–3):155–165, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2006:NIF

[685] Damas P. Gruska. Network information flow. Fundamenta Informaticae, 72(1–3):167–180, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janowska:2006:STA

[686] Agata Janowska and Paweł Janowski. Slicing of timed automata with discrete data. Fundamenta Informaticae, 72(1–3):181–195, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2006:SCR

[687] Magdalena Kacprzak. A strong completeness result for a MAS logic. Fundamenta Informaticae, 72(1–3):197–213, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2006:CBS

[688] Magdalena Kacprzak, Alessio Lomuscio, Artur Niewiadomski, Wojciech Penczek, Franco Raimondi, and Maciej Szreter. Comparing BDD and SAT based techniques for model checking Chaum’s dining cryptographers protocol. Fundamenta Informaticae, 72(1–3):215–234, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klunder:2006:SCF

[689] Barbara Klunder. Star-connected flat languages and automata. Fundamenta Informaticae, 72(1–3):235–243, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler:2006:PSD

[690] Michael Köhler and Heiko Rölke. Properties of super-dual nets. Fundamenta Informaticae, 72(1–3):245–254, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2006:QVA

[691] Manfred Kudlek. On quantum vector addition systems. Fundamenta Informaticae, 72(1–3):255–261, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurkowski:2006:QFF

[692] Mirosław Kurkowski and Marian Srebrny. A quantifier-free first-order knowledge logic of authentication. Fundamenta Informaticae, 72(1–3): 263–282, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nowak:2006:OSR

[693] Agnieszka Nowak, Alicja Wakulicz-Deja, and Sebastian Bachliński. Optimization of speech recognition by clustering of phones. Fundamenta Informaticae, 72(1–3):283–293, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2006:LSC

[694] Trung Thanh Nguyen, Claire P. Willis, Derek J. Paddon, Sinh Hoa Nguyen, and Hung Son Nguyen. Learning sunspot classification. Fundamenta Informaticae, 72(1–3):295–309, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ochmanski:2006:TNC

[695] Edward Ochmański and Joanna Pieckowska. Trace nets and conflict-free computations. Fundamenta Informaticae, 72(1–3):311–321, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ochmanski:2006:SFS

[696] Edward Ochmański and Krystyna Stawikowska. Star-free star and trace languages. Fundamenta Informaticae, 72(1–3):323–331, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2006:AOP

[697] Roman R. Redziejowski. Associative Omega-product of processes. Fundamenta Informaticae, 72(1–3):333–345, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shilov:2006:UAM

[698] N. V. Shilov, N. O. Garanina, and K.-M. Choe. Update and abstraction in model checking of knowledge and branching time. Fundamenta Informaticae, 72(1–3):347–361, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2006:CAS

[699] Andrzej Skowron, Jarosław Stepaniuk, James Peters, and Roman Swiniarski. Calculi of approximation spaces. Fundamenta Informaticae, 72(1–3):363–378, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stefanowski:2006:RSH

[700] Jerzy Stefanowski and Szymon Wilk. Rough sets for handling imbalanced data: Combining filtering and rule-based classifiers. Fundamenta Informaticae, 72(1–3):379–391, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2006:RSA

[701] Zbigniew Suraj, Neamat El Gayar, and Pawel Delimata. A rough set approach to multiple classifier systems. Fundamenta Informaticae, 72 (1–3):393–406, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2006:ACA

[702] Józef Winkowski. An axiomatic characterization of algebras of processes of Petri nets. Fundamenta Informaticae, 72(1–3):407–420, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2006:COC

[703] Marcin Wolski. Complete orders, categories and lattices of approximations. Fundamenta Informaticae, 72(1–3):421–435, September 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Collings:2006:SPV

[704] Simon Collings, Ryszard Kozera, and Lyle Noakes. Surface properties from N-views of a strictly convex solid. Fundamenta Informaticae, 72(4): 437–452, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fredriksson:2006:LAS

[705] Kimmo Fredriksson. On-line approximate string matching in natural language. Fundamenta Informaticae, 72(4):453–466, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-

8681 (electronic). URL http://content.iospress.com/articles/
fundamenta-informaticae/fi72-4-02;   http://www.cs.uku.fi/~
fredriks/pub/papers/fi06.pdf.

Paun:2006:SPM

[706] Andrei Păun and Bianca Popa. P systems with proteins on membranes. Fundamenta Informaticae, 72(4):467–483, September 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Potikas:2006:VPT

[707] Petros Potikas, Panos Rondogiannis, and Manolis Gergatsoulis. A valuepropagating transformation technique for Datalog programs based on non-deterministic constructs. Fundamenta Informaticae, 72(4):485–527, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shen:2006:MBF

[708] Haifeng Shen, Qunxia Li, Jun Guo, and Gang Liu. Model-based feature compensation for robust speech recognition. Fundamenta Informaticae, 72(4):529–539, September 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2006:TLT

[709] Anonymous. Trajectories of language theory. Fundamenta Informaticae, 73(1–2):??, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:2006:ARN

[710] Cristian S. Calude, Cezar Câmpeanu, and Monica Dumitrescu. Automata recognizing no words: A statistical approach. Fundamenta Informaticae, 73(1–2):1–18, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ceterchi:2006:ACS

[711] Rodica Ceterchi. An algebraic characterization of semi-simple splicing. Fundamenta Informaticae, 73(1–2):19–25, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2006:BCS

[712] Gabriel Ciobanu, Gheorghe Păun, and Mario J. Pérez-Jiménez. On the branching complexity of P systems. Fundamenta Informaticae, 73(1– 2):27–36, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2006:CBC

[713] Erzsébet Csuhaj-Varjú and Alexandru Mateescu. On a connection between cooperating distributed grammar systems and basic process algebra. Fundamenta Informaticae, 73(1–2):37–50, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ding:2006:SSS

[714] Cunsheng Ding and Arto Salomaa. Secret sharing schemes with nice access structures. Fundamenta Informaticae, 73(1–2):51–63, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ding:2006:SPM

[715] Cunsheng Ding and Arto Salomaa. On some problems of mateescu concerning subword occurrences. Fundamenta Informaticae, 73(1–2):65–79, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Domaratzki:2006:IT

[716] Michael Domaratzki, Grzegorz Rozenberg, and Kai Salomaa. Interpreted trajectories. Fundamenta Informaticae, 73(1–2):81–97, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Droste:2006:OSP

[717] Manfred Droste, Jarkko Kari, and Paula Steinby. Observations on the smoothness properties of real functions computed by weighted finite automata. Fundamenta Informaticae, 73(1–2):99–106, October 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dumitrescu:2006:FAD

[718] Sorina Dumitrescu. Faster algorithm for designing optimal prefix-free codes with unequal letter costs. Fundamenta Informaticae, 73(1–2):107– 117, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:2006:URL

[719] Vesa Halava, Tero Harju, and Juhani Karhumäki. Undecidability in ω-regular languages. Fundamenta Informaticae, 73(1–2):119–125, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:2006:PWI

[720] Juha Honkala. On the problem whether the image of an N-rational series equals. Fundamenta Informaticae, 73(1–2):127–132, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ibarra:2006:CPD

[721] Oscar H. Ibarra, Sara Woodworth, Hsu-Chun Yen, and Zhe Dang. On the computational power of 1-deterministic and sequential P systems. Fundamenta Informaticae, 73(1–2):133–152, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilie:2006:SCS

[722] Lucian Ilie and Cristian Popescu. The shortest common superstring problem and viral genome compression. Fundamenta Informaticae, 73(1– 2):153–164, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2006:BST

[723] Lila Kari and Elena Losseva. Block substitutions and their properties. Fundamenta Informaticae, 73(1–2):165–178, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Langille:2006:SGA

[724] Miika Langille and Ion Petre. Simple gene assembly is deterministic. Fundamenta Informaticae, 73(1–2):179–190, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manea:2006:SSB

[725] Florin Manea, Victor Mitrana, and Daniel-Claudian Voinescu. Synchronized shuffle on backbones. Fundamenta Informaticae, 73(1–2):191–202, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcus:2006:MCS

[726] Solomon Marcus. Mild context-sensitivity, after twenty years. Fundamenta Informaticae, 73(1–2):203–204, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mateescu:2006:RPM

[727] Alexandru Mateescu and Rudolf Freund. Routes and products of monoids. Fundamenta Informaticae, 73(1–2):205–211, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2006:MCE

[728] Gheorghe Păun and Radu Păun. Membrane computing and economics: Numerical P systems. Fundamenta Informaticae, 73(1–2):213–227, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paz:2006:PIR

[729] Azaria Paz. A property of independency relations induced by probabilistic distributions with binary variables. Fundamenta Informaticae, 73(1–2):229–236, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Santean:2006:NBR

[730] Nicolae Santean and Sheng Yu. Nondeterministic bimachines and rational relations with finite codomain. Fundamenta Informaticae, 73(1–2): 237–264, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Serbanuta:2006:IPM

[731] Virgil Nicolae Şerbănuţă and Traian Florin Şerbănuţă. Injectivity of the Parikh matrix mappings revisited. Fundamenta Informaticae, 73(1–2): 265–283, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stefanescu:2006:ISR

[732] Gheorghe Stefanescu. Interactive systems with registers and voices. Fundamenta Informaticae, 73(1–2):285–305, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vaida:2006:NSO

[733] Dragoş Vaida. Note on some order properties related to processes semantics (i). Fundamenta Informaticae, 73(1–2):307–319, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Analyti:2006:LPR

[734] Anastasia Analyti and Ioannis Pachoulakis. Logic programming representation of the compound term composition algebra. Fundamenta Informaticae, 73(3):321–360, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2006:LCD

[735] Liviu P. Dinu and Andrea Sgarro. A low-complexity distance for DNA strings. Fundamenta Informaticae, 73(3):361–372, October 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2006:MIE

[736] Yu-Chen Hu. Multiple images embedding scheme based on moment preserving block truncation coding. Fundamenta Informaticae, 73(3):373– 387, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2006:AMP

[737] Weiru Liu, Guilin Qi, and David A. Bell. Adaptive merging of prioritized knowledge bases. Fundamenta Informaticae, 73(3):389–407, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pinna:2006:ESD

[738] G. Michele Pinna. Event structures with disabling/enabling relation and event automata. Fundamenta Informaticae, 73(3):409–430, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brogi:2006:FCL

[739] Antonio Brogi, Jean-Marie Jacquet, and Joost Kok. Foundations on coordination languages and software architectures. Fundamenta Informaticae, 73(4):??, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brogi:2006:FAS

[740] Antonio Brogi, Jean-Marie Jacquet, and Isabelle Linden. Fully abstract semantics for a coordination model with asynchronous communication and enhanced matching. Fundamenta Informaticae, 73(4):431–478, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorrieri:2006:SSC

[741] Roberto Gorrieri, Roberto Lucchi, and Gianluigi Zavattaro. Supporting secure coordination in SecSpaces. Fundamenta Informaticae, 73(4):479– 506, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Viroli:2006:CS

[742] Mirko Viroli and Andrea Omicini. Coordination as a service. Fundamenta Informaticae, 73(4):507–534, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orzan:2006:DSS

[743] Simona Orzan and Jaco van de Pol. Distribution of a simple shared dataspace architecture. Fundamenta Informaticae, 73(4):535–559, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guillen-Scholten:2006:CCM

[744] Juan Guillen-Scholten, Farhad Arbab, Frank de Boer, and Marcello Bonsangue. A component coordination model based on mobile channels. Fundamenta Informaticae, 73(4):561–582, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vallecillo:2006:TBS

[745] Antonio Vallecillo, Vasco T. Vasconcelos, and António Ravara. Typing the behavior of software components using session types. Fundamenta Informaticae, 73(4):583–598, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engels:2006:SIS

[746] Gregor Engels, Fernando Orejas, and Francesco Parisi-Presicce. Special issue with selected papers from ICGT 2004. Fundamenta Informaticae, 74(1):??, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrig:2006:AHL

[747] Hartmut Ehrig, Julia Padberg, Ulrike Prange, and Annegret Habel. Adhesive high-level replacement systems: A new categorical framework for graph transformation. Fundamenta Informaticae, 74(1):1–29, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrig:2006:FTT

[748] Hartmut Ehrig, Karsten Ehrig, Ulrike Prange, and Gabriele Taentzer. Fundamental theory for typed attributed graphs and graph transformation based on adhesive HLR categories. Fundamenta Informaticae, 74 (1):31–61, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heckel:2006:SGT

[749] Reiko Heckel, Georgios Lajios, and Sebastian Menge. Stochastic graph transformation systems. Fundamenta Informaticae, 74(1):63–84, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chalopin:2006:LCG

[750] Jérémie Chalopin, Yves Métivier, and Wiesław Zielonka. Local computations in graphs: The case of cellular edge local computations. Fundamenta Informaticae, 74(1):85–114, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:2006:EPS

[751] Andrzej Ehrenfeucht, Jurriaan Hage, Tero Harju, and Grzegorz Rozenberg. The embedding problem for switching classes of graphs. Fundamenta Informaticae, 74(1):115–134, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrig:2006:TCA

[752] Hartmut Ehrig, Karsten Ehrig, Annegret Habel, and Karl-Heinz Pennemann. Theory of constraints and application conditions: From graphs to high-level structures. Fundamenta Informaticae, 74(1):135–166, October 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bednarczyk:2006:LM

[753] Marek A. Bednarczyk. Limits of modularity. Fundamenta Informaticae, 74(2–3):167–187, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2006:RDE

[754] Chin-Chen Chang and Chih-Yang Lin. Reversible data embedding based on prediction approach for VQ and SMVQ compressed images. Fundamenta Informaticae, 74(2–3):189–207, November 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dimov:2006:CARa

[755] Georgi Dimov and Dimiter Vakarelov. Contact algebras and region-based theory of space: a proximity approach — I. Fundamenta Informaticae, 74(2–3):209–249, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dimov:2006:CARb

[756] Georgi Dimov and Dimiter Vakarelov. Contact algebras and region-based theory of space: Proximity approach — II. Fundamenta Informaticae, 74(2–3):251–282, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2006:RRA

[757] Ivo Düntsch and Michael Winter. Rough relation algebras revisited. Fundamenta Informaticae, 74(2–3):283–300, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hitzler:2006:CVA

[758] Pascal Hitzler, Markus Krötzsch, and Guo-Qiang Zhang. A categorical view on algebraic lattices in formal concept analysis. Fundamenta Informaticae, 74(2–3):301–328, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalyanasundaram:2006:CTM

[759] Bala Kalyanasundaram and Mahe Velauthapillai. Capabilities of thoughtful machines. Fundamenta Informaticae, 74(2–3):329–340, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nadarajah:2006:LCL

[760] Saralees Nadarajah and Samuel Kotz. The linear combination of logistic and Gumbel random variables. Fundamenta Informaticae, 74(2–3):341– 350, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shah:2006:PSI

[761] Mohak Shah, Marina Sokolova, and Stan Szpakowicz. Process-specific information for learning electronic negotiation outcomes. Fundamenta Informaticae, 74(2–3):351–373, November 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szpyrka:2006:ARN

[762] Marcin Szpyrka. Analysis of RTCP-nets with reachability graphs. Fundamenta Informaticae, 74(2–3):375–390, November 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2006:MCU

[763] Maurice Margenstern. Machines, computations and universality. Fundamenta Informaticae, 74(4):??, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2006:TSU

[764] Artiom Alhazov, Rudolf Freund, Alberto Leporati, Marion Oswald, and Claudio Zandron. (tissue) P systems with unit rules and energy assigned to membranes. Fundamenta Informaticae, 74(4):391–408, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bournez:2006:RAC

[765] Olivier Bournez and Emmanuel Hainry. Recursive analysis characterized as a class of real recursive functions. Fundamenta Informaticae, 74(4): 409–433, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burderi:2006:HTD

[766] Fabio Burderi, Giuseppa Castiglione, and Antonio Restivo. Higman’s theorem on discrete sets. Fundamenta Informaticae, 74(4):435–446, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cavaliere:2006:OSR

[767] Matteo Cavaliere, Peter, and Leupold. Observation of string-rewriting systems. Fundamenta Informaticae, 74(4):447–462, December 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delvenne:2006:DUS

[768] Jean-Charles Delvenne, Petr Kůrka, and Vincent Blondel. Decidability and universality in symbolic dynamical systems. Fundamenta Informaticae, 74(4):463–490, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Durand-Lose:2006:AGC

[769] Jérôme Durand-Lose. Abstract geometrical computation 1: Embedding black hole computations with rational numbers. Fundamenta Informaticae, 74(4):491–510, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2006:VCS

[770] Markus Holzer and Martin Kutrib. Variable complexity of simple programs. Fundamenta Informaticae, 74(4):511–528, December 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leporati:2006:RSS

[771] Alberto Leporati, Claudio Zandron, and Giancarlo Mauri. Reversible P systems to simulate Fredkin circuits. Fundamenta Informaticae, 74(4): 529–548, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Neraud:2006:CCS

[772] Jean Néraud. On the completion of codes in submonoids with finite rank. Fundamenta Informaticae, 74(4):549–562, December 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okhotin:2006:CUO

[773] Alexander Okhotin. Computational universality in one-variable language equations. Fundamenta Informaticae, 74(4):563–578, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schott:2006:SWA

[774] René Schott and Jean-Claude Spehner. Shuffle of words and Araucaria trees. Fundamenta Informaticae, 74(4):579–601, December 2006. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See erratum [977].

Umeo:2006:SEM

[775] Hiroshi Umeo, Masashi Maeda, Masaya Hisaoka, and Masato Teraoka. A state-efficient mapping scheme for designing two-dimensional firing squad synchronization algorithms. Fundamenta Informaticae, 74(4):603–623, December 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:MPZ

[776] Anonymous. In memory of Professor Zdzisław Pawlak. Fundamenta Informaticae, 75(1–4):??, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2007:LWZ

[777] J. F. Peters and A. Skowron. Life and work of Zdzisław Pawlak. Fundamenta Informaticae, 75(1–4):??, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2007:ALA

[778] Philippe Balbiani, Dimiter, and Vakarelov. Arrow logic with arbitrary intersections: Applications to Pawlak’s information systems. Fundamenta Informaticae, 75(1–4):1–25, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2007:RSA

[779] Jan G. Bazan, Piotr Kruczek, Stanisława Bazan-Socha, Andrzej Skowron, and Jacek J. Pietrzyk. Rough set approach to behavioral pattern identification. Fundamenta Informaticae, 75(1–4):27–47, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardini:2007:MBS

[780] Francesco Bernardini, Robert Brijder, Grzegorz Rozenberg, and Claudio Zandron. Multiset-based self-assembly of graphs. Fundamenta Informaticae, 75(1–4):49–75, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bianucci:2007:ECE

[781] Daniela Bianucci, Gianpiero Cattaneo, and Davide Ciucci. Entropies and co-entropies of coverings with application to incomplete information systems. Fundamenta Informaticae, 75(1–4):77–105, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buszkowski:2007:RSL

[782] Wojciech Buszkowski. Rough sets and learning by unification. Fundamenta Informaticae, 75(1–4):107–121, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:2007:RDI

[783] Mihir K. Chakraborty and Mohua Banerjee. Rough dialogue and implication lattices. Fundamenta Informaticae, 75(1–4):123–139, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2007:SLG

[784] Haiming Chen, Rudolf Freund, Mihai Ionescu, Gheorghe Păun, and Mario J. Pérez-Jiménez. On string languages generated by spiking neural P systems. Fundamenta Informaticae, 75(1–4):141–162, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Demri:2007:RNI

[785] Stéphane Demri, Ewa, and Orłowska. Relative nondeterministic information logic is EXPTIME-complete. Fundamenta Informaticae, 75(1–4): 163–178, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:2007:CFB

[786] Patrick Doherty and Andrzej Szałas. A correspondence framework between three-valued logics and similarity-based approximate reasoning. Fundamenta Informaticae, 75(1–4):179–193, January 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dubois:2007:PTV

[787] Didier Dubois, Florence Dupin de Saint-Cyr, and Henri Prade. A possibility-theoretic view of formal concept analysis. Fundamenta Informaticae, 75(1–4):195–213, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2007:MML

[788] Ivo Düntsch and Beata Konikowska. A multi-modal logic for disagreement and exhaustiveness. Fundamenta Informaticae, 75(1–4):215–238, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dziubinski:2007:CIM

[789] Marcin Dziubiński, Rineke Verbrugge, and Barbara Dunin-Kȩplicz. Complexity issues in multiagent logics. Fundamenta Informaticae, 75(1–4): 239–262, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrenfeucht:2007:RS

[790] A. Ehrenfeucht and G. Rozenberg. Reaction systems. Fundamenta Informaticae, 75(1–4):263–280, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gohring:2007:CWM

[791] Daniel Göhring and Hans-Dieter Burkhard. Cooperative world modeling in dynamic multi-robot environments. Fundamenta Informaticae, 75(1– 4):281–294, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2007:TPM

[792] Ryszard Janicki and Dai Tri Man Lê. Towards a pragmatic mereology. Fundamenta Informaticae, 75(1–4):295–314, January 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2007:PIS

[793] Jouni Järvinen. Pawlak’s information systems in terms of Galois connections and functional dependencies. Fundamenta Informaticae, 75(1–4): 315–330, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcus:2007:ZPP

[794] Solomon Marcus. Z. Pawlak, a precursor of DNA computing and of picture grammars. Fundamenta Informaticae, 75(1–4):331–334, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2007:LDG

[795] Antoni Mazurkiewicz. Locally derivable graphs. Fundamenta Informaticae, 75(1–4):335–355, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2007:CPR

[796] Mikhail Ju. Moshkov, Marcin Piliszczuk, and Beata Zielosko. On construction of partial reducts and irreducible partial decision rules. Fundamenta Informaticae, 75(1–4):357–374, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Novotny:2007:RA

[797] Miroslav Novotný. Retracts of algebras. Fundamenta Informaticae, 75(1– 4):375–384, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ohsuga:2007:BGB

[798] Setsuo Ohsuga. Bridging the gap between non-symbolic and symbolic processing — how could human being acquire language? Fundamenta Informaticae, 75(1–4):385–406, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2007:NSS

[799] James F. Peters. Near sets. Special theory about nearness of objects. Fundamenta Informaticae, 75(1–4):407–433, January 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2007:CAM

[800] Lech Polkowski. Concept approximation in mathematics and computer science. An essay in homage to Zdzisław Pawlak. Fundamenta Informaticae, 75(1–4):435–451, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ramanna:2007:ACD

[801] Sheela Ramanna, James F. Peters, and Andrzej Skowron. Approaches to conflict dynamics based on rough sets. Fundamenta Informaticae, 75(1– 4):453–468, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salomaa:2007:SBB

[802] Arto Salomaa. Subword balance in binary words, languages and sequences. Fundamenta Informaticae, 75(1–4):469–482, January 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sen:2007:HTU

[803] Debashis Sen and Sankar K. Pal. Histogram thresholding using beam theory and ambiguity measures. Fundamenta Informaticae, 75(1–4):483– 504, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:2007:GAB

[804] Władysław Skarbek. On generating all binary trees. Fundamenta Informaticae, 75(1–4):505–536, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2007:BA

[805] Józef Winkowski. Behaviour algebras. Fundamenta Informaticae, 75(1– 4):537–560, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Busch:2007:OBA

[806] J. Busch. On the optimality of the binary algorithm for the Jacobi symbol. Fundamenta Informaticae, 76(1–2):1–11, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2007:TKA

[807] Ya-Fen Chang and Chin-Chen Chang. Tolerant key assignment for enforcing complicated access control policies in a hierarchy. Fundamenta Informaticae, 76(1–2):13–23, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2007:FSE

[808] Chin-Chen Chang and Chang-Chu Chen. Full-searching-equivalent vector quantization using two-bounds triangle inequality. Fundamenta Informaticae, 76(1–2):25–37, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2007:VCS

[809] Chin-Chen Chang, Tzu-Chuen Lu, and Jun-Bin Ya. VQ codebook searching algorithm based on correlation property. Fundamenta Informaticae, 76(1–2):39–57, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chunlin:2007:PAG

[810] Li Chunlin and Li Layuan. A pricing approach for grid resource scheduling with QoS guarantees. Fundamenta Informaticae, 76(1–2):59–73, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dey:2007:CPC

[811] Sabyasachi Dey, Bhargab B. Bhattacharya, Malay K. Kundu, Arijit Bishnu, and Tinku Acharya. A co-processor for computing the Euler number of a binary image using divide-and-conquer strategy. Fundamenta Informaticae, 76(1–2):75–89, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Formenti:2007:ASS

[812] Enrico Formenti, Benoît Masson, and Theophilos Pisokas. Advances in symmetric sandpiles. Fundamenta Informaticae, 76(1–2):91–112, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2007:IRL

[813] Yo-Sub Han, Kai Salomaa, and Derick Wood. Intercode regular languages. Fundamenta Informaticae, 76(1–2):113–128, January 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2007:MLO

[814] Jouni Järvinen, Michiro Kondo, and Jari Kortelainen. Modal-like operators in Boolean lattices, Galois connections and fixed points. Fundamenta Informaticae, 76(1–2):129–145, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paquet:2007:PSL

[815] Ulrich Paquet and Andries P. Engelbrecht. Particle swarms for linearly constrained optimisation. Fundamenta Informaticae, 76(1–2):147–170, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2007:RSB

[816] Suman Saha, C. A. Murthy, and Sankar K. Pal. Rough set based ensemble classifier for Web page classification. Fundamenta Informaticae, 76(1–2):171–187, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tarasyuk:2007:SPB

[817] Igor V. Tarasyuk. Stochastic Petri box calculus with discrete time. Fundamenta Informaticae, 76(1–2):189–218, January 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:DGS

[818] Anonymous. Developments in grammar systems: Preface. Fundamenta Informaticae, 76(3):??, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bel-Enguix:2007:DMM

[819] Gemma Bel-Enguix and M. Dolores Jiménez-López. Dynamic meaning membrane systems: An application to the description of semantic change. Fundamenta Informaticae, 76(3):219–237, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bensch:2007:ASP

[820] Suna Bensch and Henning Bordihn. Active symbols in pure systems. Fundamenta Informaticae, 76(3):239–254, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bordihn:2007:CDG

[821] Henning Bordihn and Markus Holzer. Cooperating distributed grammar systems as models of distributed problem solving, revisited. Fundamenta Informaticae, 76(3):255–270, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2007:GSV

[822] Erzsébet Csuhaj-Varjü, Gheorghe Păun, and György Vaszil. Grammar systems versus membrane computing: The case of CD grammar systems. Fundamenta Informaticae, 76(3):271–292, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurgen:2007:CDG

[823] Jürgen and Dassow. On cooperating distributed grammar systems with competence based start and stop conditions. Fundamenta Informaticae, 76(3):293–304, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Freund:2007:MGS

[824] Rudolf Freund and Marion Oswald. Modelling grammar systems by tissue P systems working in the sequential mode. Fundamenta Informaticae, 76(3):305–323, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grando:2007:PCB

[825] María Adela Grando and Victor Mitrana. A possible connection between two theories: Grammar systems and concurrent programming. Fundamenta Informaticae, 76(3):325–336, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kelemen:2007:CB

[826] Jozef Kelemen. On computing bodies. Fundamenta Informaticae, 76 (3):337–347, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kelemenova:2007:MHE

[827] Alica Kelemenová and Michal Tupý. Monocultures and homogeneous environment in eco-grammar systems. Fundamenta Informaticae, 76(3): 349–365, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sebestyen:2007:MMR

[828] Peter Sebestyén and Petr Sosík. Modelling multiple robots in space: An adaptive eco-grammar system. Fundamenta Informaticae, 76(3):367– 381, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sunckel:2007:MCG

[829] Bettina Sunckel. On metalinear CD grammar systems. Fundamenta Informaticae, 76(3):383–397, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2007:SIH

[830] Chin-Chen Chang, Chih-Yang Lin, and Chun-Sen Tseng. Secret image hiding and sharing based on the (t, n)-threshold. Fundamenta Informaticae, 76(4):399–411, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konstantinidis:2007:FRR

[831] Stavros Konstantinidis, Santean Nicolae, and Sheng Yu. Fuzzification of rational and recognizable sets. Fundamenta Informaticae, 76(4):413–447, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peltier:2007:RCS

[832] Nicolas Peltier. A resolution calculus with shared literals. Fundamenta Informaticae, 76(4):449–480, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2007:IRM

[833] Hsien-Chu Wu and Chin-Chen Chang. An image retrieval method based on color-complexity and spatial-histogram features. Fundamenta Informaticae, 76(4):481–493, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ziembinski:2007:ACB

[834] Radosław Ziembiński. Algorithms for context based sequential pattern mining. Fundamenta Informaticae, 76(4):495–510, December 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:ASM

[835] Anonymous. The abstract state machines method. Fundamenta Informaticae, 77(1–2):??, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abrial:2007:RDI

[836] Jean-Raymond Abrial and Stefan Hallerstede. Refinement, decomposition, and instantiation of discrete models: Application to Event-B. Fundamenta Informaticae, 77(1–2):1–28, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Banach:2007:RPB

[837] Richard Banach, Czesław Jeske, Michael Poppleton, and Susan Stepney. Retrenching the purse: The balance enquiry quandary, and generalised and (1, 1) forward refinements. Fundamenta Informaticae, 77(1–2):29– 69, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farahbod:2007:CEA

[838] Roozbeh Farahbod, Vincenzo Gervasi, and Uwe Glässer. CoreASM: An extensible ASM execution engine. Fundamenta Informaticae, 77(1–2): 71–103, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tang:2007:MCA

[839] Calvin Kai Fan Tang and Eugenia Ternovska. Model checking abstract state machines with answer set programming. Fundamenta Informaticae, 77(1–2):105–141, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Graf:2007:TSM

[840] Susanne Graf and Andreas Prinz. Time in state machines. Fundamenta Informaticae, 77(1–2):143–174, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seshadhri:2007:RSB

[841] Comandur Seshadhri, Anil Seth, and Somenath Biswas. RAM simulation of BGS model of abstract-state machines. Fundamenta Informaticae, 77 (1–2):175–185, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Badouel:2007:ATG

[842] Eric Badouel, Jules Chenou, and Goulven Guillou. An axiomatization of the token game based on Petri algebras. Fundamenta Informaticae, 77(3):187–215, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2007:NIH

[843] Chin-Chen Chang, Wei-Liang Tai, and Chia-Chen Lin. A novel image hiding scheme based on VQ and Hamming distance. Fundamenta Informaticae, 77(3):217–228, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2007:AAC

[844] Yancang Li and Wanqing Li. Adaptive ant colony optimization algorithm based on information entropy: Foundation and application. Fundamenta Informaticae, 77(3):229–242, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pechenizkiy:2007:FED

[845] Mykola Pechenizkiy, Alexey Tsymbal, Seppo Puuronen, and David Patterson. Feature extraction for dynamic integration of classifiers. Fundamenta Informaticae, 77(3):243–275, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Porwik:2007:EAA

[846] Piotr Porwik. Efficient algorithm of affine form searching for weakly specified Boolean function. Fundamenta Informaticae, 77(3):277–291, July 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:TLC

[847] Anonymous. Typed lambda calculi and applications 2005, selected papers. Fundamenta Informaticae, 77(4):??, December 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Coquand:2007:CTC

[848] Thierry Coquand. The completeness of typing for context-semantics. Fundamenta Informaticae, 77(4):293–301, December 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Felty:2007:TES

[849] Amy P. Felty. Tutorial examples of the semantic approach to foundational proof-carrying code. Fundamenta Informaticae, 77(4):303–330, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hayashi:2007:CPA

[850] Susumu Hayashi. Can proofs be animated by games? Fundamenta Informaticae, 77(4):331–343, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abel:2007:UAE

[851] Andreas Abel and Thierry Coquand. Untyped algorithmic equality for Martin-Löf’s logical framework with surjective pairs. Fundamenta Informaticae, 77(4):345–395, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ahmed:2007:LLL

[852] Amal Ahmed, Matthew Fluet, and Greg Morrisett. L^{3}: A linear language with locations. Fundamenta Informaticae, 77(4):397–449, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Damiani:2007:RIR

[853] Ferruccio Damiani. Rank 2 intersection for recursive definitions. Fundamenta Informaticae, 77(4):451–488, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

David:2007:APS

[854] René David and Karim Nour. Arithmetical proofs of strong normalization results for symmetric λ-calculi. Fundamenta Informaticae, 77(4):489– 510, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laird:2007:EAP

[855] James D. Laird. On the expressiveness of affine programs with non-local control: The elimination of nesting in SPCF. Fundamenta Informaticae, 77(4):511–531, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:FSI

[856] Anonymous. The fourth special issue on applications of concurrency to system design. Fundamenta Informaticae, 78(1):??, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clariso:2007:VCS

[857] Robert Clarisó and Jordi Cortadella. Verification of concurrent systems with parametric delays using octahedra. Fundamenta Informaticae, 78 (1):1–33, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geilen:2007:APP

[858] Marc Geilen, Twan Basten, Bart Theelen, and Ralph Otten. An algebra of Pareto points. Fundamenta Informaticae, 78(1):35–74, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorgonio:2007:ASA

[859] Kyller Gorgônio, Jordi Cortadella, Fei Xia, and Alex Yakovlev. Automating synthesis of asynchronous communication mechanisms. Fundamenta Informaticae, 78(1):75–100, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Josephs:2007:CDI

[860] Mark B. Josephs and Hemangee K. Kapoor. Controllable delayinsensitive processes. Fundamenta Informaticae, 78(1):101–130, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Potop-Butucaru:2007:CCA

[861] Dumitru Potop-Butucaru and Benoît Caillaud. Correct-by-construction asynchronous implementation of modular synchronous specifications. Fundamenta Informaticae, 78(1):131–159, September 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vogler:2007:IDS

[862] W. Vogler and B. Kangsah. Improved decomposition of signal transition graphs. Fundamenta Informaticae, 78(1):161–197, September 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bolikowski:2007:NMP

[863] Łukasz Bolikowski and Anna Gambin. New metrics for phylogenies. Fundamenta Informaticae, 78(2):199–216, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Breysse:2007:NAC

[864] Olivia Breysse and Michel De Glas. A new approach to the concepts of boundary and contact: Toward an alternative to mereotopology. Fundamenta Informaticae, 78(2):217–238, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2007:PGI

[865] Yu-Chen Hu. Predictive grayscale image coding scheme using VQ and BTC. Fundamenta Informaticae, 78(2):239–255, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2007:BPV

[866] Yu-Chen Hu, Chia-Chen Lin, and Kang-Liang Chi. Block prediction vector quantization for grayscale image compression. Fundamenta Informaticae, 78(2):257–270, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huynh:2007:AFL

[867] Van Nam Huynh, Yoshiteru Nakamori, and Tetsuya Murai. An algebraic foundation for linguistic reasoning. Fundamenta Informaticae, 78(2): 271–294, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Palka:2007:ISS

[868] Ewa Palka. An infinitary sequent system for the equational theory of *-continuous action lattices. Fundamenta Informaticae, 78(2):295–309, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:F

[869] Anonymous. Foreword. Fundamenta Informaticae, 78(3):??, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Freund:2007:CAP

[870] Rudolf Freund and Fritz Tafill. Cellular automata and parallel array systems. Fundamenta Informaticae, 78(3):311–327, November 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fuks:2007:RCB

[871] Henryk Fukś. Remarks on the critical behavior of second order additive invariants in elementary cellular automata. Fundamenta Informaticae, 78(3):329–341, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klein:2007:CDU

[872] Andreas Klein and Martin Kutrib. Cellular devices and unary languages. Fundamenta Informaticae, 78(3):343–368, November 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2007:RHP

[873] Pradipta Maji and P. Pal Chaudhuri. RBFFCA: A hybrid pattern classifier using radial basis function and fuzzy cellular automata. Fundamenta Informaticae, 78(3):369–396, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nishio:2007:ASN

[874] Hidenosuke Nishio, Maurice Margenstern, and Friedrich von Haeseler. On algebraic structure of neighborhoods of cellular automata. Fundamenta Informaticae, 78(3):397–416, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pivato:2007:SDB

[875] Marcus Pivato. Spectral domain boundaries in cellular automata. Fundamenta Informaticae, 78(3):417–447, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umeo:2007:SEO

[876] Hiroshi Umeo, Koshi Michisaka, Naoki Kamikawa, and Masaru Kanazawa. State-efficient one-bit communication solutions for some classical cellular automata problems. Fundamenta Informaticae, 78(3): 449–465, November 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:SIK

[877] Anonymous. Special issue on knowledge discovery. Fundamenta Informaticae, 78(4):??, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2007:SNF

[878] Liying Jiang and Jitender Deogun. SPICE: A new framework for data mining based on probability logic and formal concept analysis. Fundamenta Informaticae, 78(4):467–485, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Galassi:2007:HHM

[879] Ugo Galassi, Marco Botta, and Attilio Giordana. Hierarchical hidden Markov models for user/process profile learning. Fundamenta Informaticae, 78(4):487–505, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Im:2007:PAD

[880] Seunghyun Im. Privacy aware data management and chase. Fundamenta Informaticae, 78(4):507–524, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seo:2007:TES

[881] Minkoo Seo, Sanghyun Park, and Jung-Im Won. Towards efficient searching on the secondary structure of protein sequences. Fundamenta Informaticae, 78(4):525–542, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Agier:2007:UFR

[882] Marie Agier, Jean-Marc Petit, and Einoshin Suzuki. Unifying framework for rule semantics: Application to gene expression data. Fundamenta Informaticae, 78(4):543–559, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2007:VDB

[883] Shusaku Tsumoto and Shoji Hirano. Visualization of differences between rules’ syntactic and semantic similarities using multidimensional scaling. Fundamenta Informaticae, 78(4):561–573, December 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salleb-Aouissi:2007:CUD

[884] Ansaf Salleb-Aouissi and Christel Vrain. A contribution to the use of decision diagrams for loading and mining transaction databases. Fundamenta Informaticae, 78(4):575–594, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2007:PPD

[885] Xintao Wu, Yongge Wang, Songtao Guo, and Yuliang Zheng. Privacy preserving database generation for database application testing. Fundamenta Informaticae, 78(4):595–612, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2007:SIH

[886] Xin Zhang and Zbigniew W. Raś. Sound isolation by harmonic peak partition for music instrument recognition. Fundamenta Informaticae, 78(4):613–628, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bell:2007:NES

[887] Paul Bell. A note on the emptiness of semigroup intersections. Fundamenta Informaticae, 79(1–2):1–4, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cowen:2007:DPS

[888] Robert Cowen and Adam Kolany. Davis–Putnam style rules for deciding property S. Fundamenta Informaticae, 79(1–2):5–15, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chaitin:2007:ACH

[889] Gregory Chaitin. An algebraic characterization of the halting probability. Fundamenta Informaticae, 79(1–2):17–23, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eiter:2007:LBA

[890] Thomas Eiter, Esra Erdem, Wolfgang Faber, and Ján Senko. A logicbased approach to finding explanations for discrepancies in optimistic plan execution. Fundamenta Informaticae, 79(1–2):25–69, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herman:2007:PPS

[891] Grzegorz Herman, Tim Paterson, and Michael Soltys. A propositional proof system with quantification over permutations. Fundamenta Informaticae, 79(1–2):71–83, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2007:FMD

[892] Linh Anh Nguyen. Foundations of modal deductive databases. Fundamenta Informaticae, 79(1–2):85–135, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rabinovich:2007:CTG

[893] Alexander Rabinovich. Composition theorem for generalized sum. Fundamenta Informaticae, 79(1–2):137–167, January 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sadowski:2007:OPS

[894] Zenon Sadowski. Optimal proof systems, optimal acceptors and recursive presentability. Fundamenta Informaticae, 79(1–2):169–185, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vennekens:2007:PILa

[895] Joost Vennekens, Johan Wittocx, Maarten Mariën, and Marc Denecker. Predicate introduction for logics with a fixpoint semantics. Part I: Logic programming. Fundamenta Informaticae, 79(1–2):187–208, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vennekens:2007:PILb

[896] Joost Vennekens, Johan Wittocx, Maarten Mariën, and Marc Denecker. Predicate introduction for logics with fixpoint semantics. Part II: Autoepistemic logic. Fundamenta Informaticae, 79(1–2):209–227, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna:2007:BMC

[897] Bożena Woźna and Andrzej Zbrzezny. Bounded model checking for the existential fragment of TCTL_{−G} and diagonal timed automata. Fundamenta Informaticae, 79(1–2):229–256, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2007:CAP

[898] Wei-Guo Zhang and Ying-Luo Wang. A comparative analysis of possibilistic variances and covariances of fuzzy numbers. Fundamenta Informaticae, 79(1–2):257–263, January 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:CSP

[899] Anonymous. Concurrency specification and programming (CS&P). Fundamenta Informaticae, 79(3–4):??, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2007:CSR

[900] Antoni Mazurkiewicz. Compositional systems over reducible networks. Fundamenta Informaticae, 79(3–4):265–282, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2007:IN

[901] Ludwik Czaja. Interpreted nets. Fundamenta Informaticae, 79(3–4): 283–293, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dietze:2007:DPB

[902] Roxana Dietze, Manfred Kudlek, and Olaf Kummer. Decidability problems of a basic class of object nets. Fundamenta Informaticae, 79(3–4): 295–302, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2007:CTM

[903] Berndt Farwer, Manfred Kudlek, and Heiko Rölke. Concurrent Turing machines. Fundamenta Informaticae, 79(3–4):303–317, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2007:ASB

[904] Anna Gomolińska. Approximation spaces based on relations of similarity and dissimilarity of objects. Fundamenta Informaticae, 79(3–4):319–333, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2007:OBS

[905] Damas P. Gruska. Observation based system security. Fundamenta Informaticae, 79(3–4):335–346, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2007:CPA

[906] Kees van Hee, Olivia Oanea, Alexander Serebrenik, Natalia Sidorova, Marc Voorhoeve, and Irina A. Lomazova. Checking properties of adaptive workflow nets. Fundamenta Informaticae, 79(3–4):347–362, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jakubowska:2007:MCT

[907] Gizela Jakubowska and Wojciech Penczek. Modelling and checking timed authentication of security protocols. Fundamenta Informaticae, 79(3–4): 363–378, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janowska:2007:PCT

[908] Agata Janowska and Wojciech Penczek. Path compression in timed automata. Fundamenta Informaticae, 79(3–4):379–399, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler:2007:RPO

[909] Michael Köhler. The reachability problem for object nets. Fundamenta Informaticae, 79(3–4):401–413, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler:2007:FMM

[910] Michael Köhler. A formal model of multi-agent organisations. Fundamenta Informaticae, 79(3–4):415–430, February 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryvyy:2007:ATM

[911] Sergiy Kryvyy and Lyudmila Matvyeyeva. Algorithm of translation of MSC-specified system into Petri net. Fundamenta Informaticae, 79(3–4): 431–445, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2007:SSC

[912] Manfred Kudlek. On semilinear sets over commutative semirings. Fundamenta Informaticae, 79(3–4):447–452, February 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurkowski:2007:VSP

[913] Mirosław Kurkowski and Wojciech Penczek. Verifying security protocols modelled by networks of automata. Fundamenta Informaticae, 79(3–4): 453–471, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomuscio:2007:VTP

[914] Alessio Lomuscio, Franco Raimondi, and Bożena Woźna. Verification of the TESLA protocol in MCMAS-X. Fundamenta Informaticae, 79(3–4): 473–486, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2007:LPT

[915] Antoni Mazurkiewicz. Local properties of triangular graphs. Fundamenta Informaticae, 79(3–4):487–495, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2007:NOE

[916] James F. Peters, Andrzej Skowron, and Jarosław Stepaniuk. Nearness of objects: Extension of approximation space model. Fundamenta Informaticae, 79(3–4):497–512, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2007:PEG

[917] Roman R. Redziejowski. Parsing expression grammar as a primitive recursive-descent parser with backtracking. Fundamenta Informaticae, 79(3–4):513–524, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:2007:RDR

[918] Jarosław Stepaniuk. Relational data and rough sets. Fundamenta Informaticae, 79(3–4):525–539, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2007:AAP

[919] Zbigniew Suraj and Barbara Fryc. Analysis of approximate Petri nets by means of occurrence graphs. Fundamenta Informaticae, 79(3–4):541–551, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2007:TEC

[920] Zbigniew Suraj and Krzysztof Pancerz. Towards efficient computing consistent and partially consistent extensions of information systems. Fundamenta Informaticae, 79(3–4):553–566, February 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2007:ASN

[921] Marcin Wolski. Approximation spaces and nearness type structures. Fundamenta Informaticae, 79(3–4):567–577, February 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2007:SBR

[922] Andrzej Zbrzezny and Agata Półrola. SAT-based reachability checking for timed automata with discrete data. Fundamenta Informaticae, 79(3– 4):579–593, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:AMH

[923] Anonymous. Antoni Mazurkiewicz — half a century of inspirational research. Fundamenta Informaticae, 80(1–3):??, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diekert:2007:FOF

[924] Volker Diekert, Martin Horsch, and Manfred Kufleitner. On first-order fragments for Mazurkiewicz traces. Fundamenta Informaticae, 80(1–3): 1–29, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardinello:2007:PSB

[925] Luca Bernardinello, Elena Monticelli, and Lucia Pomello. On preserving structural and behavioural properties by composing net systems on interfaces. Fundamenta Informaticae, 80(1–3):31–47, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Billington:2007:FTD

[926] Jonathan Billington and Bing Han. Formalising TCP’s data transfer service language: A symbolic automaton and its properties. Fundamenta Informaticae, 80(1–3):49–74, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Best:2007:MPN

[927] Eike Best, Philippe Darondeau, and Harro Wimmel. Making Petri nets safe and free of internal transitions. Fundamenta Informaticae, 80(1–3): 75–90, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Conforti:2007:SBU

[928] Giovanni Conforti, Damiano Macedonio, and Vladimiro Sassone. Static BiLog: a unifying language for spatial structures. Fundamenta Informaticae, 80(1–3):91–110, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Engelfriet:2007:FCT

[929] Joost Engelfriet and Hendrik Jan Hoogeboom. Finitary compositions of two-way finite-state transductions. Fundamenta Informaticae, 80(1– 3):111–123, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carette:2007:CPN

[930] Jacques Carette and Ryszard Janicki. Computing properties of numerical imperative programs by symbolic computation. Fundamenta Informaticae, 80(1–3):125–146, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Genest:2007:CAB

[931] Blaise Genest, Dietrich Kuske, and Anca Muscholl. On communicating automata with bounded channels. Fundamenta Informaticae, 80(1–3): 147–167, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gastin:2007:USP

[932] Paul Gastin and Dietrich Kuske. Uniform satisfiability in PSPACE for local temporal logics over Mazurkiewicz traces. Fundamenta Informaticae, 80(1–3):169–197, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kleijn:2007:PPN

[933] Jetty Kleijn and Maciej Koutny. Processes of Petri nets with range testing. Fundamenta Informaticae, 80(1–3):199–219, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chalopin:2007:EMP

[934] Jérémie Chalopin and Yves Métivier. An efficient message passing election algorithm based on Mazurkiewicz’s algorithm. Fundamenta Informaticae, 80(1–3):221–246, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2007:MRS

[935] Mikhail Ju. Moshkov, Andrzej Skowron, and Zbigniew Suraj. On minimal rule sets for almost all binary information systems. Fundamenta Informaticae, 80(1–3):247–258, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ochmanski:2007:EMT

[936] Edward Ochmański and Joanna Pieckowska. On ethics of Mazurkiewicz traces. Fundamenta Informaticae, 80(1–3):259–272, March 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Reisig:2007:DPE

[937] Wolfgang Reisig. The decent philosophers: An exercise in concurrent behaviour. Fundamenta Informaticae, 80(1–3):273–281, March 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2007:PAV

[938] Xu Wang and Marta Kwiatkowska. On process-algebraic verification of asynchronous circuits. Fundamenta Informaticae, 80(1–3):283–310, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2007:TFM

[939] Józef Winkowski. Towards a framework for modelling behaviours of hybrid systems. Fundamenta Informaticae, 80(1–3):311–332, March 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2007:MCS

[940] Jan A. Bergstra and Cornelis A. Middelburg. Maurer computers with single-thread control. Fundamenta Informaticae, 80(4):333–362, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bui:2007:HCT

[941] Alain Bui and Devan Sohier. How to compute times of random walks based distributed algorithms. Fundamenta Informaticae, 80(4):363–378, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Flasinski:2007:IPG

[942] Mariusz Flasiński. Inference of parsable graph grammars for syntactic pattern recognition. Fundamenta Informaticae, 80(4):379–413, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hebert:2007:DMF

[943] Céline Hébert, Alain Bretto, and Bruno Crémilleux. A data mining formalization to improve hypergraph minimal transversal computation. Fundamenta Informaticae, 80(4):415–433, December 2007. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2007:FCA

[944] Feng Jiang, Yuefei Sui, and Cungen Cao. Formal concept analysis in relational database and rough relational database. Fundamenta Informaticae, 80(4):435–451, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2007:SAP

[945] Fagen Li and Yupu Hu. Security analysis of the Pomykała–Barabasz scheme. Fundamenta Informaticae, 80(4):453–459, December 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See [622].

Lu:2007:NPL

[946] Rongxing Lu, Zhenfu Cao, and Xiaolei Dong. A new practical limited identity-based encryption scheme. Fundamenta Informaticae, 80(4):461– 474, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2007:RHC

[947] Pradipta Maji and Sankar K. Pal. RFCM: A hybrid clustering algorithm using rough and fuzzy sets. Fundamenta Informaticae, 80(4):475–496, December 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2007:P

[948] Anonymous. Preface. Fundamenta Informaticae, 81(1–3):??, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krajewski:2007:AGL

[949] Stanisław Krajewski and Jan Woleński. Andrzej Grzegorczyk: Logic and philosophy. Fundamenta Informaticae, 81(1–3):1–17, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arnold:2007:CSG

[950] André Arnold and Damian Niwiński. Continuous separation of game languages. Fundamenta Informaticae, 81(1–3):19–28, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2007:MLR

[951] Philippe Balbiani, Tinko Tinchev, and Dimiter Vakarelov. Modal logics for region-based theories of space. Fundamenta Informaticae, 81(1–3): 29–82, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cegielski:2007:ATP

[952] Patrick Cégielski, Denis Richard, and Maxim Vsemirnov. On the additive theory of prime numbers. Fundamenta Informaticae, 81(1–3):83–96, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czelakowski:2007:IPS

[953] Janusz Czelakowski. Induction principles for sets. Fundamenta Informaticae, 81(1–3):97–122, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See corrigendum [1302].

Gottwald:2007:MFL

[954] Siegfried Gottwald. Mathematical fuzzy logic: An invitation to interesting research areas. Fundamenta Informaticae, 81(1–3):123–137, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruszczynski:2007:PS

[955] Rafał Gruszczyński and Andrzej Pietruszczak. Pieri’s structures. Fundamenta Informaticae, 81(1–3):139–154, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hajek:2007:MFL

[956] Petr Hájek. Mathematical fuzzy logic and natural numbers. Fundamenta Informaticae, 81(1–3):155–163, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kossak:2007:SUN

[957] Roman Kossak. On some uses of nonstandard models. Fundamenta Informaticae, 81(1–3):165–172, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krajewski:2007:GTM

[958] Stanisław Krajewski. On Gödel’s Theorem and mechanism: Inconsistency or unsoundness is unavoidable in any attempt to ’out-Gödel’ the mechanist. Fundamenta Informaticae, 81(1–3):173–181, May 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krynicki:2007:FA

[959] Michał Krynicki, Marcin Mostowski, and Konrad Zdanowski. Finite arithmetics. Fundamenta Informaticae, 81(1–3):183–202, May 2007. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maksimova:2007:MGL

[960] Larisa Maksimova. On modal Grzegorczyk logic. Fundamenta Informaticae, 81(1–3):203–210, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marek:2007:CPS

[961] Victor W. Marek and Jeffrey B. Remmel. Compactness properties for stable semantics of logic programs. Fundamenta Informaticae, 81(1–3): 211–239, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mostowski:2007:PIC

[962] Marcin Mostowski. Potential infinity and the Church thesis. Fundamenta Informaticae, 81(1–3):241–248, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murawski:2007:DLN

[963] Roman Murawski. Did Leibniz and Newton discover or create the calculus? Fundamenta Informaticae, 81(1–3):249–256, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romanowska:2007:ABA

[964] Anna B. Romanowska, Jonathan D. H. Smith, and Ewa Orłowska. Abstract barycentric algebras. Fundamenta Informaticae, 81(1–3):257–273, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orlowska:2007:DDH

[965] Ewa Orłowska and Ingrid Rewitzky. Discrete dualities for Heyting algebras with operators. Fundamenta Informaticae, 81(1–3):275–295, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rybakov:2007:DWR

[966] Vladimir V. Rybakov. Decidability w.r.t. logical consecutions of linear temporal logic extended by since and previous. Fundamenta Informaticae, 81(1–3):297–313, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salwicki:2007:AGC

[967] Andrzej Salwicki. Andrzej Grzegorczyk’s contribution to computer science. Fundamenta Informaticae, 81(1–3):315–323, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Srebrny:2007:PPE

[968] Marian Srebrny and Lidia Stepień. A propositional programming environment for linear algebra. Fundamenta Informaticae, 81(1–3):325–345, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Svejdar:2007:IRA

[969] Vítězslav Švejdar. An interpretation of Robinson arithmetic in its Grzegorczyk’s weaker variant. Fundamenta Informaticae, 81(1–3):347–354, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolenski:2007:LRP

[970] Jan Woleński. Logic, rationality and psychologism. Fundamenta Informaticae, 81(1–3):355–365, May 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bedregal:2007:PAF

[971] Benjamín René Callejas Bedregal. Pushdown automata free of explicit nondeterminism and an infinite hierarchy of context-free languages. Fundamenta Informaticae, 81(4):367–377, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diaconescu:2007:TOM

[972] Denisa Diaconescu and George Georgescu. Tense operators on MValgebras and Łukasiewicz–Moisil algebras. Fundamenta Informaticae, 81(4):379–408, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernandez-Camacho:2007:BST

[973] María-Inés Fernández-Camacho and José-Ramón Sánchez-Couso. Binary search trees, recurrent properties and wave equations. Fundamenta Informaticae, 81(4):409–439, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2007:OFR

[974] Yo-Sub Han and Derick Wood. Outfix-free regular languages and prime outfix-free decomposition. Fundamenta Informaticae, 81(4):441–457, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manea:2007:GAP

[975] Florin Manea and Călina Ploscaru. A generalization of the assignment problem, and its application to the rank aggregation problem. Fundamenta Informaticae, 81(4):459–471, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marciniec:2007:UCL

[976] Jacek Marciniec. From unification compactness to learnability. Fundamenta Informaticae, 81(4):473–484, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schott:2007:ESW

[977] René Schott and Jean-Claude Spehner. Erratum for “Shuffle of Words and Araucaria Trees”. Fundamenta Informaticae, 81(4):485–490, June 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See [774].

Chang:2008:TPL

[978] Chin-Chen Chang, Chih-Yang Lin, and Yi-Pei Hsieh. Three-phase lossless data hiding method for the VQ index table. Fundamenta Informaticae, 82(1–2):1–13, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chiang:2008:CPB

[979] Yung-Kuei Chiang, Piyu Tsai, and Feng-Long Huang. Codebook partition based steganography without member restriction. Fundamenta Informaticae, 82(1–2):15–27, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2008:ECS

[980] Erzsébet Csuhaj-Varjú, Antonio Di Nola, Gheorghe Păun, Mario J. Pérez-Jiménez, and György Vaszil. Editing configurations of P systems. Fundamenta Informaticae, 82(1–2):29–46, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drewes:2008:PLR

[981] Frank Drewes and Brink van der Merwe. Path languages of random permitting context tree grammars are regular. Fundamenta Informaticae, 82(1–2):47–60, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esik:2008:PTA

[982] Zoltán Ésik and Szabolcs Iván. Products of tree automata with an application to temporal logic. Fundamenta Informaticae, 82(1–2):61–78, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esik:2008:SVF

[983] Zoltán Ésik and Szabolcs Iván. Some varieties of finite tree automata related to restricted temporal logics. Fundamenta Informaticae, 82(1–2): 79–103, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Galinski:2008:ISU

[984] Grzegorz Galiński and Władysław Skarbek. Image search using dominant colors. Fundamenta Informaticae, 82(1–2):105–112, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilcinkas:2008:IAB

[985] David Ilcinkas and Andrzej Pelc. Impact of asynchrony on the behavior of rational selfish agents. Fundamenta Informaticae, 82(1–2):113–125, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:2008:IUS

[986] Michael Kaminski. Invariance under stuttering in a temporal logic without the “until” operator. Fundamenta Informaticae, 82(1–2):127–140, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kucera:2008:CSF

[987] Antonín Kučera and Oldřich Stražovský. On the controller synthesis for finite-state Markov decision processes. Fundamenta Informaticae, 82 (1–2):141–153, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2008:HPR

[988] Ching-Chiuan Lin, Nien-Lin Hsueh, and Wen-Hsiang Shen. Highperformance reversible data hiding. Fundamenta Informaticae, 82(1–2): 155–169, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moura:2008:CTA

[989] Arnaldo V. Moura and Guilherme A. Pinto. Classes of timed automata and the undecidability of universality. Fundamenta Informaticae, 82(1– 2):171–184, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2008:ITS

[990] Tomasz Terlikowski. Information in a two-stage decision process. General idea and classification. Fundamenta Informaticae, 82(1–2):185–212, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:RRM

[991] Anonymous. Rules and rule markup languages for the semantic Web. Fundamenta Informaticae, 82(3):??, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deBruijn:2008:RBD

[992] Jos de Bruijn and Stijn Heymans. On the relationship between description logic-based and F -logic-based ontologies. Fundamenta Informaticae, 82(3):213–236, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Behrends:2008:EEA

[993] Erik Behrends, Oliver Fritzen, Wolfgang May, and Franz Schenk. Embedding event algebras and process for ECA rules for the semantic Web. Fundamenta Informaticae, 82(3):237–263, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Damasio:2008:RUR

[994] Carlos Viegas Damásio, Jeff Z. Pan, Giorgos Stoilos, and Umberto Straccia. Representing uncertainty in RuleML. Fundamenta Informaticae, 82 (3):265–288, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lukasiewicz:2008:FDL

[995] Thomas Lukasiewicz. Fuzzy description logic programs under the answer set semantics for the semantic Web. Fundamenta Informaticae, 82(3): 289–310, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:BCC

[996] Anonymous. Behavior of composed concurrent systems: Logic and reasoning. Fundamenta Informaticae, 82(4):??, July 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stehr:2008:CTC

[997] Mark-Oliver Stehr. Compositionality for tightly coupled systems: A new application of the propositions-as-types interpretation. Fundamenta Informaticae, 82(4):311–340, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Johnsen:2008:VBC

[998] Einar Broch Johnsen, Olaf Owe, and Arild B. Torjusen. Validating behavioral component interfaces in rewriting logic. Fundamenta Informaticae, 82(4):341–359, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clarke:2008:BLR

[999] Dave Clarke. A basic logic for reasoning about connector reconfiguration. Fundamenta Informaticae, 82(4):361–390, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abraham:2008:DPS

[1000] Erika Ábrahám, Frank S. de Boer, Willem-Paul de Roever, and Martin Steffen. A deductive proof system for multithreaded Java with exceptions. Fundamenta Informaticae, 82(4):391–463, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esmaeilsabzali:2008:IAC

[1001] Shahram Esmaeilsabzali, Nancy A. Day, and Farhad Mavaddat. Interface automata with complex actions: Limiting interleaving in interface automata. Fundamenta Informaticae, 82(4):465–512, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:ACA

[1002] Anonymous. Algorithmic complexity and applications. Fundamenta Informaticae, 83(1–2):??, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Archibald:2008:RRS

[1003] Margaret Archibald, Vasco Brattka, and Clemens Heuberger. Randomness with respect to the signed-digit representation. Fundamenta Informaticae, 83(1–2):1–19, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bienvenu:2008:SPM

[1004] Laurent Bienvenu, Wolfgang Merkle, and Alexander Shen. A simple proof of Miller–Yu theorem. Fundamenta Informaticae, 83(1–2):21–24, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bridges:2008:PC

[1005] D. S. Bridges and L. S. Vîţă. Proximal connectedness. Fundamenta Informaticae, 83(1–2):25–34, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2008:NCS

[1006] Jürgen Dassow and Ralf Stiebe. Nonterminal complexity of some operations on context-free languages. Fundamenta Informaticae, 83(1–2): 35–49, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2008:BCA

[1007] Henning Fernau and Ralf Stiebe. Blind counter automata on ω-words. Fundamenta Informaticae, 83(1–2):51–64, January 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Freund:2008:RLD

[1008] Rudolf Freund and Marion Oswald. Regular ω-languages defined by finite extended spiking neural P systems. Fundamenta Informaticae, 83 (1–2):65–73, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gao:2008:SCT

[1009] Yuan Gao, Kai Salomaa, and Sheng Yu. The state complexity of two combined operations: Star of catenation and star of reversal. Fundamenta Informaticae, 83(1–2):75–89, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gulliver:2008:GAP

[1010] T. Aaron Gulliver, Isaiah Makwakwa, and Ulrich Speidel. On the generation of aperiodic and periodic necklaces via T -augmentation. Fundamenta Informaticae, 83(1–2):91–107, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gossel:2008:NLS

[1011] Michael Gössel and Egor Sogomonyan. A non-linear split error detection code. Fundamenta Informaticae, 83(1–2):109–115, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hay:2008:SC

[1012] Nicholas J. Hay. Simulation complexity. Fundamenta Informaticae, 83(1– 2):117–140, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hertling:2008:CTP

[1013] Peter Hertling and Christoph Spandl. Computability theoretic properties of the entropy of gap shifts. Fundamenta Informaticae, 83(1–2):141–157, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jain:2008:PLI

[1014] Sanjay Jain, Frank Stephan, and Ye Nan. Prescribed learning of indexed families. Fundamenta Informaticae, 83(1–2):159–175, January 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ryabko:2008:AKC

[1015] Boris Ryabko. Applications of Kolmogorov complexity and universal codes to nonparametric estimation of characteristics of time series. Fundamenta Informaticae, 83(1–2):177–196, January 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Selivanov:2008:CTP

[1016] Victor L. Selivanov and Klaus W. Wagner. Complexity of topological properties of regular ω-languages. Fundamenta Informaticae, 83(1–2): 197–217, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zheng:2008:HCE

[1017] Xizhong Zheng. A hierarchy of computably enumerable reals. Fundamenta Informaticae, 83(1–2):219–230, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghosh:2008:NDR

[1018] Ashish Ghosh and Mrinal Kanti Das. Non-dominated rank based sorting genetic algorithms. Fundamenta Informaticae, 83(3):231–252, August 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grimberg:2008:FOG

[1019] Benjamin Grimberg and Michael Kaminski. First-order ground nonmonotonic modal logic. Fundamenta Informaticae, 83(3):253–276, August 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laugerotte:2008:WWR

[1020] Éric Laugerotte and Djelloul Ziadi. Weighted word recognition. Fundamenta Informaticae, 83(3):277–298, August 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2008:EBI

[1021] Chia-Chen Lin and Shing-Shoung Wang. An edge-based image copy detection scheme. Fundamenta Informaticae, 83(3):299–318, August 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2008:GPV

[1022] Jen-Chang Liu and Ming-Hong Shih. Generalizations of pixel-value differencing steganography for data hiding in images. Fundamenta Informaticae, 83(3):319–335, August 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farina:2008:PPC

[1023] Fabio Farina and Alberto Dennunzio. A predator-prey cellular automaton with parasitic interactions and environmental effects. Fundamenta Informaticae, 83(4):337–353, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meghini:2008:AAI

[1024] Carlo Meghini, Yannis Tzitzikas, and Nicolas Spyratos. Abduction for accessing information sources. Fundamenta Informaticae, 83(4):355–387, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2008:ASB

[1025] Daniel Simson and Mariusz Wojewódzki. An algorithmic solution of a Birkhoff type problem. Fundamenta Informaticae, 83(4):389–410, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Subramanian:2008:AGC

[1026] K. G. Subramanian, Do Long Van, P. Helen Chandra, and Nghiem Do Quyen. Array grammars with contextual operations. Fundamenta Informaticae, 83(4):411–428, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winter:2008:RAT

[1027] Michael Winter. A relation-algebraic theory of bisimulations. Fundamenta Informaticae, 83(4):429–449, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2008:MLP

[1028] Philippe Balbiani, Petar Iliev, and Dimiter Vakarelov. A modal logic for Pawlak’s approximation spaces with rough cardinality n. Fundamenta Informaticae, 83(4):451–464, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:AWC

[1029] Anonymous. Australasian workshop on combinatorial algorithms. Fundamenta Informaticae, 84(1):??, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amir:2008:PEC

[1030] Amihood Amir, Avivit Levy, and Liron Reuveni. The practical efficiency of convolutions in pattern matching algorithms. Fundamenta Informaticae, 84(1):1–15, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Roh:2008:EMA

[1031] Kangho Roh, Maxime Crochemore, Costas S. Iliopoulos, and Kunsoo Park. External memory algorithms for string problems. Fundamenta Informaticae, 84(1):17–32, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2008:ACR

[1032] Hui Zhang, Qing Guo, and Costas S. Iliopoulos. Algorithms for computing the λ-regularities in strings. Fundamenta Informaticae, 84(1):33–49, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kelarev:2008:ABC

[1033] Andrei Kelarev. An algorithm for BCH codes extended with finite state automata. Fundamenta Informaticae, 84(1):51–60, September 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Luo:2008:SMC

[1034] Hongwei Luo and Kathy Horadam. Simulation and measurement of complex network growth: the interplay of uniform and preferential attachment. Fundamenta Informaticae, 84(1):61–79, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Park:2008:IBC

[1035] Eunhui Park and Kunsoo Park. An improved Boolean circuit for maximum matching in a convex bipartite graph. Fundamenta Informaticae, 84 (1):81–97, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lefevre:2008:PRH

[1036] James G. Lefevre, Diane Donovan, and Aleš Drápal. Permutation representation of 3 and 4-homogeneous Latin bitrades. Fundamenta Informaticae, 84(1):99–110, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bueno:2008:PTM

[1037] Orestes Bueno and Prabhu Manyem. Polynomial-time maximisation classes: Syntactic hierarchy. Fundamenta Informaticae, 84(1):111–133, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Weng:2008:EIQ

[1038] J. F. Weng, J. MacGregor Smith, M. Brazil, and D. A. Thomas. Equivalence, indicators, quasi-indicators and optimal Steiner topologies on four points in space. Fundamenta Informaticae, 84(1):135–149, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:LP

[1039] Anonymous. Logic for pragmatics. Fundamenta Informaticae, 84(2):??, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Asher:2008:TDT

[1040] Nicholas Asher. A type driven theory of predication with complex types. Fundamenta Informaticae, 84(2):151–183, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Biasi:2008:TAP

[1041] Corrado Biasi and Federico Aschieri. A term assignment for polarized biintuitionistic logic and its strong normalization. Fundamenta Informaticae, 84(2):185–205, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Oliva:2008:KRI

[1042] Paulo Oliva and Thomas Streicher. On Krivine’s realizability interpretation of classical second-order arithmetic. Fundamenta Informaticae, 84(2):207–220, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Power:2008:CTS

[1043] John Power and Miki Tanaka. Category theoretic semantics for typed binding signatures with recursion. Fundamenta Informaticae, 84(2):221– 240, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Raffalli:2008:RAC

[1044] Christophe Raffalli and Frédéric Ruyer. Realizability of the axiom of choice in HOL. (an analysis of Krivine’s work). Fundamenta Informaticae, 84(2):241–258, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

White:2008:DRA

[1045] Graham White. Davidson and Reiter on actions. Fundamenta Informaticae, 84(2):259–289, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:2008:SAP

[1046] Alexander Clark and Chris Watkins. Some alternatives to Parikh matrices using string kernels. Fundamenta Informaticae, 84(3–4):291–303, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Droste:2008:MVM

[1047] Manfred Droste, Werner Kuich, and George Rahonis. Multi-valued MSO logics OverWords and trees. Fundamenta Informaticae, 84(3–4):305–327, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamzeh:2008:NAL

[1048] Ali Hamzeh and Adel Rahmani. A new architecture for learning classifier systems to solve POMDP problems. Fundamenta Informaticae, 84(3–4): 329–351, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huang:2008:NCK

[1049] Hui-Feng Huang and Chin-Chen Chang. A novel cryptographic key assignment scheme with ID-based access control in a hierachy. Fundamenta Informaticae, 84(3–4):353–361, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ishdorj:2008:CEI

[1050] Tseren-Onolt Ishdorj, Remco Loos, and Ion Petre. Computational efficiency of intermolecular gene assembly. Fundamenta Informaticae, 84 (3–4):363–373, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kazda:2008:CRS

[1051] Alexandr Kazda. The chain relation in sofic subshifts. Fundamenta Informaticae, 84(3–4):375–390, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2008:AEB

[1052] Ching-Chiuan Lin, Nien-Lin Hsueh, and Wen-Hsiang Shen. Adaptively embedding binary data in an image. Fundamenta Informaticae, 84(3–4): 391–402, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2008:RRE

[1053] Yong Liu, Congfu Xu, Qiong Zhang, and Yunhe Pan. Rough rule extracting from various conditions: Incremental and approximate approaches for inconsistent data. Fundamenta Informaticae, 84(3–4):403–427, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Patra:2008:CDR

[1054] Swarnajyoti Patra, Susmita Ghosh, and Ashish Ghosh. Change detection of remote sensing images with semi-supervised multilayer perceptron. Fundamenta Informaticae, 84(3–4):429–442, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ranalter:2008:SAL

[1055] Kurt Ranalter. A semantic analysis of a logic for pragmatics with assertions, obligations, and causal implication. Fundamenta Informaticae, 84(3–4):443–470, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2008:FSB

[1056] Sriparna Saha and Sanghamitra Bandyopadhyay. Fuzzy symmetry based real-coded genetic clustering technique for automatic pixel classification in remote sensing imagery. Fundamenta Informaticae, 84(3–4):471–492, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:CSP

[1057] Anonymous. Concurrency specification and programming (CS&P). Fundamenta Informaticae, 85(1–4):??, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mirkowska:2008:ALS

[1058] Grażyna Mirkowska, Andrzej Salwicki, and Oskar Swida. Algorithmic logic + SpecVer = the methodology for high integrity programming — invited paper. Fundamenta Informaticae, 85(1–4):1–17, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bednarczyk:2008:MFA

[1059] Marek A. Bednarczyk, Wiesław Pawłowski, Luca Bernardinello, Lucia Pomello, and Tomasz Borzyszkowski. A multi-facet approach to dynamic agent systems. Fundamenta Informaticae, 85(1–4):19–33, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2008:MPP

[1060] Marco Bellia and M. Eugenia Occhiuto. Methods as parameters: A preprocessing approach to higher order in Java. Fundamenta Informaticae, 85(1–4):35–50, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Budzynska:2008:LRA

[1061] Katarzyna Budzyńska and Magdalena Kacprzak. A logic for reasoning about persuasion. Fundamenta Informaticae, 85(1–4):51–65, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chrzastowski-Wachtel:2008:TDS

[1062] Piotr Chrzastowski-Wachtel, Paweł Findeisen, and Grzegorz Wolny. Time distribution in structural workflow nets. Fundamenta Informaticae, 85(1–4):67–87, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2008:BAN

[1063] Ludwik Czaja. Boolean algebra of nets, their synthesis and analysis. Fundamenta Informaticae, 85(1–4):89–95, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delimata:2008:FSA

[1064] Pawel Delimata and Zbigniew Suraj. Feature selection algorithm for multiple classifier systems: A hybrid approach. Fundamenta Informaticae, 85(1–4):97–110, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farwer:2008:PNC

[1065] Berndt Farwer, Matthias Jantzen, Manfred Kudlek, Heiko Rölke, and Georg Zetzsche. Petri net controlled finite automata. Fundamenta Informaticae, 85(1–4):111–121, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gohring:2008:CBW

[1066] Daniel Göhring, Heinrich Mellmann, Kataryna Gerasymova, and Hans-Dieter Burkhard. Constraint based world modeling. Fundamenta Informaticae, 85(1–4):123–137, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2008:SFS

[1067] Anna Gomolińska. Satisfiability of formulas from the standpoint of object classification: The RST approach. Fundamenta Informaticae, 85(1–4): 139–153, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gora:2008:CBP

[1068] Grzegorz Góra, Piotr Kruczek, Andrzej Skowron, Jan G. Bazan, Stanisława Bazan-Socha, and Jacek J. Pietrzyk. Case-based planning of treatment of infants with respiratory failure. Fundamenta Informaticae, 85(1–4):155–172, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2008:PIF

[1069] Damas P. Gruska. Probabilistic information flow security. Fundamenta Informaticae, 85(1–4):173–187, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2008:SML

[1070] Damas P. Gruska, Andrea Maggiolo-Schettini, and Paolo Milazzo. Security in a model for long-running transactions. Fundamenta Informaticae, 85(1–4):189–203, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzymala-Busse:2008:ASL

[1071] Jerzy W. Grzymala-Busse and Wojciech Rzasa. Approximation space and LEM2-like algorithms for computing local coverings. Fundamenta Informaticae, 85(1–4):205–217, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2008:THP

[1072] Kees van Hee, Alexander Serebrenik, and Natalia Sidorova. Token history Petri nets. Fundamenta Informaticae, 85(1–4):219–234, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janowska:2008:TIL

[1073] Agata Janowska, Paweł Janowski, and Dobiesław Wróblewski. Translation of intermediate language to timed automata with discrete data. Fundamenta Informaticae, 85(1–4):235–248, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jankowski:2008:ODC

[1074] Andrzej Jankowski, James F. Peters, Andrzej Skowron, and Jarosław Stepaniuk. Optimization in discovery of compound granules. Fundamenta Informaticae, 85(1–4):249–265, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jantzen:2008:LCD

[1075] Matthias Jantzen, Manfred Kudlek, and Georg Zetzsche. Language classes defined by concurrent finite automata. Fundamenta Informaticae, 85(1–4):267–280, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jolkowska:2008:TEB

[1076] Joanna Jółkowska and Edward Ochmański. On trace-expressible behaviour of Petri nets. Fundamenta Informaticae, 85(1–4):281–295, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jungel:2008:MBS

[1077] Matthias Jüngel and Heinrich Mellmann. Memory-based stateestimation. Fundamenta Informaticae, 85(1–4):297–311, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2008:VMC

[1078] Magdalena Kacprzak, Wojciech Nabiałek, Artur Niewiadomski, Wojciech Penczek, Agata Półrola, Maciej Szreter, Bożena Woźna, and Andrzej Zbrzezny. VerICS 2007 — a model checker for knowledge and real-time. Fundamenta Informaticae, 85(1–4):313–328, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2008:LPZ

[1079] Michael Köhler-Bußmeier and Manfred Kudlek. Linear properties of zerosafe nets with debit tokens. Fundamenta Informaticae, 85(1–4):329–342, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Langmaack:2008:DAI

[1080] Hans Langmaack, Andrzej Salwicki, and Marek Warpechowski. A deterministic algorithm for identifying direct superclasses in Java. Fundamenta Informaticae, 85(1–4):343–357, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomuscio:2008:LFM

[1081] Alessio Lomuscio and Wojciech Penczek. LDYIS: a framework for model checking security protocols. Fundamenta Informaticae, 85(1–4):359–375, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Middelkoop:2008:SVI

[1082] Ronald Middelkoop, Cornelis Huizing, Ruurd Kuiper, and Erik J. Luit. Specification and verification of invariants by exploiting layers in OO designs. Fundamenta Informaticae, 85(1–4):377–398, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mikulski:2008:PRM

[1083] Łukasz Mikulski. Projection representation of Mazurkiewicz traces. Fundamenta Informaticae, 85(1–4):399–408, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nabialek:2008:TTP

[1084] Wojciech Nabiałek, Agata Janowska, and Paweł Janowski. Translation of timed Promela to timed automata with discrete data. Fundamenta Informaticae, 85(1–4):409–424, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2008:SBU

[1085] Wojciech Penczek and Maciej Szreter. SAT-based unbounded model checking of timed automata. Fundamenta Informaticae, 85(1–4):425–440, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2008:SAP

[1086] Roman R. Redziejowski. Some aspects of parsing expression grammar. Fundamenta Informaticae, 85(1–4):441–451, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Strzalka:2008:PSL

[1087] Dominik Strzałka and Franciszek Grabowski. Processes in systems with limited resources in the context of non-extensive thermodynamics. Fundamenta Informaticae, 85(1–4):455–464, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Varpaaniemi:2008:MSQ

[1088] Kimmo Varpaaniemi and Leo Ojala. Modelling and simulation of quantum teleportation and dense coding using predicate/transition-nets. Fundamenta Informaticae, 85(1–4):465–479, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2008:AFD

[1089] Józef Winkowski. An algebraic framework for defining random concurrent behaviours. Fundamenta Informaticae, 85(1–4):481–496, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2008:DMI

[1090] Marcin Wolski. Distance measures induced by finite approximation spaces and approximation operators. Fundamenta Informaticae, 85(1–4): 497–512, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2008:ITE

[1091] Andrzej Zbrzezny. Improving the translation from ECTL to SAT. Fundamenta Informaticae, 85(1–4):513–531, September 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2008:TVJ

[1092] Andrzej Zbrzezny and Bożena Woźna. Towards verification of Java programs in perICS. Fundamenta Informaticae, 85(1–4):533–548, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zielosko:2008:GAA

[1093] Beata Zielosko and Marcin Piliszczuk. Greedy algorithm for attribute reduction. Fundamenta Informaticae, 85(1–4):549–561, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bijak:2008:GAA

[1094] Katarzyna Bijak. Genetic algorithms as an alternative method of parameter estimation and finding most likely sequences of states of hidden Markov chains for HMMs and hybrid HMM/ANN models. Fundamenta Informaticae, 86(1–2):1–17, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bloom:2008:PCI

[1095] Stephen L. Bloom, Zoltan Ésik, and Werner Kuich. Partial Conway and iteration semirings. Fundamenta Informaticae, 86(1–2):19–40, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Denis:2008:RSL

[1096] François Denis and Yann Esposito. On rational stochastic languages. Fundamenta Informaticae, 86(1–2):41–77, October 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2008:MCD

[1097] Liviu P. Dinu and Marius Popescu. A multi-criteria decision method based on rank distance. Fundamenta Informaticae, 86(1–2):79–91, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamzeh:2008:AAR

[1098] Ali Hamzeh and Adel Rahmani. Approximating arbitrary reinforcement signal by learning classifier systems using micro genetic algorithm. Fundamenta Informaticae, 86(1–2):93–111, October 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2008:GIH

[1099] Yu-Chen Hu, Chih-Chiang Tsou, and Bing-Hwang Su. Grayscale image hiding based on modulus function and greedy method. Fundamenta Informaticae, 86(1–2):113–126, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jonoska:2008:ISJ

[1100] Nataša Jonoska, Lila Kari, and Kalpana Mahalingam. Involution solid and join codes. Fundamenta Informaticae, 86(1–2):127–142, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2008:CAB

[1101] Pradipta Maji. On characterization of attractor basins of fuzzy multiple attractor cellular automata. Fundamenta Informaticae, 86(1–2):143–168, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nadarajah:2008:NC

[1102] Saralees Nadarajah. On the normalized cut. Fundamenta Informaticae, 86(1–2):169–173, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ramadoss:2008:MDV

[1103] Balakrishnan Ramadoss and Kannan Rajkumar. Modeling the dance video semantics using regular tree automata. Fundamenta Informaticae, 86(1–2):175–189, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Terlikowski:2008:SII

[1104] Tomasz Terlikowski. Some issues on inner structure of information. A definition of the decision making problem. Fundamenta Informaticae, 86(1–2):191–212, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tzacheva:2008:TBC

[1105] Angelina A. Tzacheva and Li-Shiang Tsay. Tree-based construction of low-cost action rules. Fundamenta Informaticae, 86(1–2):213–225, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:FSI

[1106] Anonymous. The fifth special issue on application of concurrency to system design. Fundamenta Informaticae, 86(3):??, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:2008:CCC

[1107] Javier Esparza, Petr Jančar, and Alexander Miller. On the complexity of consistency and complete state coding for signal transition graphs. Fundamenta Informaticae, 86(3):227–253, October 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Juhas:2008:CSA

[1108] Gabriel Juhás, Robert Lorenz, and Sebastian Mauser. Causal semantics of algebraic Petri nets distinguishing concurrency and Synchronicity. Fundamenta Informaticae, 86(3):255–298, October 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khomenko:2008:REC

[1109] Victor Khomenko, Agnes Madalinski, and Alex Yakovlev. Resolution of encoding conflicts by signal insertion and concurrency reduction based on STG unfoldings. Fundamenta Informaticae, 86(3):299–323, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2008:SAP

[1110] Cong Liu, Alex Kondratyev, Yosinori Watanabe, Jörg Desel, and Alberto Sangiovanni-Vincentelli. Schedulability analysis of Petri nets based on structural properties. Fundamenta Informaticae, 86(3):325–341, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seceleanu:2008:MCS

[1111] Tiberiu Seceleanu and Axel Jantsch. Modeling communication with synchronized environments. Fundamenta Informaticae, 86(3):343–369, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abdelli:2008:TBS

[1112] Abdelkrim Abdelli and Nadjib Badache. Towards building the state class graph of the TSPN model. Fundamenta Informaticae, 86(4):371–409, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mukhopadhyay:2008:UPC

[1113] Anirban Mukhopadhyay and Ujjwal Maulik. Unsupervised pixel classification in satellite imagery: A two-stage fuzzy clustering approach. Fundamenta Informaticae, 86(4):411–428, October 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chaitin:2008:HPW

[1114] Gregory Chaitin. The halting probability via Wang tiles. Fundamenta Informaticae, 86(4):429–433, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2008:HCD

[1115] Chin-Chen Chang, Yung-Chen Chou, and The Duc Kieu. High capacity data hiding for grayscale images. Fundamenta Informaticae, 86(4):435– 446, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2008:SNI

[1116] Chien-Chang Chen and Yu-Wei Chien. Sharing numerous images secretly with reduced possessing load. Fundamenta Informaticae, 86(4):447–458, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drewes:2008:BCT

[1117] Frank Drewes, Christine du Toit, Sigrid Ewert, Brink van der Merwe, and Andries van der Walt. Bag context tree grammars. Fundamenta Informaticae, 86(4):459–480, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ha:2008:DID

[1118] Minghu Ha, Witold Pedrycz, Aiquan Zhang, and Yijun Fan. A development of inclusion-degree-based rough fuzzy random sets. Fundamenta Informaticae, 86(4):481–502, October 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:SIMa

[1119] Anonymous. Special issue on membrane computing. Fundamenta Informaticae, 87(1):??, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:SIMb

[1120] Anonymous. Special issue on membrane computing. Fundamenta Informaticae, 87(1):??, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ceterchi:2008:ISN

[1121] Rodica Ceterchi and Alexandru Ioan Tomescu. Implementing sorting networks with spiking neural P systems. Fundamenta Informaticae, 87 (1):35–48, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2008:CCS

[1122] Gabriel Ciobanu and Andreas Resios. Computational complexity of simple P systems. Fundamenta Informaticae, 87(1):49–59, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leporati:2008:SSS

[1123] Alberto Leporati and Miguel A. Gutiérrez-Naranjo. Solving SUBSET SUM by spiking neural P systems with pre-computed resources. Fundamenta Informaticae, 87(1):61–77, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zandron:2008:CEP

[1124] Claudio Zandron, Alberto Leporati, Claudio Ferretti, Giancarlo Mauri, and Mario J. Pérez-Jiménez. On the computational efficiency of polarizationless recognizer P systems with strong division and dissolution. Fundamenta Informaticae, 87(1):79–91, November 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2008:QIE

[1125] Ge-Xiang Zhang, Marian Gheorghe, and Chao-Zhong Wu. A quantuminspired evolutionary algorithm based on P systems for knapsack problem. Fundamenta Informaticae, 87(1):93–116, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2008:SUS

[1126] Xingyi Zhang, Xiangxiang Zeng, and Linqiang Pan. Smaller universal spiking neural P systems. Fundamenta Informaticae, 87(1):117–136, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:CAT

[1127] Anonymous. Cellular automata: Theory and application in artificial intelligence. Fundamenta Informaticae, 87(2):??, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ganguly:2008:ECS

[1128] Niloy Ganguly, Biplab K. Sikdar, and P. Pal Chaudhuri. Exploring cycle structures of additive cellular automata. Fundamenta Informaticae, 87(2):137–154, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umeo:2008:NTO

[1129] Hiroshi Umeo and Hiroki Uchino. A new time-optimum synchronization algorithm for rectangle arrays. Fundamenta Informaticae, 87(2):155–164, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Flocchini:2008:RVC

[1130] Paola Flocchini and Vladimir Cezar. Radial view of continuous cellular automata. Fundamenta Informaticae, 87(2):165–183, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Basu:2008:DTL

[1131] Subhasree Basu and Sumita Basu. Different types of linear fuzzy cellular automata and their applications. Fundamenta Informaticae, 87(2):185– 205, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bandini:2008:NGF

[1132] Stefania Bandini, Leonardo Vanneschi, Andrew Wuensche, and Alessandro Bahgat Shehata. A neuro-genetic framework for pattern recognition in complex systems. Fundamenta Informaticae, 87(2):207–226, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moitra:2008:CAB

[1133] Debashis Moitra, S. Dhar, J. M. Mallick, S. Sadhu, and Biplab K. Sikdar. Cellular automata based design of cost optimal steel building frames. Fundamenta Informaticae, 87(2):227–245, November 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rongo:2008:LFH

[1134] Rocco Rongo, William Spataro, Donato D’Ambrosio, Maria Vittoria Avolio, Giuseppe A. Trunfio, and Salvatore Di Gregorio. Lava flow hazard evaluation through cellular automata and genetic algorithms: an application to Mt. Etna volcano. Fundamenta Informaticae, 87(2):247–267, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Domaratzki:2008:CDB

[1135] Michael Domaratzki. Characterizing DNA bond shapes using trajectories. Fundamenta Informaticae, 87(3–4):269–285, November 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fokkink:2008:TBB

[1136] Wan Fokkink, Jun Pang, and Anton Wijs. Is timed branching bisimilarity a congruence indeed? Fundamenta Informaticae, 87(3–4):287–311, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2008:NBR

[1137] Yu-Chen Hu, Piyu Tsai, and Chun-Chi Lo. New bit reduction of vector quantization using block prediction and relative addressing. Fundamenta Informaticae, 87(3–4):313–329, November 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Juhas:2008:CPS

[1138] Gabriel Juhás, Robert Lorenz, and Sebastian Mauser. Complete process semantics of Petri nets. Fundamenta Informaticae, 87(3–4):331–365, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Macia:2008:SME

[1139] Hermenegilda Macià, Valentín Valero, Fernando Cuartero, and M. Carmen Ruiz. sPBC: A Markovian extension of Petri box calculus with immediate multiactions. Fundamenta Informaticae, 87(3–4):367–406, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Masopust:2008:DCP

[1140] Tomáš Masopust and Alexander Meduna. On descriptional complexity of partially parallel grammars. Fundamenta Informaticae, 87(3–4):407–415, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitra:2008:NPA

[1141] Suman K. Mitra, Malay K. Kundu, C. A. Murthy, Bhargab B. Bhattacharya, and Tinku Acharya. A new probabilistic approach for fractal based image compression. Fundamenta Informaticae, 87(3–4):417–433, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Oprocha:2008:AAD

[1142] Piotr Oprocha. Algorithmic approach to Devaney chaos in shift spaces. Fundamenta Informaticae, 87(3–4):435–446, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsai:2008:FVC

[1143] Piyu Tsai, Yu-Chen Hu, and Hsiu-Lien Yeh. Fast VQ codebook generation method using codeword stability check and finite state concept. Fundamenta Informaticae, 87(3–4):447–463, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xiao:2008:AUS

[1144] Kai Xiao, Sooi Hock Ho, and Aboul ella Hassanien. Automatic unsupervised segmentation methods for MRI based on modified fuzzy C-means. Fundamenta Informaticae, 87(3–4):465–481, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yoon:2008:RUP

[1145] Eun-Jun Yoon and Kee-Young Yoo. Robust user password change scheme based on the elliptic curve cryptosystem. Fundamenta Informaticae, 87 (3–4):483–492, November 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aceto:2008:CTB

[1146] Luca Aceto, Wan Fokkink, and Anna Ingolfsdottir. A cancellation theorem for BCCSP. Fundamenta Informaticae, 88(1–2):1–21, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2008:ADE

[1147] Chin-Chen Chang, Chia-Chen Lin, and Yi-Hui Chen. Adaptive data embedding scheme based on similar and dissimilar clustering strategies. Fundamenta Informaticae, 88(1–2):23–45, January 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chou:2008:LPD

[1148] Yung-Chen Chou, Chin-Chen Chang, and Kuan-Ming Li. A large payload data embedding technique for color images. Fundamenta Informaticae, 88 (1–2):47–61, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Damiani:2008:ATE

[1149] Ferruccio Damiani, Elena Giachino, and Paola Giannini. Alias types and effects for “environment-aware” computations. Fundamenta Informaticae, 88(1–2):63–107, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haghighi:2008:NC

[1150] Hassan Haghighi and Seyed-Hassan Mirian-Hosseinabadi. Nondeterminism in constructive Z. Fundamenta Informaticae, 88(1–2):109–134, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hashemi:2008:CSF

[1151] Sattar Hashemi, Mohammadreza Kangavari, and Ying Yang. Class specific fuzzy decision trees for mining high speed data streams. Fundamenta Informaticae, 88(1–2):135–160, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2008:SOF

[1152] Pradipta Maji, Malay K. Kundu, and Bhabatosh Chanda. Second order fuzzy measure and weighted co-occurrence matrix for segmentation of brain MR images. Fundamenta Informaticae, 88(1–2):161–176, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vyas:2008:GTI

[1153] Vibha S. Vyas and Priti P. Rege. Geometric transform invariant texture analysis based on modified Zernike moments. Fundamenta Informaticae, 88(1–2):177–192, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xiong:2008:ICS

[1154] Hu Xiong, Zhiguang Qin, and Fagen Li. An improved certificateless signature scheme secure in the standard model. Fundamenta Informaticae, 88(1–2):193–206, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2008:NPE

[1155] Zhanbo Yang. A new proof on embedding the category of proximity spaces into the category of nearness spaces. Fundamenta Informaticae, 88(1–2):207–223, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:P

[1156] Anonymous. Preface. Fundamenta Informaticae, 88(3):??, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boyer:2008:CEA

[1157] Marc Boyer and Olivier H. Roux. On the compared expressiveness of arc, place and transition time Petri nets. Fundamenta Informaticae, 88(3): 225–249, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lazic:2008:NTW

[1158] Ranko Lazić, Tom Newcomb, Joël Ouaknine, A. W. Roscoe, and James Worrell. Nets with tokens which carry data. Fundamenta Informaticae, 88(3):251–274, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ganty:2008:MPF

[1159] Pierre Ganty, Jean-François Raskin, and Laurent Van Begin. From many places to few: Automatic abstraction refinement for Petri nets. Fundamenta Informaticae, 88(3):275–305, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:2008:SEN

[1160] Maciej Koutny and Marta Pietkiewicz-Koutny. Synthesis of elementary net systems with context arcs and localities. Fundamenta Informaticae, 88(3):307–328, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rosa-Velardo:2008:NCV

[1161] Fernando Rosa-Velardo and David de Frutos-Escrig. Name creation vs. replication in Petri net systems. Fundamenta Informaticae, 88(3):329– 356, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Billington:2008:CPN

[1162] Jonathan Billington and Somsak Vanit-Anunchai. Coloured Petri net modelling of an evolving Internet standard: the datagram congestion control protocol. Fundamenta Informaticae, 88(3):357–385, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2008:WPI

[1163] Kees van Hee, Alexander Serebrenik, Natalia Sidorova, and Wil van der Aalst. Working with the past: Integrating history in Petri nets. Fundamenta Informaticae, 88(3):387–409, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:ACS

[1164] Anonymous. Application of concurrency to system design, the sixth special issue. Fundamenta Informaticae, 88(4):??, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beal:2008:HCT

[1165] Frédéric Béal, Tomohiro Yoneda, and Chris J. Myers. Hazard checking of timed asynchronous circuits revisited. Fundamenta Informaticae, 88(4): 411–435, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenthum:2008:SPN

[1166] Robin Bergenthum, Jörg Desel, Robert Lorenz, and Sebastian Mauser. Synthesis of Petri nets from finite partial languages. Fundamenta Informaticae, 88(4):437–468, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boucheneb:2008:MET

[1167] Hanifa Boucheneb and Hind Rakkay. A more efficient time Petri net state space abstraction useful to model checking timed linear properties. Fundamenta Informaticae, 88(4):469–495, December 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cassez:2008:FDS

[1168] Franck Cassez and Stavros Tripakis. Fault diagnosis with static and dynamic observers. Fundamenta Informaticae, 88(4):497–540, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khomenko:2008:ODA

[1169] Victor Khomenko, Mark Schaefer, and Walter Vogler. Outputdeterminacy and asynchronous circuit synthesis. Fundamenta Informaticae, 88(4):541–579, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sokolov:2008:ASD

[1170] Danil Sokolov, Ivan Poliakov, and Alex Yakovlev. Analysis of static data flow structures. Fundamenta Informaticae, 88(4):581–610, December 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:PMR

[1171] Anonymous. Progress on multi-relational data mining. Fundamenta Informaticae, 89(1):??, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeKnijf:2008:ECD

[1172] Jeroen De Knijf and Ad Feelders. An experimental comparison of different inclusion relations in frequent tree mining. Fundamenta Informaticae, 89(1):1–22, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esposito:2008:MDR

[1173] Floriana Esposito, Nicola Di Mauro, Teresa M. A. Basile, and Stefano Ferilli. Multi-dimensional relational sequence mining. Fundamenta Informaticae, 89(1):23–43, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fonseca:2008:CHS

[1174] Nuno A. Fonseca, Rui Camacho, Ricardo Rocha, and Vítor Santos Costa. Compile the hypothesis space: Do it once, use it often. Fundamenta Informaticae, 89(1):45–67, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guo:2008:LSC

[1175] Hongyu Guo and Herna L. Viktor. Learning from skewed class multirelational databases. Fundamenta Informaticae, 89(1):69–94, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuzelka:2008:RSE

[1176] Ondřej Kuželka and Filip Železný. A restarted strategy for efficient subsumption testing. Fundamenta Informaticae, 89(1):95–109, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Landwehr:2008:RTB

[1177] Niels Landwehr, Bernd Gutmann, Ingo Thon, Luc De Raedt, and Matthai Philipose. Relational transformation-based tagging for activity recognition. Fundamenta Informaticae, 89(1):111–129, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meert:2008:LGC

[1178] Wannes Meert, Jan Struyf, and Hendrik Blockeel. Learning ground CPlogic theories by leveraging Bayesian network learning techniques. Fundamenta Informaticae, 89(1):131–160, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cheng:2008:BAF

[1179] Chien-Fu Cheng, Shu-Ching Wang, and Tyne Liang. Byzantine agreement & fault diagnosis agreement in dynamic ad-hoc environment. Fundamenta Informaticae, 89(2–3):161–187, January 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dembczynski:2008:EPW

[1180] Krzysztof Dembczyński, Wojciech Kotłowski, and Marcin Sydow. Effective prediction of Web user behaviour with user-level models. Fundamenta Informaticae, 89(2–3):189–206, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fulop:2008:PMT

[1181] Zoltán Fülöp and Loránd Muzamel. Pebble macro tree transducers with strong pebble handling. Fundamenta Informaticae, 89(2–3):207–257, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grahne:2008:PRP

[1182] Gösta Grahne, Alex Thomo, and William W. Wadge. Preferential regular path queries. Fundamenta Informaticae, 89(2–3):259–288, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurdzinski:2008:BCG

[1183] Tomasz Jurdziński. The Boolean closure of growing context-sensitive languages. Fundamenta Informaticae, 89(2–3):289–305, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leivant:2008:LUM

[1184] Daniel Leivant. Logical undecidabilities made easy. Fundamenta Informaticae, 89(2–3):307–312, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nepomniaschaya:2008:EII

[1185] Anna Nepomniaschaya. Efficient implementation of the Italiano algorithms for updating the transitive closure on associative parallel processors. Fundamenta Informaticae, 89(2–3):313–329, January 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2008:PSL

[1186] Indranil Saha, Bhargab B. Bhattacharya, Sheng Zhang, and Sharad C. Seth. Planar straight-line embedding of double-tree scan architecture on a rectangular grid. Fundamenta Informaticae, 89(2–3):331–344, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Svrcek:2008:ASS

[1187] Filip Švrček. On the axiomatic system of SBL a̸lgebras. Fundamenta Informaticae, 89(2–3):345–368, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2008:FSE

[1188] Anonymous. Fundamentals of software engineering 2007: Selected contributions. Fundamenta Informaticae, 89(4):??, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bruckner:2008:SA

[1189] Ingo Brückner, Klaus Dräger, Bernd Finkbeiner, and Heike Wehrheim. Slicing abstractions. Fundamenta Informaticae, 89(4):369–392, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dashti:2008:NDP

[1190] Mohammad Torabi Dashti, Srijith Krishnan Nair, and Hugo Jonker. Nuovo DRM paradiso: Designing a secure, verified, fair exchange DRM scheme. Fundamenta Informaticae, 89(4):393–417, January 2008. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abdulla:2008:UAO

[1191] Parosh Aziz Abdulla, Johann Deneux, Joël Ouaknine, Karin Quaas, and James Worrell. Universality analysis for one-clock timed automata. Fundamenta Informaticae, 89(4):419–450, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bravetti:2008:FTC

[1192] Mario Bravetti and Gianluigi Zavattaro. A foundational theory of contracts for multi-party service composition. Fundamenta Informaticae, 89 (4):451–478, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bonchi:2008:UBE

[1193] Filippo Bonchi, Antonio Brogi, Sara Corfini, and Fabio Gadducci. On the use of behavioural equivalences for Web services’ development. Fundamenta Informaticae, 89(4):479–510, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Montangero:2008:LBC

[1194] Carlo Montangero, Stephan Reiff-Marganiec, and Laura Semini. Logicbased conflict detection for distributed policies. Fundamenta Informaticae, 89(4):511–538, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hashemian:2008:LRA

[1195] Seyyed Vahid Hashemian and Farhad Mavaddat. A logical reasoning approach to automatic composition of stateless components. Fundamenta Informaticae, 89(4):539–577, January 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2009:ECE

[1196] Jean-Marc Champarnaud, Faissal Ouardi, and Djelloul Ziadi. An efficient computation of the equation K-automaton of a regular K-expression. Fundamenta Informaticae, 90(1–2):1–16, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2009:FDI

[1197] Chin-Chen Chang and Yung-Chen Chou. A fragile digital image authentication scheme inspired by wet paper codes. Fundamenta Informaticae, 90(1–2):17–26, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cirulis:2009:RSA

[1198] Jānis Cīrulis. Rough set algebras as description domains. Fundamenta Informaticae, 90(1–2):27–41, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferilli:2009:GSF

[1199] S. Ferilli, T. M. A. Basile, M. Biba, N. Di Mauro, and F. Esposito. A general similarity framework for Horn clause logic. Fundamenta Informaticae, 90(1–2):43–66, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giesen:2009:AS

[1200] Joachim Giesen, Eva Schuberth, and Miloš Stojaković. Approximate sorting. Fundamenta Informaticae, 90(1–2):67–72, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorawski:2009:IST

[1201] Marcin Gorawski and Adam Dyga. Indexing of spatio-temporal telemetric data based on adaptive multi-dimensional bucket index. Fundamenta Informaticae, 90(1–2):73–86, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halas:2009:NAS

[1202] Radomír Halaš. A note on axiom system for SBL^{¬}-algebras. Fundamenta Informaticae, 90(1–2):87–92, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2009:NSC

[1203] Yo-Sub Han, Kai Salomaa, and Derick Wood. Nondeterministic state complexity of basic operations for prefix-free regular languages. Fundamenta Informaticae, 90(1–2):93–106, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2009:SSB

[1204] Sriparna Saha and Sanghamitra Bandyopadhyay. Some symmetry based classifiers. Fundamenta Informaticae, 90(1–2):107–123, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:NWA

[1205] Chung-Chuan Wang, Chin-Chen Chang, and Jinn-Ke Jan. Novel watermarking authentication schemes for binary images based on dual-pair block pixel patterns. Fundamenta Informaticae, 90(1–2):125–155, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:CLS

[1206] Xia Wang and Wenxiu Zhang. Concept lattices of subcontexts of a context. Fundamenta Informaticae, 90(1–2):157–169, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2009:ESV

[1207] Chih-Chia Yao, Pao-Ta Yu, and Ruo-Wei Hung. Extractive support vector algorithm on support vector machines for image restoration. Fundamenta Informaticae, 90(1–2):171–190, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2009:NCS

[1208] Leyou Zhang, Yupu Hu, and Qing Wu. New construction of short hierarchical ID-based signature in the standard model. Fundamenta Informaticae, 90(1–2):191–201, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:PCIa

[1209] Yingxu Wang, Du Zhang, and Shusaku Tsumoto. Preface: Cognitive informatics, cognitive computing, and their denotational mathematical foundations (i). Fundamenta Informaticae, 90(3):??, March 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:DCI

[1210] Yingxu Wang, Witold Kinsner, James A. Anderson, Du Zhang, Yiyu Yao, Philip Sheu, Jeffrey Tsai, Witold Pedrycz, Jean-Claude Latombe, Lotfi A. Zadeh, Dilip Patel, and Christine Chan. A doctrine of cognitive informatics (CI). Fundamenta Informaticae, 90(3):203–228, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

tenBosch:2009:CML

[1211] Louis ten Bosch, Lou Boves, Hugo Van Hamme, and Roger K. Moore. A computational model of language acquisition: the emergence of words. Fundamenta Informaticae, 90(3):229–249, March 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2009:SID

[1212] Shusaku Tsumoto and Shoji Hirano. Statistical independence and determinants in a contingency table — interpretation of Pearson residuals based on linear algebra. Fundamenta Informaticae, 90(3):251–267, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peng:2009:DIE

[1213] Jun Peng, Du Zhang, and Xiaofeng Liao. A digital image encryption algorithm based on hyper-chaotic cellular neural network. Fundamenta Informaticae, 90(3):269–282, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:PDM

[1214] Yingxu Wang. Paradigms of denotational mathematics for cognitive informatics and cognitive computing. Fundamenta Informaticae, 90(3): 283–303, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Llana:2009:TSR

[1215] Luis Llana and Manuel Núuñez. Testing semantics for RTPA. Fundamenta Informaticae, 90(3):305–335, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamandi:2009:CMB

[1216] Ali Kamandi and Jafar Habibi. A comparison of metric-based and empirical approaches for cognitive analysis of modeling languages. Fundamenta Informaticae, 90(3):337–352, March 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:PCIb

[1217] Yingxu Wang, Du Zhang, and Shusaku Tsumoto. Preface: Cognitive informatics, cognitive computing, and their denotational mathematical foundations (II). Fundamenta Informaticae, 90(4):??, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:FSN

[1218] Yingxu Wang. A formal syntax of natural languages and the deductive grammar. Fundamenta Informaticae, 90(4):353–368, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abe:2009:IAC

[1219] Hidenao Abe and Shusaku Tsumoto. Investigating accuracies of classifications for randomized imbalanced class distributions. Fundamenta Informaticae, 90(4):369–378, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kwolek:2009:OSV

[1220] Bogdan Kwolek. Object segmentation in video via graph cut built on superpixels. Fundamenta Informaticae, 90(4):379–393, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:DOD

[1221] Guoyin Wang and Yan Wang. 3DM: Domain-oriented data-driven data mining. Fundamenta Informaticae, 90(4):395–426, December 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2009:CMT

[1222] Shusaku Tsumoto and Shoji Hirano. Contingency matrix theory II: Degree of dependence as granularity. Fundamenta Informaticae, 90(4):427– 442, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2009:TFM

[1223] Yingxu Wang. Toward formal models of the theoretical framework of fundamental economics. Fundamenta Informaticae, 90(4):443–459, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:Pa

[1224] Anonymous. Preface. Fundamenta Informaticae, 91(1):??, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2009:NEP

[1225] Artiom Alhazov, Carlos Martín-Vide, Bianca Truthe, Jürgen Dassow, and Yurii Rogozhin. On networks of evolutionary processors with nodes of two types. Fundamenta Informaticae, 91(1):1–15, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2009:PHM

[1226] Artiom Alhazov, Marion Oswald, Rudolf Freund, and Sergey Verlan. Partial halting and minimal parallelism based on arbitrary rule partitions. Fundamenta Informaticae, 91(1):17–34, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baumeister:2009:SPR

[1227] Dorothea Baumeister and Jörg Rothe. Satisfiability parsimoniously reduces to the Tantrix^{T M} rotation puzzle problem. Fundamenta Informaticae, 91(1):35–51, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burgin:2009:UTM

[1228] Mark Burgin and Eugene Eberbach. Universality for Turing machines, inductive Turing machines and evolutionary algorithms. Fundamenta Informaticae, 91(1):53–77, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fisher:2009:SM

[1229] John Fisher and Marc Bezem. Skolem machines. Fundamenta Informaticae, 91(1):79–103, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruber:2009:MSH

[1230] Hermann Gruber, Markus Holzer, and Martin Kutrib. More on the size of Higman–Haines sets: Effective constructions. Fundamenta Informaticae, 91(1):105–121, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Neary:2009:FSU

[1231] Turlough Neary and Damien Woods. Four small universal Turing machines. Fundamenta Informaticae, 91(1):123–144, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Subramanian:2009:ASC

[1232] K. G. Subramanian, Rosihan M. Ali, Atulya K. Nagar, and Maurice Margenstern. Array P systems and t-communication. Fundamenta Informaticae, 91(1):145–159, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umeo:2009:SFS

[1233] Hiroshi Umeo and Takashi Yanagihara. A small five-state non-optimum-time solution to the firing squad synchronization problem. Fundamenta Informaticae, 91(1):161–178, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Woods:2009:SSW

[1234] Damien Woods and Turlough Neary. Small semi-weakly universal Turing machines. Fundamenta Informaticae, 91(1):179–195, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:Pb

[1235] Anonymous. Preface. Fundamenta Informaticae, 91(2):??, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arrighi:2009:IUO

[1236] Pablo Arrighi, Renan Fargetton, and Zizhu Wang. Intrinsically universal one-dimensional quantum cellular automata in two flavours. Fundamenta Informaticae, 91(2):197–230, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertoni:2009:CUT

[1237] Alberto Bertoni, Massimiliano Goldwurm, and Violetta Lonati. The complexity of unary tiling recognizable picture languages: Nondeterministic and unambiguous cases. Fundamenta Informaticae, 91(2):231–249, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bliudze:2009:MCS

[1238] Simon Bliudze and Daniel Krob. Modelling of complex systems: Systems as dataflow machines. Fundamenta Informaticae, 91(2):251–274, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bozga:2009:FPC

[1239] Marius Bozga, Radu Iosif, and Yassine Lakhnech. Flat parametric counter automata. Fundamenta Informaticae, 91(2):275–303, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2009:HUP

[1240] Olivier Finkel. Highly undecidable problems about recognizability by tiling systems. Fundamenta Informaticae, 91(2):305–323, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Friedl:2009:QTH

[1241] Katalin Friedl, Miklos Santha, Frédéric Magniez, and Pranab Sen. Quantum testers for hidden group properties. Fundamenta Informaticae, 91 (2):325–340, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Istrate:2009:DSB

[1242] Gabriel Istrate. On the dynamics of social balance on general networks (with an application to XOR-SAT). Fundamenta Informaticae, 91(2): 341–356, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lippi:2009:UHI

[1243] Sylvain Lippi. Universal hard interaction for clockless computation. Fundamenta Informaticae, 91(2):357–394, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lisitsa:2009:CPQ

[1244] Alexei Lisitsa and Igor Potapov. On the computational power of querying the history. Fundamenta Informaticae, 91(2):395–409, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stewart:2009:PSQ

[1245] Iain A. Stewart. Program schemes, queues, the recursive spectrum and zero-one laws. Fundamenta Informaticae, 91(2):411–435, April 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

terBeek:2009:AIS

[1246] Maurice H. ter Beek and Jetty Kleijn. Associativity of infinite synchronized shuffles and team automata. Fundamenta Informaticae, 91(3–4): 437–461, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernot:2009:BPB

[1247] Gilles Bernot and Fariza Tahi. Behaviour preservation of a biological regulatory network when embedded into a larger network. Fundamenta Informaticae, 91(3–4):463–485, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhowmick:2009:RPC

[1248] Partha Bhowmick and Bhargab B. Bhattacharya. Real polygonal covers of digital discs — some theories and experiments. Fundamenta Informaticae, 91(3–4):487–505, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Damasevicius:2009:ACG

[1249] Robertas Damaševičius. Analysis of components for generalization using multidimensional scaling. Fundamenta Informaticae, 91(3–4):507–522, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gao:2009:FIO

[1250] Chong zhi Gao and Zheng an Yao. A further improved online/offline signature scheme. Fundamenta Informaticae, 91(3–4):523–532, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gillibert:2009:HGL

[1251] Luc Gillibert and Alain Bretto. Hypergraphs for generic lossless image compression. Fundamenta Informaticae, 91(3–4):533–546, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hristea:2009:ASD

[1252] Florentina Hristea and Marius Popescu. Adjective sense disambiguation at the border between unsupervised and knowledge-based techniques. Fundamenta Informaticae, 91(3–4):547–562, August 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huang:2009:FVC

[1253] Chang-Chin Huang, Du-Shiau Tsai, and Gwoboa Horng. A fast VQ codebook generation algorithm based on Otsu histogram threshold. Fundamenta Informaticae, 91(3–4):563–579, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2009:RDH

[1254] Ching-Chiuan Lin, Shun-Ping Yang, and Nien-Lin Hsueh. Reversible data hiding based on three-circular-pixel difference expansion. Fundamenta Informaticae, 91(3–4):581–595, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lu:2009:IHS

[1255] Tzu-Chuen Lu, Chin-Chen Chang, and Yi-Long Liu. An informationhiding scheme based on quantization-based embedding technique. Fundamenta Informaticae, 91(3–4):597–610, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moreira:2009:SPA

[1256] Nelma Moreira and Rogério Reis. Series-parallel automata and short regular expressions. Fundamenta Informaticae, 91(3–4):611–629, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mu:2009:HID

[1257] Kedian Mu, Weiru Liu, Zhi Jin, Ruqian Lu, Anbu Yue, and David Bell. Handling inconsistency in distributed software requirements specifications based on prioritized merging. Fundamenta Informaticae, 91(3–4): 631–670, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bouyer:2009:URT

[1258] Patricia Bouyer, Serge Haddad, and Pierre-Alain Reynier. Undecidability results for timed automata with silent transitions. Fundamenta Informaticae, 92(1–2):1–25, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2009:PCC

[1259] Chin-Chen Chang, Tzung-Her Chen, and Li-Jen Liu. Preventing cheating in computational visual cryptography. Fundamenta Informaticae, 92(1– 2):27–42, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czerniak:2009:EAD

[1260] Jacek Czerniak. Evolutionary approach to data discretization for rough sets theory. Fundamenta Informaticae, 92(1–2):43–61, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fredriksson:2009:SRA

[1261] Kimmo Fredriksson and Fedor Nikitin. Simple random access compression. Fundamenta Informaticae, 92(1–2):63–81, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hartmann:2009:IWM

[1262] Sven Hartmann and Sebastian Link. On inferences of weak multivalued dependencies. Fundamenta Informaticae, 92(1–2):83–102, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hogberg:2009:BMW

[1263] Johanna Högberg, Andreas Maletti, and Heiko Vogler. Bisimulation minimisation of weighted automata on unranked trees. Fundamenta Informaticae, 92(1–2):103–130, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hung:2009:FBA

[1264] Kuo Lung Hung and Shin-Wei He. Feature based affine invariant watermarking robust to geometric distortions. Fundamenta Informaticae, 92(1–2):131–143, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Verma:2009:CNF

[1265] Rakesh Verma. Complexity of normal form properties and reductions for term rewriting problems complexity of normal form properties and reductions for term rewriting problems. Fundamenta Informaticae, 92(1– 2):145–168, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Waszkiewicz:2009:DTG

[1266] Paweł Waszkiewicz. On domain theory over Girard quantales. Fundamenta Informaticae, 92(1–2):169–192, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Foustoucos:2009:MSO

[1267] Eugénie Foustoucos and Labrini Kalantzi. The monadic second-order logic evaluation problem on finite colored trees: a database-theoretic approach. Fundamenta Informaticae, 92(3):193–231, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holscher:2009:AUM

[1268] Karsten Hölscher, Hans-Jörg Kreowski, and Sabine Kuske. Autonomous units to model interacting sequential and parallel processes. Fundamenta Informaticae, 92(3):233–257, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2009:GAC

[1269] Mikhail Ju. Moshkov, Marcin Piliszczuk, and Beata Zielosko. Greedy algorithm for construction of partial association rules. Fundamenta Informaticae, 92(3):259–277, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sroka:2009:TFSa

[1270] Jacek Sroka and Jan Hidders. Towards a formal semantics for the process model of the Taverna Workbench. Part I. Fundamenta Informaticae, 92 (3):279–299, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yinbin:2009:PPC

[1271] Lei Yinbin and Luo Maokang. A proof of Plotkin’s conjecture. Fundamenta Informaticae, 92(3):301–306, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Conradie:2009:ACC

[1272] Willem Conradie, Valentin Goranko, and Dimiter Vakarelov. Algorithmic correspondence and completeness in modal logic. III. Extensions of the algorithm SQEMA with substitutions. Fundamenta Informaticae, 92(4):

307–343, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halder:2009:APD

[1273] Anindya Halder, Ashish Ghosh, and Susmita Ghosh. Aggregation pheromone density based pattern classification. Fundamenta Informaticae, 92(4):345–362, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Spatharis:2009:CTB

[1274] Anthony Spatharis, Ilias Foudalis, Martha Sideri, and Christos Papadimitriou. Comparing trade-off based models of the Internet. Fundamenta Informaticae, 92(4):363–372, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sroka:2009:TFSb

[1275] Jacek Sroka and Jan Hidders. Towards a formal semantics for the process model of the Taverna Workbench. Part II. Fundamenta Informaticae, 92 (4):373–396, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2009:CLW

[1276] Cheng-Hsing Yang, Chi-Yao Weng, Shiuh-Jeng Wang, and Hung-Min Sun. Codebook-linked watermarking scheme for digital images. Fundamenta Informaticae, 92(4):397–409, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhong:2009:DA

[1277] Sheng Zhong. On distributed k-anonymization. Fundamenta Informaticae, 92(4):411–431, June 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:CSP

[1278] Anonymous. Concurrency specification and programming (CS&P). Fundamenta Informaticae, 93(1–3):??, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balicki:2009:FDX

[1279] Krzysztof Balicki and Marcin Szpyrka. Formal definition of XCCS modelling language. Fundamenta Informaticae, 93(1–3):1–15, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2009:STD

[1280] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Simone Tini. P systems with transport and diffusion membrane channels. Fundamenta Informaticae, 93(1–3):17–31, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barylska:2009:LPP

[1281] Kamila Barylska and Edward Ochmański. Levels of persistency in place/ transition nets. Fundamenta Informaticae, 93(1–3):33–43, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2009:JSI

[1282] Marco Bellia and M. Eugenia Occhiuto. JavaΩ: The structures and the implementation of a preprocessor for Java with m and mc parameters. Fundamenta Informaticae, 93(1–3):45–64, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Budzynska:2009:PSA

[1283] Katarzyna Budzyńska, Magdalena Kacprzak, and Paweł Rembelski. Perseus. Software for analyzing persuasion process. Fundamenta Informaticae, 93(1–3):65–79, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bulling:2009:WAC

[1284] Nils Bulling and Wojciech Jamroga. What agents can probably enforce. Fundamenta Informaticae, 93(1–3):81–96, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2009:ASN

[1285] Ludwik Czaja and Manfred Kudlek. Analysis and synthesis of net structures and transition graphs. Fundamenta Informaticae, 93(1–3):97–110, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doroshenko:2009:UOA

[1286] Anatoliy Doroshenko and Olena Yatsenko. Using ontologies and algebra of algorithms for formalized development of parallel programs. Fundamenta Informaticae, 93(1–3):111–125, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gribovskaya:2009:TDB

[1287] Natalya Gribovskaya and Irina Virbitskaite. Timed delay bisimulation is an equivalence relation for timed transition systems. Fundamenta Informaticae, 93(1–3):127–142, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grochowalski:2009:OOR

[1288] Piotr Grochowalski and Krzysztof Pancerz. The outline of an ontology for the rough set theory and its applications. Fundamenta Informaticae, 93(1–3):143–154, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2009:QST

[1289] Damas P. Gruska. Quantifying security for timed process algebras. Fundamenta Informaticae, 93(1–3):155–169, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2009:GDT

[1290] Kees M. van Hee, Natalia Sidorova, Marc Voorhoeve, and Jan Martijn van derWerf. Generation of database transactions with Petri nets. Fundamenta Informaticae, 93(1–3):171–184, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jakubowska:2009:SSP

[1291] Gizela Jakubowska, Piotr Dembiński, Wojciech Penczek, and Maciej Szreter. Simulation of security protocols based on scenarios of attacks. Fundamenta Informaticae, 93(1–3):185–203, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2009:ECC

[1292] Michael Kohler-Bußmeier and Frank Heitmann. On the expressiveness of communication channels for object nets. Fundamenta Informaticae, 93(1–3):205–219, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2009:MPA

[1293] Manfred Kudlek, Patrick Totzke, and Georg Zetzsche. Multiset pushdown automata. Fundamenta Informaticae, 93(1–3):221–233, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2009:PML

[1294] Manfred Kudlek, Patrick Totzke, and Georg Zetzsche. Properties of multiset language classes defined by multiset pushdown automata. Fundamenta Informaticae, 93(1–3):235–244, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurkowski:2009:TAB

[1295] Mirosław Kurkowski and Wojciech Penczek. Timed automata based model checking of timed security protocols. Fundamenta Informaticae, 93(1–3):245–259, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2009:MIR

[1296] Mikhail Ju. Moshkov, Andrzej Skowron, and Zbigniew Suraj. On minimal inhibitory rules for almost all k-valued information systems. Fundamenta Informaticae, 93(1–3):261–272, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2009:ETP

[1297] Linh Anh Nguyen. An efficient tableau prover using global caching for the description logic ALC. Fundamenta Informaticae, 93(1–3):273–288, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2009:NAM

[1298] Artur Niewiadomski, Wojciech Penczek, and Maciej Szreter. A new approach to model checking of UML state machines. Fundamenta Informaticae, 93(1–3):289–303, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rataj:2009:TJP

[1299] Artur Rataj, Bożena Woźna, and Andrzej Zbrzezny. A translator of Java programs to TADDs. Fundamenta Informaticae, 93(1–3):305–324, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2009:ACC

[1300] Roman R. Redziejowski. Applying classical concepts to parsing expression grammar. Fundamenta Informaticae, 93(1–3):325–336, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wegener:2009:PNT

[1301] Jan-Thierry Wegener and Louchka Popova-Zeugmann. Petri nets with time windows: A comparison to classical Petri nets. Fundamenta Informaticae, 93(1–3):337–352, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czelakowski:2009:AEC

[1302] Janusz Czelakowski. AC is equivalent to the coherence principle. Corrigendum to my paper “Induction Principles for Sets”. Fundamenta Informaticae, 93(4):353–356, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See [953].

Dinu:2009:IGM

[1303] Liviu P. Dinu. On insertion grammars with maximum parallel derivation. Fundamenta Informaticae, 93(4):357–369, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanGlabbeek:2009:BBE

[1304] Rob van Glabbeek, Bas Luttik, and Nikola Trčka. Branching bisimilarity with explicit divergence. Fundamenta Informaticae, 93(4):371–392, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lei:2009:NSR

[1305] Yuxia Lei, Yuefei Sui, and Cungen Cao. Normalized-scale relations and their concept lattices in relational databases. Fundamenta Informaticae, 93(4):393–409, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paszynski:2009:PSAa

[1306] Maciej Paszyński. On the parallelization of self-adaptive hp-finite element methods Part I. Composite programmable graph grammar model. Fundamenta Informaticae, 93(4):411–434, December 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paszynski:2009:PSAb

[1307] Maciej Paszyński. On the parallelization of self-adaptive hp-finite element methods Part II. Partitioning communication agglomeration mapping (PCAM) analysis. Fundamenta Informaticae, 93(4):435–457, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2009:NAO

[1308] Ling Zhang, Fu gui He, Yan ping Zhang, and Shu Zhao. A new algorithm for optimal path finding in complex networks based on the quotient space. Fundamenta Informaticae, 93(4):459–469, December 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2009:TA

[1309] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Luca Tesei. Timed P automata. Fundamenta Informaticae, 94(1):1–19, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gore:2009:CTM

[1310] Rajeev Goré and Linh Anh Nguyen. Clausal tableaux for multimodal logics of belief. Fundamenta Informaticae, 94(1):21–40, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kondo:2009:APG

[1311] Michiro Kondo. Algebraic properties of generalized rough sets. Fundamenta Informaticae, 94(1):41–48, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Madalinska-Bugaj:2009:FOG

[1312] Ewa Madalińska-Bugaj and Witold Łukaszewicz. First-order generalization of the MPMA belief update operator. Fundamenta Informaticae, 94 (1):49–61, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2009:IGF

[1313] Maurice Margenstern. The injectivity of the global function of a cellular automaton in the hyperbolic plane is undecidable. Fundamenta Informaticae, 94(1):63–99, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2009:GAW

[1314] Mikhail Ju. Moshkov, Marcin Piliszczu, and Beata Zielosko. Greedy algorithms with weights for construction of partial association rules. Fundamenta Informaticae, 94(1):101–120, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:FKT

[1315] Anonymous. Fundamentals of knowledge technology. Fundamenta Informaticae, 94(2):??, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blajdo:2009:ECS

[1316] Piotr Blajdo, Zdzisław S. Hippe, Teresa Mroczek, Jerzy W. Grzymala-Busse, Maksymilian Knap, and Lukasz Piatek. An extended comparison of six approaches to discretization — a rough set approach. Fundamenta Informaticae, 94(2):121–131, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chan:2009:DBR

[1317] Chien-Chung Chan and Gwo-Hshiung Tzeng. Dominance-based rough sets using indexed blocks as granules. Fundamenta Informaticae, 94(2): 133–146, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2009:AAF

[1318] Davide Ciucci. Approximation algebra and framework. Fundamenta Informaticae, 94(2):147–161, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dembczynski:2009:LRE

[1319] Krzysztof Dembczyński, Wojciech Kotłowski, and Roman Słowiński. Learning rule ensembles for ordinal classification with monotonicity constraints. Fundamenta Informaticae, 94(2):163–178, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gong:2009:SIB

[1320] Xun Gong, Guoyin Wang, and Lili Xiong. Single 2D image-based 3D face reconstruction and its application in pose estimation. Fundamenta Informaticae, 94(2):179–195, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2009:PCB

[1321] Ryszard Janicki. Pairwise comparisons based non-numerical ranking. Fundamenta Informaticae, 94(2):197–217, April 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2009:TPM

[1322] Huaxiong Li, Yiyu Yao, Xianzhong Zhou, and Bing Huang. A two-phase model for learning rules from incomplete data. Fundamenta Informaticae, 94(2):219–232, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lingras:2009:SSR

[1323] Pawan Lingras, Min Chen, and Duoqian Miao. Semi-supervised rough cost/benefit decisions. Fundamenta Informaticae, 94(2):233–244, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2009:IAI

[1324] Dun Liu, Tianrui Li, Da Ruan, and Weili Zou. An incremental approach for inducing knowledge from dynamic information systems. Fundamenta Informaticae, 94(2):245–260, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Song:2009:ICT

[1325] Jing Song, Tianrui Li, and Da Ruan. An integration of cloud transform and rough set theory to induction of decision trees. Fundamenta Informaticae, 94(2):261–273, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:FIP

[1326] Anonymous. Fundamenta Informaticae (2009) Petri Nets 2008. Fundamenta Informaticae, 94(3–4):??, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Darondeau:2009:SNS

[1327] Philippe Darondeau, Maciej Koutny, Marta Pietkiewicz-Koutny, and Alex Yakovlev. Synthesis of nets with step firing policies. Fundamenta Informaticae, 94(3–4):275–303, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bonchi:2009:NBA

[1328] Filippo Bonchi, Antonio Brogi, Sara Corfini, and Fabio Gadducci. A netbased approach to Web services publication and replaceability. Fundamenta Informaticae, 94(3–4):305–330, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenthum:2009:USP

[1329] Robin Bergenthum, Sebastian Mauser, Robert Lorenz, and Gabriel Juhas. Unfolding semantics of Petri nets based on token flows. Fundamenta Informaticae, 94(3–4):331–360, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kristensen:2009:MVS

[1330] Lars M. Kristensen and Paul Fleischer. Modelling and validation of secure connection establishment in a generic access network scenario. Fundamenta Informaticae, 94(3–4):361–386, August 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderWerf:2009:PDU

[1331] J. M. E. M. van derWerf, B. F. van Dongen, C. A. J. Hurkens, and A. Serebrenik. Process discovery using integer linear programming. Fundamenta Informaticae, 94(3–4):387–412, August 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamez:2009:BEM

[1332] Alexandre Hamez, Yann Thierry-Mieg, and Fabrice Kordon. Building efficient model checkers using hierarchical set decision diagrams and automatic saturation. Fundamenta Informaticae, 94(3–4):413–437, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meyer:2009:PAV

[1333] Roland Meyer, Victor Khomenko, and Tim Strazny. A practical approach to verification of mobile systems using net unfoldings. Fundamenta Informaticae, 94(3–4):439–471, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:ACS

[1334] Anonymous. Application of concurrency to system design, the seventh special issue. Fundamenta Informaticae, 95(1):??, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wist:2009:AIC

[1335] Dominic Wist, Ralf Wollowski, Mark Schaefer, and Walter Vogler. Avoiding irreducible CSC conflicts by internal communication. Fundamenta Informaticae, 95(1):1–29, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Toosizadeh:2009:ACA

[1336] Navid Toosizadeh and Safwat G. Zaky. Application of concurrency in the asynchronous design of write-after-read operations. Fundamenta Informaticae, 95(1):31–52, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2009:TMA

[1337] Shufan Yang, Steve B. Furber, Yebin Shi, and Luis A. Plana. A tokenmanaged admission control system for QoS provision on a best-effort GALS interconnect. Fundamenta Informaticae, 95(1):53–72, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guidi:2009:DEH

[1338] Claudio Guidi, Ivan Lanese, Fabrizio Montesi, and Gianluigi Zavattaro. Dynamic error handling in service oriented applications. Fundamenta Informaticae, 95(1):73–102, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meng:2009:QDS

[1339] Sun Meng and Farhad Arbab. QoS-driven service selection and composition using quantitative constraint automata. Fundamenta Informaticae, 95(1):103–128, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hahn:2009:TBM

[1340] E. Moritz Hahn, Holger Hermanns, Björn Wachter, and Lijun Zhang. Time-bounded model checking of infinite-state continuous-time Markov chains. Fundamenta Informaticae, 95(1):129–155, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Markovski:2009:PED

[1341] Jasen Markovski and Erik P. de Vink. Performance evaluation of distributed systems based on a discrete real- and stochastic-time process algebra. Fundamenta Informaticae, 95(1):157–186, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenthum:2009:SPN

[1342] Robin Bergenthum, Jörg Desel, Sebastian Mauser, and Robert Lorenz. Synthesis of Petri nets from term based representations of infinite partial languages. Fundamenta Informaticae, 95(1):187–217, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Madalinski:2009:MCF

[1343] Agnes Madalinski and Eric Fabre. Modular construction of finite and complete prefixes of Petri net unfoldings. Fundamenta Informaticae, 95 (1):219–244, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amirani:2009:ETA

[1344] Mehdi Chehel Amirani and Ali Asghar Beheshti Shirazi. Evaluation of the texture analysis using spectral correlation function. Fundamenta Informaticae, 95(2–3):245–262, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Byrski:2009:SME

[1345] Aleksander Byrski and Robert Schaefer. Stochastic model of evolutionary and immunological multi-agent systems: Mutually exclusive actions. Fundamenta Informaticae, 95(2–3):263–285, April 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2009:RTL

[1346] Olivier Finkel and Pierre Simonnet. On recognizable tree languages beyond the Borel hierarchy. Fundamenta Informaticae, 95(2–3):287–303, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mirkowska:2009:VCC

[1347] Grażyna Mirkowska, Andrzej Salwicki, and Oskar Świda. Verifying a class: combining testing and proving. Fundamenta Informaticae, 95(2– 3):305–324, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schaefer:2009:SME

[1348] Robert Schaefer, Aleksander Byrski, and Maciej Smołka. Stochastic model of evolutionary and immunological multi-agent systems: Parallel execution of local actions. Fundamenta Informaticae, 95(2–3):325– 348, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Trojanowski:2009:PQP

[1349] Krzysztof Trojanowski. Properties of quantum particles in multi-swarms for dynamic optimization. Fundamenta Informaticae, 95(2–3):349–380, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vagvolgyi:2009:CGE

[1350] Sándor Vágvölgyi. Congruences generated by extended ground term rewrite systems. Fundamenta Informaticae, 95(2–3):381–399, April 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:SI

[1351] Anonymous. Swarm intelligence. Fundamenta Informaticae, 95(4):??, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:2009:IHS

[1352] Prithwish Chakraborty, Gourab Ghosh Roy, Swagatam Das, Dhaval Jain, and Ajith Abraham. An improved harmony search algorithm with differential mutation operator. Fundamenta Informaticae, 95(4):401–426, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cui:2009:IPS

[1353] Zhihua Cui and Xingjuan Cai. Integral particle swarm optimization with dispersed accelerator information. Fundamenta Informaticae, 95(4):427– 447, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kwolek:2009:PSO

[1354] Bogdan Kwolek. Particle swarm optimization based object tracking. Fundamenta Informaticae, 95(4):449–463, December 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2009:MSA

[1355] Hongbo Liu, Ajith Abraham, and Zuwen Wang. A multi-swarm approach to multi-objective flexible job-shop scheduling problems. Fundamenta Informaticae, 95(4):465–489, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2009:RCP

[1356] Ming-Tsung Liu and Pao-Ta Yu. Robust candidate pruning approach based on the PSO-SVM for fast corner detection with noise tolerance in gray-level images. Fundamenta Informaticae, 95(4):491–510, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pant:2009:LDI

[1357] Millie Pant, Radha Thangaraj, and Ajith Abraha. Low discrepancy initialized particle swarm optimization for solving constrained optimization problems. Fundamenta Informaticae, 95(4):511–531, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ramanna:2009:MRB

[1358] Sheela Ramanna and Amir H. Meghdadi. Measuring resemblances between swarm behaviours: A perceptual tolerance near set approach. Fundamenta Informaticae, 95(4):533–552, December 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bandyopadhyay:2009:MLD

[1359] Sanghamitra Bandyopadhyay and Malay Bhattacharyya. Mining the largest dense vertexlet in a weighted scale-free graph. Fundamenta Informaticae, 96(1–2):1–25, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2009:ISD

[1360] Jan A. Bergstra and Cornelis A. Middelburg. Instruction sequences with dynamically instantiated instructions. Fundamenta Informaticae, 96(1– 2):27–48, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chan:2009:ULM

[1361] Chi-Shiang Chan. On using LSB matching function for data hiding in pixels. Fundamenta Informaticae, 96(1–2):49–59, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2009:FCV

[1362] Tzung-Her Chen, Kai-Hsiang Tsao, and Yan-Ting Yang. Friendly color visual secret sharing by random grids. Fundamenta Informaticae, 96(1– 2):61–70, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2009:SMC

[1363] Jürgen Dassow, Gema M. Martín, and Francisco J. Vico. A similarity measure for cyclic unary regular languages. Fundamenta Informaticae, 96 (1–2):71–88, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janodet:2009:BCB

[1364] Jean-Christophe Janodet, Marc Sebban, and Henri-Maxime Suchier. Boosting classifiers built from different subsets of features. Fundamenta Informaticae, 96(1–2):89–109, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutras:2009:MEA

[1365] Costas D. Koutras and Yorgos Zikos. On a modal epistemic axiom emerging from McDermott–Doyle logics. Fundamenta Informaticae, 96(1–2): 111–125, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lanotte:2009:DPL

[1366] Ruggero Lanotte and Danièle Beauquier. A decidable probability logic for timed probabilistic systems. Fundamenta Informaticae, 96(1–2):127–151, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rital:2009:HCU

[1367] Soufiane Rital. Hypergraph cuts & unsupervised representation for image segmentation. Fundamenta Informaticae, 96(1–2):153–179, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudnicki:2009:NSI

[1368] Witold R. Rudnicki, Aleksander Jankowski, Aleksander Modzelewski, Aleksander Piotrowski, and Adam Zadrożny. The new SIMD implementation of the Smith–Waterman algorithm on cell microprocessor. Fundamenta Informaticae, 96(1–2):181–194, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2009:SGS

[1369] Sriparna Saha and Sanghamitra Bandyopadhyay. Semi-GAPS: A semisupervised clustering method using point symmetry. Fundamenta Informaticae, 96(1–2):195–209, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shen:2009:SMM

[1370] Guohua Shen, Zhiqiu Huang, Yuping Zhang, Xiaodong Zhu, and Jun Yang. A semantic model for matchmaking of Web services based on description logics. Fundamenta Informaticae, 96(1–2):211–226, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:ACL

[1371] Anonymous. Advances in computational logic (CILC08). Fundamenta Informaticae, 96(3):??, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenti:2009:IFD

[1372] Federico Bergenti, Alessandro Dal Palù, and Gianfranco Rossi. Integrating finite domain and set constraints into a set-based constraint language. Fundamenta Informaticae, 96(3):227–252, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bozzato:2009:ACS

[1373] Loris Bozzato, Mauro Ferrari, and Paola Villa. Actions over a constructive semantics for description logics. Fundamenta Informaticae, 96(3): 253–269, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bria:2009:NFN

[1374] Annamaria Bria, Wolfgang Faber, and Nicola Leone. Normal form nested programs. Fundamenta Informaticae, 96(3):271–295, August 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DalPalu:2009:GAS

[1375] Alessandro Dal Palù, Agostino Dovier, Enrico Pontelli, and Gianfranco Rossi. GASP: Answer set programming with lazy grounding. Fundamenta Informaticae, 96(3):297–322, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferrante:2009:MCG

[1376] Alessandro Ferrante, Margherita Napoli, and Mimmo Parente. Model checking for graded CTL. Fundamenta Informaticae, 96(3):323–339, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordano:2009:APE

[1377] Laura Giordano, Nicola Olivetti, Valentina Gliozzic, and Gian Luca Pozzato. ALC + T: a preferential extension of description logics. Fundamenta Informaticae, 96(3):341–372, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Senni:2009:FRE

[1378] Valerio Senni, Alberto Pettorossi, and Maurizio Proietti. A folding rule for eliminating existential variables from constraint logic programs. Fundamenta Informaticae, 96(3):373–393, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:AAI

[1379] Anonymous. Advances in artificial intelligence and applications. Fundamenta Informaticae, 96(4):??, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rhee:2009:MSC

[1380] Sang Keun Rhee, Jihye Lee, Myon-Woong Park, Michał Szymczak, Grzegorz Frackowiak, Maria Ganzha, and Marcin Paprzycki. Measuring semantic closeness of ontologically demarcated resources. Fundamenta Informaticae, 96(4):395–418, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

ElJelali:2009:DMB

[1381] Soufiane El Jelali, Abdelouahid Lyhyaoui, and Aníbal R. Figueiras-Vidal. Designing model based classifiers by emphasizing soft targets. Fundamenta Informaticae, 96(4):419–433, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paradowski:2009:IRW

[1382] Mariusz Paradowski and Halina Kwasnicka. Improved resulted word counts optimizer for automatic image annotation problem. Fundamenta Informaticae, 96(4):435–463, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brzostowski:2009:DAM

[1383] Krzysztof Brzostowski and Jerzy Światek. Different approaches to model relationship between EMG signals and force moments in human skeletal muscle. Analysis for diagnosis of neuronmuscular disorders. Fundamenta Informaticae, 96(4):465–475, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kotowski:2009:SCT

[1384] Stefan Kotowski, Witold Kosiński, Zbigniew Michalewicz, Piotr Synak, and Łukasz Brocki. On stability and classification tools for genetic algorithms. Fundamenta Informaticae, 96(4):477–491, December 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diqun:2009:QSP

[1385] Yan Diqun, Wang Rangding, and Zhang Liguang. Quantization step parity-based steganography for MP3 audio. Fundamenta Informaticae, 97(1–2):1–14, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamzeh:2009:RCS

[1386] Ali Hamzeh, Sattar Hashemi, Ashkan Sami, and Adel Rahmani. A recursive classifier system for partially observable environments. Fundamenta Informaticae, 97(1–2):15–40, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:2009:SON

[1387] Maciej Koutny and Brian Randell. Structured occurrence nets: A formalism for aiding system failure prevention and analysis techniques. Fundamenta Informaticae, 97(1–2):41–91, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2009:EME

[1388] Meng-Tsan Li, Nien-Ching Huang, Kuo-Chen Wu, Chin-Kai Jan, and Chung-Ming Wang. An effective message embedding scheme for 3D models. Fundamenta Informaticae, 97(1–2):93–109, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Magnin:2009:EPNa

[1389] Morgan Magnin, Pierre Molinaro, and Olivier (H.) Roux. Expressiveness of Petri nets with stopwatches. Dense-time part. Fundamenta Informaticae, 97(1–2):111–138, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Magnin:2009:EPNb

[1390] Morgan Magnin, Pierre Molinaro, and Olivier (H.) Roux. Expressiveness of Petri nets with stopwatches. Discrete-time part. Fundamenta Informaticae, 97(1–2):139–176, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mani:2009:ASS

[1391] A. Mani. Algebraic semantics of similarity-based bitten rough set theory. Fundamenta Informaticae, 97(1–2):177–197, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2009:MBI

[1392] Sriparna Saha and Sanghamitra Bandyopadhyay. MR brain image segmentation using a multi-seed based automatic clustering technique. Fundamenta Informaticae, 97(1–2):199–214, January 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2009:TFC

[1393] Suman Saha, C. A. Murthy, and Sankar K. Pal. Tensor framework and combined symmetry for hypertext mining. Fundamenta Informaticae, 97(1–2):215–234, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2009:AFDa

[1394] Józef Winkowski. An algebraic framework for defining behaviours of concurrent systems. Part 1: The constructive presentation. Fundamenta Informaticae, 97(1–2):235–273, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zeng:2009:HSN

[1395] Xiangxiang Zeng, Xingyi Zhang, and Linqiang Pan. Homogeneous spiking neural P systems. Fundamenta Informaticae, 97(1–2):275–294, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2009:SIS

[1396] Anonymous. Special issue on stringology. Fundamenta Informaticae, 97 (3):??, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Daykin:2009:CUM

[1397] David E. Daykin, Jacqueline W. Daykin, and W. F. (Bill) Smyth. Combinatorics of unique maximal factorization families (UMFFs). Fundamenta Informaticae, 97(3):295–309, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Iliopoulos:2009:FAC

[1398] Costas S. Iliopoulos, W. F. Smyth, and Munina Yusufu. Faster algorithms for computing maximal multirepeats in multiple sequences. Fundamenta Informaticae, 97(3):311–320, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gagie:2009:NAB

[1399] Travis Gagie. A new algorithm for building alphabetic minimax trees. Fundamenta Informaticae, 97(3):321–329, August 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Allali:2009:TGF

[1400] Julien Allali, Pascal Ferraro, Pierre Hanna, Costas Iliopoulos, and Matthias Robine. Toward a general framework for polyphonic comparison. Fundamenta Informaticae, 97(3):331–346, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2009:CAA

[1401] Ivo Düntsch and Ian Pratt-Hartmann. Complex algebras of arithmetic. Fundamenta Informaticae, 97(4):347–367, December 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kieu:2009:IAB

[1402] The Duc Kieu and Chin-Chen Chang. An image authentication based on discrete Fourier transform. Fundamenta Informaticae, 97(4):369–379, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suresh:2009:DCU

[1403] Kaushik Suresh, Debarati Kundu, Sayan Ghosh, Swagatam Das, and Ajith Abraham. Data clustering using multi-objective differential evolution algorithms. Fundamenta Informaticae, 97(4):381–403, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vitoria:2009:MRP

[1404] Aida Vitória, Jan Małuszyński, and Andrzej Szałas. Modeling and reasoning with paraconsistent rough sets. Fundamenta Informaticae, 97(4): 405–438, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2009:AFDb

[1405] Józef Winkowski. An algebraic framework for defining behaviours of concurrent systems. Part 2: The axiomatic presentation. Fundamenta Informaticae, 97(4):439–470, December 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2009:IVI

[1406] Zhiming Zhang. An interval-valued intuitionistic fuzzy rough set model. Fundamenta Informaticae, 97(4):471–498, December 2009. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).